### v1

Abstract
Prompt-tuning has demonstrated parameter-efficiency in fusing unimodal foundation models for multimodal tasks. However, its limited adaptivity and expressiveness lead to suboptimal performance when compared with other tuning methods. In this paper, we address this issue by disentangling the vanilla prompts to adaptively capture dataset-level and instance-level features. Building upon this disentanglement, we introduce the mixture of prompt experts (MoPE) technique to enhance expressiveness. MoPE leverages multimodal pairing priors to route the most effective prompt on a per-instance basis. Compared to vanilla prompting, our MoPE-based conditional prompting exhibits greater expressiveness for multimodal fusion, scaling better with the training data and the overall number of trainable parameters. We also study a regularization term for expert routing, leading to emergent expert specialization, where different experts focus on different concepts, enabling interpretable soft prompting. Extensive experiments across three multimodal datasets demonstrate that our method achieves state-of-the-art results, matching or even surpassing the performance of fine-tuning, while requiring only 0.8% of the trainable parameters. Code will be released: https://github.com/songrise/MoPE.
Keywords:
Multimodal Fusion Prompt Tuning Mixture of Experts
1 Introduction
Figure 1: High-level motivation of MoPE-based conditional prompting. a) vanilla prompt tuning learn a global shared prompt for all instances, which may not be optimal for each instance; b) MoPE-based conditional prompt tuning optimize multiple specialized prompt experts and is instance-wise adaptative.
Empowered with billion-scale data and highly scalable model architectures, recent unimodal pre-trained models [35, 28, 3, 4, 17], also known as foundation models, have demonstrated their transferable power to various downstream tasks [26, 30, 51]. However, transferring multimodal foundation models for downstream applications is less flexible. Recent explorations, such as CLIP [34], employ two-tower designs that entangle the two encoders, meaning that replacing either one would necessitate expensive paired pre-training from scratch. This limitation hinders the broader application that would benefit from a specific unimodal architecture, such as longer context windows. To democratize foundation models, a compelling question arises: How can we efficiently combine separately pretrained unimodal foundation models for downstream multimodal tasks?
Recently, prompt tuning [20, 11] has emerged as a novel paradigm for parameter-efficient transfer learning. While originally proposed for low-shot adaptation to tasks with the same modality as they were pretrained on [20, 11, 50], it has also proven effective for multimodal fusion. Typically, this is achieved by abstracting the representation from one modality into one or a few prompts, which are fed into the frozen Transformer of another modality for fusion [39, 22, 21, 45].
Despite the effectiveness of prompt tuning in inter-modality transfer learning, directly adopting it for multimodal fusion could yield suboptimal results, particularly due to its limited adaptivity and scalability. Notably, while prompt tuning generally performs well in low-data regimes, it can be less effective when applied to full-shot training on the entire dataset with a challenging objective [7, 22, 39, 44, 19]. This reduced efficacy could be attributed to at least two factors. First, the vanilla prompt tuning used in previous prompt-based fusion methods [39, 21, 22] is not designed for multimodal tasks. These methods simply optimize a globally shared prompt for all instances [11, 20, 39], neglecting the multimodal interplay and potentially failing to capture the locally optimal representation for each input instance. Second, the relatively small number of trainable parameters (compared with fine-tuning) can lead to underfitting in multimodal datasets with a long-tail distribution and complex cross-modal mapping [27, 32, 42].
To address these challenges, increasing the number of learnable prompts, known as “length-scaling,” appears to be a plausible solution. Despite its apparent simplicity, a lot of research indicates that the performance gains from length-scaling quickly reach saturation in both unimodal [7, 19] and multimodal cases [22, 39, 44]. Furthermore, over-length prompts may even lead to worse results [11, 44, 14, 20, 16]. Recent theoretical analyses have substantiated these empirical observations, highlighting the limited expressiveness of prompt tuning [42, 32]. A direct implication of their analysis is that simply increasing prompt length is insufficient to address the issues of adaptivity and expressiveness.
Based on the above observations, we aim to enhance the expressiveness of the vanilla prompt used in multimodal fusion to improve adaptivity and scalability. Our motivation is to utilize the multimodal input pairing information as a prior, and to scale up the prompt in another dimension—the number of prompt experts—rather than length, as illustrated in Fig. 1. To accomplish this, we first propose a conditional prompt tuning method, which disentangles the vanilla global-shared prompt vector into static and dynamic prompts to adaptively capture global and local features. The cornerstone of our dynamic prompt generation is our innovative MoPE (Mixture of Prompt Experts) design. In this design, we fix the prompt length used in attention and scale up expressiveness by learning multiple prompt experts and a router at each layer. We then use the representation from another modality as a prior to route the most effective dynamic prompt for each instance, enabling adaptive fusion.
We show that our MoPE-based fusion method alleviates the two issues in previous prompt-based fusion methods by allowing instance-wise adaptive prompting and avoiding over-length deterioration for better scalability. We systematically compare our MoPE-based fusion method with previous fusion methods on three multimodal datasets, UPMC Food-101 [41], SNLI-VE [43] and MM-IMDB [1]. Not only does our method deliver state-of-the-art results on all datasets, it also has higher parameter-efficiency compared with methods based on vanilla prompting, due to our improved adaptivity and expressiveness. Specifically, we achieve results comparable to fine-tuning while requiring only $\sim 0.8\%$ of the trainable parameters, and we outperform previous state-of-the-art (SOTA) prompt-based methods by a significant margin. Additionally, we demonstrate the superior scalability of our method. Our analysis indicates that increasing the number of experts (i.e. “expert-scaling”) tends to be a better alternative than length-scaling, which avoids performance deterioration with an overly-long prompt. In addition to parameter scaling, our method also demonstrates better scalability with increased training data. Finally, as the first endeavor to use prompts as experts, we analyze how different methods of expert routing can affect expert specialization. Thanks to the proposed regularization term, we observe the emergence of specialized prompt experts after end-to-end training, achieving high interpretability compared to black-box conditional prompting. We summarize our contributions as follows:
  • We design a conditional prompt tuning method for multimodal fusion, which disentangles the vanilla prompt as static, dynamic and mapped prompts for better adaptiveness.
  • We introduce the MoPE technique for instance-wise dynamic prompt generation, which scales up the expressiveness of prompt tuning for multimodal fusion.
  • A combination of regularization terms is studied to aid expert specialization.
  • Extensive experiments on UPMC Food-101, SNLI-VE and MM-IMDB datasets demonstrate state-of-the-art performance and parameter-efficiency for multimodal fusion.
2 Related Works
Prompt Tuning for Transfer Learning. Prompt tuning involves learning continuous token embedding as additional input to a frozen pretrained model for transfer learning. This technique was first popularized in NLP [20, 19], then quickly introduced to CV society [11, 2, 25, 23] and multimodal learning [12, 14, 49, 47]. For either modality, researchers empirically report good transfer learning performance in low-shot adaptation, yet its performance is less comparable to fine-tuning when abundant training instances are available [20, 7]. Moreover, increasing prompt length quickly reaches performance saturation, and over-length prompts might have an negative impact on performance [44, 11, 16]. Recent theoretical analysis [32, 42] reveal that the expressiveness of vanilla prompt tuning is lower than fine-tuning and LoRA [10]. In this work, we scale up the expressiveness of vanilla prompting by employing a Mixture of Experts (MoE)-like design.
Multimodal Fusion via Prompt. Apart from adapting models, prompts can also be used to fuse separately pretrained unimodal models for multimodal tasks. This is a different task from the commonly known “multimodal prompt tuning”, which usually refers to using prompts for transfer learning of multimodal pretrained foundation models. Frozen [39] first introduces a method where the visual representation is mapped as a few input tokens to query frozen language models (LM). PromptFuse and BlindPrompt [22] improved upon this by introducing tunable prompts to the LM for cross-model alignment. PICa [45] translate images into discrete text captions to prompt frozen LMs. These methods treat tokens from different modalities independently and lack cross-modal interaction. Recently, PMF [21] introduced an interactive prompt-tuning method, but their interactivity is built on the strong assumption that both encoders are white-box and in the same Transformer architecture. In this work, we further investigate how to efficiently utilize cross-modal features for prompting while maintaining high modularity by treating one modality as a black box.
MoE in Transformers. MoE is a powerful technique to scale up models, including Transformers [18, 36, 31, 6]. The fundamental approach involves inserting MoE layers, composed of multiple feed-forward networks (FFNs) acting as experts, into the standard Transformer architecture. A router is learned to route each token embedding to the most suitable expert(s). Due to the significant computational cost associated with FFN forwarding, a sparse gating function is typically employed to limit the number of experts used per token [33, 36, 38]. Thanks to those designs, the MoE layer can efficiently scale up model capacity while maintaining computational efficiency. Inspired by previous FFN-based MoE, we explore scaling up the vanilla prompting with an MoE design.
3 Method
Our objective is to adaptively and efficiently fuse unimodal pretrained models for solving multimodal tasks via prompt tuning. In this section, we first review the working mechanism and the limitation of vanilla prompt tuning in Sec. 3.1. In Sec. 3.2, we explain how we achieve conditional prompt-tuning via a sequential pipeline and disentangled prompt tuning. In Sec. 3.3 we introduce the proposed MoPE, consisting of multiple prompt experts and a mutimodal router for expressiveness scaling. Finally, we introduce the regularization methods in Sec. 3.4.
Figure 2: Architecture overview. We illustrate the proposed MoPE-based multimodal fusion method when applied to one frozen Transformer layer. a) We sequentially fuse two modalities, using modality y to guide the prompting of modality x through a three-type prompt disentanglement process. (b) MoPE is introduced, utilizing the multimodal pairing as a prior to instance-wisely generate the dynamic prompt; (c) Inside the multimodal router, we project the representation from each modalities. The concatenated feature (along channel dimension) is used to match the routing embedding paired with each experts for routing score calculation. Better viewed with color.
3.1 Preliminary: Vanilla Prompt Tuning
Consider a Transformer [40] or its variants used to extract features from an embedded input sequence $\mathbf{x}^{0}\in\mathbb{R}_{s\times d_{x}}$, where $s$ is the sequence length and $d_{x}$ is the embedding dimension of the Transformer. Prompt-tuning [^1] In this paper we assume that soft prompts are applied to all layers, which corresponds to “VPT-deep” in CV and “prefix-tuning” in NLP. freezes all pre-trained model weights and optimize a small amount of continuous embeddings (i.e., prompts) $\mathbf{P}\in\mathbb{R}_{l\times d}$ inserted into the input, where $l$ is the number of prompts or “prompt length”. The input of the $i$-th layer layer $L^{i}$ could be denoted as:
$\mathbf{\hat{x}}^{i}=[x^{i-1}_{0},\mathbf{P},\mathbf{T}^{i-1}]$ (1)
where $x^{i-1}_{0}\in\mathbb{R}_{d_{x}}$ denotes the [CLS] token, $\mathbf{T}^{i-1}\in\mathbb{R}_{s\times d_{x}}$ is the token embedding from the previous layer, and [,][,][ , ] denotes the concatenation operation.
Succinctly, prompt tuning works by biasing the pretrained attention pattern in each Transformer layers [32]. As opposed to fine-tuning, this biasing has strictly limited expressiveness [32, 42], meaning that there are tasks that are un-learnable even with $l\to\infty$. The empirical expressiveness of vanilla prompting is even lower than its theoretical upper bound, due to competing optimization when $l>1$, as postulated by Petrov et al. [32]. In other words, although longer prompts provide the model with more capability, finding such a long prompt becomes challenging, and a suboptimal long prompt can lead to performance degradation. In this paper, we do not aim to increase the theoretical upper bound of prompt-tuning; instead, we approach the problem by narrowing the gap between its empirical and theoretical expressiveness.
3.2 Conditional Prompt Tuning for Multimodal Fusion
Previous prompt-based multimodal fusion methods [22, 21] optimize a global prompt that is shared across all instances. However, this approach neglects the interplay between different modalities. With paired multimodal input as a prior, we aim to condition the prompting of one modality on the other for more adaptive and efficient transfer learning.
Sequential Pipeline. A sequential pipeline for multimodal fusion is adopted to achieve prompt conditioning. Specifically, let $(x,y)$ represent a pair of multimodal inputs, and $\mathcal{E}_{y}$ be the encoder of modality $y$. Depending on the intrinsics of the specific task, a fusion direction could be assigned from the complementary modality to the main modality, and the proposed method will be used for tuning the main modality only. Without loss of generality, we consider $x$ to be the main modality in the following discussion.
Prompt Disentanglement. Based on the sequential fusing pipeline, we disentangle the vanilla prompt vector $\mathbf{P}$ used in Eqn. 1 into three types of specialized prompts $[\mathbf{P}_{s},\mathbf{P}_{d},P_{m}]$. The static prompt $\mathbf{P}_{s}\in\mathbb{R}_{l\times d_{x}}$ is a globally-shared prompt vector that is non-conditional to input. To synthesize the dynamic prompt $\mathbf{P}_{d}\in\mathbb{R}_{l\times d_{x}}$, we first extract the global-level feature from the complementary modality $\mathcal{E}_{y}(y)=\psi_{y}\in\mathbb{R}_{d_{y}}$, and utilize a MoPE module $R(\cdot)$ to synthesize the instance-wise dynamic prompt (entailed in Sec 3.3). Additionally, we apply a lightweight mapper $f_{m}(\cdot)$ to map the complementary feature $\psi_{y}$ into a single prompt $P_{m}\in\mathbb{R}_{d_{x}}$, which injects fine-grained cross-modal information. In summary, the input of layer $L^{i}$ of $\mathcal{E}_{x}$ becomes:
$\mathbf{\hat{x}}^{i}=[x^{i-1}_{0},\underbrace{\mathbf{P}_{s},R(\psi_{y}),f_{m}(\psi_{y})}_{\text{Disentangled prompts}},\mathbf{T}^{i-1}]$ (2)
It is important to note that our three types of prompting serve as a plug-in module to replace the vanilla prompt vector of the main modality, while treating the encoder of the complementary modality $\mathcal{E}_{y}$ as a black-box for modularity. The whole process is illustrated in Fig. 2-(a).
Figure 3: Interpretable prompting through MoPE routing. Different experts in MoPE learn to specialize in specific types concepts. In this example, expert-4 is specialized for children while expert-12 focuses on crowds.
3.3 Mixture of Prompt Experts
As length-scaling has been proven ineffective due to quick saturation and over-length deterioration, we propose to fix the prompt length used in attention to suppress the negative impact of longer prompts. To inject more learnable parameters into the framework and increase the expressiveness of the prompts, we condition the prompts on multiple prompt experts for end-to-end training.
Prompt as Expert. We randomly initialize $k$ prompt experts and their associated routing embedding $\{(\mathbf{E}_{j},\mathbf{v}_{j})\}_{j=1}^{k}$ to be optimized end-to-end in each prompt-tuned Transformer layer $L^{i}$. For each pair, $\mathbf{E}_{i}\in\mathbb{R}_{l\times d}$ is the expert and $\mathbf{v}_{i}\in\mathbb{R}_{d_{r}}$ is the routing embedding, where $d_{r}$ is dimension of routing embedding.
Multimodal Expert Routing. In multimodal tasks, missing information from either modality can lead to ambiguous and incorrect results. Therefore, we route each instance based on representations from both modalities. Specifically, for layer $L^{i}$, we apply a layer-wise learned linear transformation $\mathbf{W}_{y}^{i}\in\mathbb{R}_{d_{y}\times d_{c}}$ to get the cross-modal routing embedding based on cross-modal feature $\psi_{y}$, where $d_{c}$ is the embedding dimension for cross-modal routing. We also utilize the global-level feature from previous layer $L^{i-1}$ to compute a inter-modal routing embedding through $\mathbf{W}_{x}^{i}\in\mathbb{R}_{d_{x}\times d_{i}}$, where $d_{i}$ is the inter-modal routing embedding, and $d_{i}+d_{c}=d_{r}$. Finally, we concatenate both embeddings along the embedding dimension to get an adjoint embedding $\mathbf{k}\in\mathbb{R}_{d_{r}}$, and dot product them with the routing embedding of all available experts. The routing score $\mathbf{r}$ is calculated by:
$\mathbf{k}=[\psi_{y}\mathbf{W}_{y}^{i},x_{0}^{i}\mathbf{W}_{x}^{i}],\quad\mathbf{r}_{j}=\frac{\operatorname{exp}(\mathbf{k}^{\top}\mathbf{v}_{j}/\tau+\epsilon)}{\sum_{n=1}^{k}\operatorname{exp}(\mathbf{k}^{\top}\mathbf{v}_{n}/\tau+\epsilon)}$ (3)
where $\tau=0.1$ is the temperature hyper-parameter, and $\epsilon$ is sampled noise [36]. The dynamic prompt is calculated by a convex combination of all experts at this layer according to the routing score:
$\mathbf{P}_{d}=\sum_{j=1}^{k}\mathbf{r}_{j}\mathbf{E}_{j}$ (4)
3.4 Regularizing Expert Routing
The proposed MoPE scales up vanilla prompt tuning by dividing the problem space into subspaces governed by specialized experts. However, we empirically found that a few experts tend to dominate across all instances during training, resulting in non-specialization. This phenomenon is self-reinforcing and is due to degenerated routing, where experts receiving high scores during initialization are exposed to larger gradient scales, causing them to be selected more frequently during optimization [5, 38]. In this section, we introduce methods to circumvent this issue and aid expert specialization.
Orthogonal Routing Embedding. Inspired by the analysis of degenerated routing, we freeze the routing embeddings to avoid self-reinforcement. We find that an orthogonal initialized [37] and non-learnable routing embedding $\{\mathbf{v}\}_{j=1}^{k}$ performs as well as (and even slightly better than) learned embeddings, while requiring fewer trainable parameters.
Importance Loss. To aid expert specialization, we add an additional importance loss [38, 36] to penalize dominant experts. For a batch of input $\mathbf{B}=\{(x_{0},y_{0}),(x_{1},y_{1}),...,(x_{b},y_{b})\}$, the importance of expert-$j$ is defined as the summed routing score in this batch. The importance loss is calculated as the mean coefficient of variation of all experts’ importance averaged across all layers::
$\operatorname{Imp}(\mathbf{E}_{j})=\sum_{(x,y)\in\mathbf{B}}\mathbf{r}_{j},\quad\mathcal{L}_{imp}=\operatorname{stopgrad}\left(\left(\frac{\operatorname{std}(\{\operatorname{Imp}(\mathbf{E}_{j})\}_{j}^{k})}{\operatorname{mean}(\{\operatorname{Imp}(\mathbf{E}_{j})\}_{j}^{k})}\right)^{2};\gamma\right)$ (5)
where $\operatorname{stopgrad}(\cdot)$ is the stop-gradient operator to prevent error propagation of this loss term when the coefficient of variation is less than a pre-defined threshold $\gamma=0.1$. While this loss was initially proposed for balancing computational budgets [38, 36], we adapt its use for promoting expert specialization. We add an additional threshold constraint due to our instance-wise routing, which is more likely to have a larger coefficient of variation than per-token routing.
4 Experiments
4.1 Implementation Details
Architecture Details. For all experiments, unless otherwise specified, we use Swin-B [28] as the vision encoder and Bert-base [3] as the text encoder. Vision is used as main modality. Following the experiment setup in [11, 21], we also finetune a linear head for each dataset. We implement the mapper $f_{m}(\cdot)$ as a two-layer MLP with GeLU [9] nonlinearily. Regarding prompt tuning, we use $l=6$ prompts and $k=16$ experts by default, and the tunable prompt is applied to all layers of the main modality encoder. Vanilla prompt tuning [20] with $l^{\prime}=4$ is used to tune the complementary modality.
Training Details. All images are resized and cropped to the size of $224\times 224$. All models are trained for 15 epoches, using the AdamW [29] optimizer with a learning rate of $4\times 10^{-4}$ for vision and $5\times 10^{-4}$ for text. We use a constant step decay scheduler that scales the learning rate at epochs 3 and 6 by a factor of $\gamma=0.4$. All models are trained on a server with a RTX4090 GPU. The mean and standard deviation of metrics, measured over 4 different experiments are reported when possible.
4.2 Datasets
MM-IMDB [1] is a multimodal movie classification dataset. It consists of 25,956 image-text pairs, where each pair includes a movie poster and a plot summary The dataset supports multi-label genre classification across a spectrum of 23 genres with a long-tail distribution.
UPMC Food-101 [41] is a comprehensive multimodal dataset designed for fine-grained recipe classification with 90,840 image-text pairs for 101 food classes.
SNLI-VE [43] is a large-scale multimodal dataset with 565,286 image-text pairs. The task for this dataset is visual entailment, which requires the model to infer whether a hypothesis matches the given premise. In our experiments, following [21], we only use the image premise. However, it is worth noting that some other works may also include the text premise.
4.3 Compared Methods
First, we compare our method with several baselines for multimodal fusion.
ImgOnly / TextOnly. Only fine-tune one encoder, and the input of the other modality is discarded.
P-ImgOnly / P-TextOnly. Only prompt-tune one encoder, and the input of the other modality is discarded.
LateConcat. This baseline involves fine-tuning both encoders, concatenating their features, and learn a linear head for classification.
P-LateConcat. Similar as LateConcat but prompt-tune each encoder instead of fine-tuning. However, the linear head is still fine-tuned.
SequentialFuse. This method first extracts features from the complementary modality and maps them to the embedding space of the main modality encoder in the same way as our mapped prompt for end-to-end training. Both encoders are fine-tuned. This is a strong baseline that can be considered as our method without MoPE, but with all parameters fine-tuned.
P-SequentialFuse. Prompt-tune each encoders in SequentialFuse.
For a fair comparison, all the prompt-tuning methods listed above use the same prompt length ($l=6$) as our method. In addition to these baselines, we also compare our methods with existing prompt-based fusion methods, including MMBT [15], Frozen [39], PromptFuse and BlindPrompt [22], and PMF [21]. Among these methods, the setting in PromptFuse [22] is the most similar to ours, as it also assumes one modality to be a black-box.
Table 1: Quantitative results on three multimodal dataset. We compare all baseline methods, existing prompt-based fusion methods, and our method with different expert numbers. Our method achieve state-of-the-art performance and parameter-efficiency. (†): Our implementation with prefix-tuning.
Method Param SNLI-VE($\uparrow$) Food-101($\uparrow$) MM-IMDB($\uparrow$)
finetuning ImgOnly 86.9M 33.34$\pm$.03 75.64$\pm$.21 39.21$\pm$.32/53.85$\pm$.21
TextOnly 109.0M 69.58$\pm$.24 86.92$\pm$.21 58.80$\pm$.31/65.37$\pm$.13
LateConcat 196.0M 72.01$\pm$.23 93.19$\pm$.21 60.43$\pm$.68/67.77$\pm$.36
SequentialFuse 197.0M 74.28$\pm$.46 93.73$\pm$.16 59.22$\pm$.38/66.34$\pm$.19
MMBT [15] 196.5M 67.58 94.10 60.80/66.10
prompt-tuning P-ImgOnly 0.1M 33.34$\pm$.01 76.65$\pm$.07 33.70$\pm$.55/50.04$\pm$.27
P-TextOnly 0.1M 64.86$\pm$.32 81.01$\pm$.50 52.19$\pm$.32/61.16$\pm$.43
P-LateConcat 1.3M 64.29$\pm$.28 90.27$\pm$.40 56.95$\pm$.37/64.23$\pm$.26
P-SequentialFuse 1.1M 65.01$\pm$.18 81.27$\pm$.22 55.57$\pm$.63/63.98$\pm$.35
P-MMBT [21] 0.9M 67.58 81.07 52.95/59.30
PromptFuse [22] <0.1M 64.53 82.21 48.59/54.49
BlindPrompt [22] <0.1M 65.54 84.56 50.18/56.46
PromptFuse(†) [22] 0.1M 64.94 82.14 50.78/60.96
PMF [21] 2.5M 71.92 91.51 58.77/64.51
Ours (ViT-b, $k=4$) 1.6M 73.47$\pm$.11 91.55$\pm$.12 62.37$\pm$.35/68.73$\pm$.24
Ours (Swin-b, $k=4$) 1.6M 73.14±.$\pm.21$ 91.54$\pm$.21 61.93$\pm$.37/68.19$\pm$.14
Ours (Swin-b, $k=16$) 2.6M 73.59$\pm$.15 92.05$\pm$.11 62.01$\pm$.21/68.24$\pm$.12
Table 2: Ablation on three types of prompt disentanglement. The results from all combinations of mapped prompt, dynamic prompt, and static prompt are presented. Our full method with all prompts achieves the best result.
Prompt SNLI-VE($\uparrow$) Food-101($\uparrow$) MM-IMDB($\uparrow$)
[…]
$[\mathbf{P}_{s},\mathbf{P}_{d},P_{m}]$ $\mathbf{73.59\!\pm\!.15}$ $\mathbf{92.05\!\pm\!.11}$ $\mathbf{62.01\!\pm\!.21}$/$\mathbf{68.24\!\pm\!.12}$
4.4 Main Results
The quantitative results of all baselines, compared methods, and our method with different expert numbers $k$ are summarized in Tab 1. The metric on SNLI-VE and Food-101 is accuracy (%), and MM-IMDB is F1-Macro and F1-Micro (%). We also list the number of trainable parameters (million) for each method.
Our method outperforms all prompt-based methods, and is competitive with fine-tuning. Specifically, when compared with the fine-tuning baselines, SequentialFuse and LateConcat, our method delivers competitive accuracy on the Food-101 dataset and superior results on the SNLI-VE and MM-IMDB datasets, while requiring as few as 0.8% of the trainable parameters.
Apart from surpassing baselines, the proposed method also outperforms existing prompt-based fusion methods, including PromptFuse [22], BlindPrompt [22], and PMF [21]. Notably, our method outperforms the current SOTA, PMF [21], by a significant margin on all datasets, while requiring $37\%$ fewer parameters. Moreover, PMF assumes both encoders share the same Transformer architecture. By contrast, our method has weaker assumptions, which could be easily extended to any Transfomer-based architecture, such as ViT [4] and Swin [28].
4.5 Qualitative Result of MoPE Routing
During end-to-end tuning, prompt experts in MoPE spontaneously specialize in different concepts. Fig. 3 presents two examples of this specialization. Specifically, we visualize the expert with the highest score as the routed expert for each instance pair. In the provided examples, we observe that expert-4 specialize in children-related concepts, while expert-12 focuses on crowds. These examples also demonstrate the interpretability of our MoPE-based prompting, as opposed to black-box conditional prompt such as CoOp [50] and CoCoOp [49]. We encourage readers for a more comprehensive visualizations in Section 0.C.2.
5 Analysis and Discussion
In this section, we systematically analyze the effects of our proposed design through ablation studies. We show that our MoPE-based fusion method is more adaptive and scalable for multimodal fusion compared to existing approaches.
Our three-types of disentangled prompts are collaborative. We first ablate each type of prompt in our prompt disentanglement, and the results are summarized in Tab. 2. Our full method achieves the best performance, indicating that the prompts collaborate effectively with each other. Specifically, the dynamic prompt is essential for multimodal fusion as it learns the interplay between different modalities. This is in contrast with previous methods [22, 39, 45], where each modality do not explicitly interact. Adding the dynamic prompt induces a significant gain of 13.5%, 13.26, 8.4%/6.5% on all datasets. Without the mapped prompt, the fine-grained information from complimentary modality is lost, leading to a significant drop in all datasets. The static prompt is also necessary, serving as a special expert that is always routed. Adding this special expert allows the other experts to focus more on capturing instance-level features, resulting in a slight performance gain.
Figure 4: More experts v.s. longer prompts. We compare increasing the number of experts, $k$, versus lengthen prompt, $l$. Expert-scaling consistently outperforms length-scaling, exhibiting a linear growth trend and do not deteriorate even when $k\to\infty$. Conversely, excessive prompt length has a negative impact on accuracy and speed.
Expert-scaling is more expressive than length-scaling. To study the expressiveness of MoPE, we compare expert-scaling with length-scaling. Our starting point for MoPE is $l=6$ prompts and $k=2$ experts (when $k\leq 1$, MoPE degenerate into vanilla prompt), which account for $(2+1)\times 6+1=19$ tunable prompts in total per-layer. We increase the number of $k$, and at each step we also report the result of simply increasing $l$ in vanilla prompt, up to the same total number of prompts. The results are summarised in Fig. 4.
As illustrated in the figure, adding the MoPE design with as few as $k=2$ experts boosts performance, and increasing the number of experts leads to a monotonic performance improvement. Although our method may also saturate when the number of experts becomes excessive, it does not face the over-length deterioration issue observed in vanilla prompt tuning. In contrast, merely augmenting the number of prompts $l$ in vanilla prompt tuning does not lead to a linear performance improvement and is prone to performance deterioration. This observation is consistent with previous findings [11, 44, 14, 20]. This makes the prompt length $l$ a hyperparameter that is difficult to tune manually. In addition to accuracy considerations, longer prompts in vanilla prompt tuning can exacerbate computational overhead due to the quadratic time complexity of self-attention. This is why we cannot experiment with $l>199$ for the other two methods. On the other hand, MoPE scales by conditioning the dynamic prompt on more experts while maintaining a fixed length of learnable prompts during self-attention. As a result, our method maintains a nearly constant time complexity.
Table 3: Result of dense routing v.s. sparse routing. Dense routing achieves slightly better results on all three datasets. Routing SNLI-VE $\uparrow$ Food-101 $\uparrow$ MM-IMDB $\uparrow$ Dense 73.55 91.74 62.01/68.25 Sparse 71.15 90.89 60.28/66.01 Table 4: Different way of generating routing embedding. Frozen routing embedding with orthogonal initialization is as good as learned one. Routing Embed SNLI-VE $\uparrow$ Food-101 $\uparrow$ MM-IMDB $\uparrow$ Frozen 73.55 91.74 62.01/68.25 Learned 73.13 91.20 61.64/67.97
[Uncaptioned image] Figure 5: Scaling performance with increased training data. We show the performance of our method and representative methods as we progressively increase the amount of training data, or “#shots”. Our method outperform other prompt-tuning methods at all data scales.
MoPE scales better when more training data is available. Previous prompting methods have been shown to not scale well with respect to increased training data [7, 22, 39, 44]. We study how the proposed MoPE method performs on different data scales to assess its scalability. Specifically, we sub-sample the training-set with different shots to represent a range of low-shot to high-shot learning scenarios. We then train our method and other prompt-tuning and fine-tuning methods with the same subsampled data. The mean accuracy and F1-Micro score on SNLI-VE and MM-IMDB, respectively, are reported in Fig. 5.
As the figure shows, our MoPE-based method is the most scalable, achieving superior results across all data scales. Specifically, we consistently match the results of the fine-tuning method, SequentialFuse, on the SNLI-VE dataset, and we achieve superior results on the MM-IMDB dataset. By contrast, the method based on vanilla-prompt does not scale well with respect to increased training data, and there is a consistent performance gap between ours and the compared method. This performance gap becomes more significant on larger datasets, such as when training with $10^{5}$ shots or using the full training set.
Refer to caption Figure 6: Routing determinism. We show the entropy of routing score during training. Dense routing are more deterministic as optimization progress. Better viewed with color.
Figure 7: Effect of the importance loss. We visualize how the importance (z-axis) of all experts (x-axis) in the last Transformer layer changes during training (y-axis). (a) Without importance loss, only a few experts are used throughout training (b) The importance loss ensures balanced utilization of all experts.
Dense routing with frozen routing embedding achieves the best result. Previous MoE with FFN acting as experts usually employ a sparse gating function that only select one or a few experts per token, while we use a dense router (Eqn. 4) that weigh all experts. To compare the effect, we follow  [36] to implement a sparse router by adding a Top-1 gate: $\mathbf{r}^{\prime}=\operatorname{TOP}_{1}(\mathbf{r})$.
The results of using sparse routing versus dense routing are reported in Tab. 5. Our experiments indicate that using the dense router yields marginally better results.
To further understand how dense routing benefits the model, we evaluated the degree of uncertainty in routing by calculating the empirical entropy of the routing score $\mathcal{H}(\mathbf{r})=-\sum_{i=1}^{n}r_{i}log_{2}(r_{i})$, averaged across training batches and all layers. In essence, high entropy indicates that routing behaves more randomly. The relationship between entropy and optimization steps is illustrated in Fig. 6. As demonstrated in the figure, dense routing is more deterministic.
In addition to routing sparsity, we also investigate the effect of using frozen routing embeddings. As Tab. 5 shows, fixed routing embedding could achieve the same result as the trained one.
Importance loss aids expert specialization. The importance loss is crucial for avoiding degenerate routing solutions and, consequently, aids expert specialization. In Fig. 7, we visualize how the importance of each expert (i.e., average routing score) changes when training on the SNLI-VE dataset. Without the importance loss, routing adheres to its initial state, resulting in a skewed distribution where a few experts dominate. The importance loss allows other experts to be used by penalizing highly unbalanced expert importance.
Architecture Tuning MM-IMDB ($\uparrow$)↑ )
BoW* Fine-tune 48.20/57.50
Bert [3] Frozen 58.86/66.13
GPT-2 [35] Frozen 34.03/50.84
Longformer [3] Bitfit [46] 60.34/67.27
Table 5: Our fusing method is highly modular. We treat complementary modality as black-box, allowing different models to be plug-in for fusion. (*): Bag-of-word initialized with Bert word embeddings.
Our fusion method are highly modular. Our method abstracts the complimentary modality as a black-box, allowing high modality of both modalities. In particular, our method allow arbitrary models to be seamless plug-in for multimodal fusion. We give some results in Tab. 5.
How could MoPE be more expressive & Limitations. We have demonstrated that the MoPE-based method is more expressive for multimodal fusion. Compared with existing prompt-based methods, our approach scales better with data and trainable parameters. This increased expressiveness comes from two factors. First, the prompts used in previous methods are not adaptive, meaning that a single prompt is used for all instances. Finding a instance-wise optimal prompt could lead to better results. Secondly, we scale the prompt by using more experts, since a long prompt can be difficult to optimize [32] for optimal performance. By doing so, we shift the difficulty from learning a universal long prompt to learning a parameterized router with multiple short prompts. As a result, our empirical expressiveness is closer to the theoretical upper bound of prompt-tuning.
Despite improving empirical expressiveness, our prompt still functions in the same way as vanilla prompts during attention, resulting in same theoretical upper-bound of expressiveness. This is also reflected by the small gap between ours and fine-tuning in large datasets such as SNLI-VE and UPMC Food-101. Future studies could focus on designing a more expressive prompting technique.
6 Conclusion
In this paper, we enhance the scalability and adaptivity of prompts used in multimodal fusion. Our method involves disentangling the prompt into static and dynamic components to enable instance-adaptive prompt learning. To circumvent performance deterioration due to overly long prompts, we introduce the mixture-of-prompt-expert technique, which improves the expressiveness of prompt-tuning. Extensive experiments demonstrate that our method is highly parameter-efficient and scales better with dataset size and prompt length. We have also shown that our method is interpretable and highly modular.
Appendix 0.A Additional Experiment Details
In this section, we provide additional details of our implementation and training.
Prompt Vector. Our prompt implementation closely follows [11]. Specifically, for static prompts and prompt experts, we use uniform initialization $\mathcal{U}\sim(-\eta,\eta)$, where $\eta$ is calculated according to the embedding dimension and patch size of the Transformer [11]. Dropout with $p=0.1$ is applied to all prompts. However, we do not use the reparameterization trick for prompts [20], as the gradients of our dynamic prompt and mapped prompt are already rectified by MLPs. For Transformer architectures that employ a window attention mechanism, we duplicate the same prompt and prepend it to all windows for attention calculation, following the approach in  [11].
Mapper. The main objective of the mapper is to map representations from dimension $y$ to the embedding dimension of modality $x$. Generally speaking, the mapper is implemented as a two-block MLP, which shares similarities with previous work [21]. In our experiments, we set the bottleneck dimension as half of the dimension of the complementary representation, i.e., $d_{bot}=\lceil d_{y}/2\rceil$. Then, we apply a batch normalization layer and a GeLU activation to obtain a bottleneck feature $\psi_{bot}\in\mathbb{R}{d_{bot}}$.
After obtaining this bottleneck feature, we apply another MLP to project it to the embedding dimension of $d_{x}$. However, some Transformer architectures use inconsistent $d_{x}$ in their internal layers. For example, in Swin-b[28], the embedding dimension changes from $[128,256,512,1024]$, making it challenging to fit a single MLP. To circumvent this, we instead learn four separate linear projections, each used to project the shared bottleneck feature to different embedding dimensions. As a result, there is one single down-sampling MLP and multiple up-sampling MLPs. To ensure fairness, we also use a similar design in plain Transformers such as ViT[4] and BERT [3].
MoPE. As discussed in the main body of this paper, we learn a per-layer linear projection to obtain the cross-model embedding. Here, we would like to further clarify that this projection weight is not shared with the one used in the mapper. In our experiments, we use $d_{c}=8$ and $d_{i}=2$, resulting in $d_{r}=10$. For the importance loss, we scale it by a factor of $0.01$ and apply it together with the task-specific losses.
Appendix 0.B Extended Analysis and Discussion
0.B.1 Analysis on Adaptivity of vanilla prompt and MoPE
In the main body of the paper, we have empirically demonstrated that the dynamic prompt generated by MoPE exhibits greater adaptivity compared to global-shared prompt tuning. In this section, we aim to further characterize and quantify the adaptiveness of different prompt techniques. Through this analysis, we will also provide intuitive insights into the significance of expert specialization in MoPE-based fusion.
Following the approach adopted in previous papers [42, 32], we will focus our analysis on the case of a single prompt within a single Transformer layer. Since our method is designed to generate an effective prompt while leaving the attention calculation the same as previous prompt tuning, our MoPE will only affect the forward behavior up to the calculation of attention map. Therefore, we will concentrate on analyzing how different prompts influence the attention pattern of a pretrained Transformer layer.
Let us define the attention map produced by the self-attention operation as $\mathbf{A}(\mathbf{X},\mathbf{P})$, which is a function of the input $\mathbf{X}$ and the prompt $\mathbf{P}$. The objective is to find a prompt $\mathbf{P}$ that enables the attention map $\mathbf{A}(\mathbf{X},\mathbf{P})$ to closely match the desired target attention pattern for each input instance.
Theorem 0.B.1 (Limited Adaptivity of Global-shared Prompt)
Let $\mathcal{X}$ be the input space, and $\mathcal{A}$ be the space of attention matrices. For any input $\mathbf{x}\in\mathcal{X}$, let $\mathcal{A}^{*}(\mathbf{x})\subseteq\mathcal{A}$ denote the set of optimal attention patterns that minimize the downstream task loss. Define the attention mapping induced by a prompt $\mathbf{P}$ as $\mathcal{A}(\mathbf{x},\mathbf{P})\subseteq\mathcal{A}$. Then, for vanilla prompting with a single shared prompt $\mathbf{P}$, there exists no $\mathbf{P}\in\mathcal{P}$ such that $\mathcal{A}^{*}(\mathbf{x})\subseteq\mathcal{A}(\mathbf{x},\mathbf{P})$ for all $\mathbf{x}\in\mathcal{X}$, where $\mathcal{P}$ is the prompt space.
Proof
Let $\mathbf{x}_{1},\mathbf{x}_{2}\in\mathcal{X}$ be two distinct input instances with disjoint optimal attention sets, i.e., $\mathcal{A}^{*}(\mathbf{x}_{1})\cap\mathcal{A}^{*}(\mathbf{x}_{2})=\emptyset$. Define the attention discrepancy for an instance $\mathbf{x}$ and prompt $\mathbf{P}$ as:
$\Delta(\mathbf{x},\mathbf{P})=\inf_{\mathbf{A}\in\mathcal{A}(\mathbf{x},\mathbf{P})}\|\mathbf{A}-\mathcal{A}^{*}(\mathbf{x})\|_{\mathcal{A}}$ (6)
where $\|\cdot\|_{\mathcal{A}}$ is a suitable distance metric on $\mathcal{A}$.
Let $\mathbf{P}_{1}^{*}$ and $\mathbf{P}_{2}^{*}$ be the locally optimal prompts for instances $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$, respectively, i.e.,
$\mathbf{P}_{1}^{*}=\arg\min_{\mathbf{P}\in\mathcal{P}}\Delta(\mathbf{x}_{1},\mathbf{P})$ (7)
$\mathbf{P}_{2}^{*}=\arg\min_{\mathbf{P}\in\mathcal{P}}\Delta(\mathbf{x}_{2},\mathbf{P})$ (8)
For vanilla prompting with a shared prompt $\mathbf{P}$, the globally optimal shared prompt $\mathbf{P}_{\text{shared}}^{*}$ minimizes the accumulated attention discrepancy:
$\mathbf{P}_{\text{shared}}^{*}=\arg\min_{\mathbf{P}\in\mathcal{P}}\Delta(\mathbf{x}_{1},\mathbf{P})+\Delta(\mathbf{x}_{2},\mathbf{P})$ (9)
Due to the disjointness of optimal attention sets for each instance, the accumulated attention discrepancy for the globally optimal shared prompt $\mathbf{P}_{\text{shared}}^{*}$ is lower-bounded by the sum of the locally optimal attention discrepancies for $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$:
$\Delta(\mathbf{x}_{1},\mathbf{P}_{\text{shared}}^{*})+\Delta(\mathbf{x}_{2},\mathbf{P}_{\text{shared}}^{*})\geq\Delta(\mathbf{x}_{1},\mathbf{P}_{1}^{*})+\Delta(\mathbf{x}_{2},\mathbf{P}_{2}^{*})$ (10)
Furthermore, by the limited expressiveness of prompt-tuning [32], we have:
$\Delta(\mathbf{x}_{2},\mathbf{P}_{2}^{*})\geq 0;\\ \Delta(\mathbf{x}_{1},\mathbf{P}_{1}^{*})\geq 0$ (11)
Hence, the following relationship holds:
$\Delta(\mathbf{x}_{1},\mathbf{P}_{\text{shared}}^{*})+\Delta(\mathbf{x}_{2},\mathbf{P}_{\text{shared}}^{*})\geq\Delta(\mathbf{x}_{1},\mathbf{P}_{1}^{*})+\Delta(\mathbf{x}_{2},\mathbf{P}_{2}^{*})\geq 0$ (12)
Theorem 0.B.2 (Improved Adaptivity of MoPE)
Let =$\mathcal{E}=\{\mathcal{E}_{1},\mathcal{E}_{2},\ldots,\mathcal{E}_{k}\}$ be a set of $k$ expert attention mappings, where each $\mathcal{E}_{i}:\mathcal{X}\rightarrow\mathcal{A}$ maps an input $\mathbf{x}\in\mathcal{X}$ to a set of attention patterns $\mathcal{E}_{i}(\mathbf{x})\subseteq\mathcal{A}$. Define the induced attention mapping of MoPE as:
[…]
where $\mathbf{r}(\mathbf{x})=[r_{1}(\mathbf{x}),\ldots,r_{k}(\mathbf{x})]$ are the routing weights for input $\mathbf{x}$.
Let $\mathcal{X}^{\prime}\subseteq\mathcal{X}$ be a set of instances. If the convex hull of the expert attention mappings, denoted as:
[…]
$\sum_{\mathbf{x}\in\mathcal{X}^{\prime}}\Delta(\mathbf{x},\mathcal{E},\mathbf{r}^{*})=\sum_{\mathbf{x}\in\mathcal{X}^{\prime}}\inf_{\mathbf{A}\in\mathcal{A}^{*}(\mathbf{x})}\|\mathbf{A}-\mathbf{A}^{*}\|_{\mathcal{A}}$ (15)
Proof
When there are more experts than instances, i.e. $|\mathcal{X}^{\prime}|\leq k$ , the proof is trivial. This is because the the optimal prompt for each instance could simply be “stored” in one or a few experts.
When cardinality of the instances is greater than the number of experts, i.e., $|\mathcal{X}^{\prime}|>k$. Let $\mathcal{X}^{\prime}\subseteq\mathcal{X}$ be a set of instances with $|\mathcal{X}^{\prime}|>k$, and assume that the convex hull of the expert attention mappings, $\text{conv}(\mathcal{E})$, contains an optimal attention pattern for each instance in $\mathcal{X}^{\prime}$, i.e., $\mathcal{A}^{*}(\mathbf{x})\subseteq\text{conv}(\mathcal{E}),\forall\mathbf{x}\in\mathcal{X}^{\prime}$. We call this premise as the specialized experts condition.
Since the routing weights $\mathbf{r}(\mathbf{x})$ are convex combinations of the expert attention patterns, the induced attention mapping of MoPE, $\mathcal{A}_{\text{MoPE}}(\mathbf{x},\mathcal{E},\mathbf{r})$, is equal to $\text{conv}(\mathcal{E})$:
$\mathcal{A}_{\text{MoPE}}(\mathbf{x},\mathcal{E},\mathbf{r})=\text{conv}(\mathcal{E}),\quad\forall\mathbf{x}\in\mathcal{X}^{\prime}$ (16)
Therefore, for each instance $\mathbf{x}\in\mathcal{X}^{\prime}$, there exists an attention pattern $\mathbf{A}^{*}\in\mathcal{A}^{*}(\mathbf{x})$ that is also contained in $\mathcal{A}_{\text{MoPE}}(\mathbf{x},\mathcal{E},\mathbf{r}^{*})$ for some routing score $\mathbf{r}^{*}$, i.e.,
$\exists\mathbf{A}^{*}\in\mathcal{A}^{*}(\mathbf{x})\cap\mathcal{A}_{\text{MoPE}}(\mathbf{x},\mathcal{E},\mathbf{r}^{*}),\quad\forall\mathbf{x}\in\mathcal{X}^{\prime}$ (17)
This implies that the attention discrepancy for each instance $\mathbf{x}\in\mathcal{X}^{\prime}$ under the routing score $\mathbf{r}^{*}$ is equal to the optimal instance-wise attention discrepancy:
$\Delta(\mathbf{x},\mathcal{E},\mathbf{r}^{*})=\inf_{\mathbf{A}\in\mathcal{A}_{\text{MoPE}}(\mathbf{x},\mathcal{E},\mathbf{r}^{*})}\|\mathbf{A}-\mathbf{A}^{*}\|_{\mathcal{A}}=\inf_{\mathbf{A}\in\mathcal{A}^{*}(\mathbf{x})}\|\mathbf{A}-\mathbf{A}^{*}\|_{\mathcal{A}},\quad\forall\mathbf{x}\in\mathcal{X}^{\prime}$ (18)
Therefore, the accumulated attention discrepancy for MoPE under the routing score $\mathbf{r}^{*}$ across instances in $\mathcal{X}^{\prime}$ is equal to the sum of the optimal instance-wise attention discrepancies:
$\sum_{\mathbf{x}\in\mathcal{X}^{\prime}}\Delta(\mathbf{x},\mathcal{E},\mathbf{r}^{*})=\sum_{\mathbf{x}\in\mathcal{X}^{\prime}}\inf_{\mathbf{A}\in\mathcal{A}^{*}(\mathbf{x})}\|\mathbf{A}-\mathbf{A}^{*}\|_{\mathcal{A}}$ (19)
which is equivalent as:
$\sum_{\mathbf{x}\in\mathcal{X}^{\prime}}\Delta(\mathbf{x},\mathcal{E},\mathbf{r}^{*})=\sum_{\mathbf{x}\in\mathcal{X}^{\prime}}\Delta(\mathbf{x},\mathbf{P}_{\mathbf{x}}^{*})\geq 0$ (20)
Thus, when the number of instances exceeds the number of experts, if the convex hull of the expert attention mappings contains an optimal attention pattern for each instance, then there exists a routing score $\mathbf{r}^{*}$ that allows MoPE to achieve an accumulated attention discrepancy equal to the sum of the optimal instance-wise attention discrepancies across those instances.
Putting it together, the Theorem 0.B.1 state that the global-shared prompt (i.e. vanilla prompt) could not achieve the best result, when the input instances require different attention patterns to perform well. Equation 15 state that it is possible for MoPE to achieve the best result on all of the input instances, conditioned on the appropriate specialization of experts.
0.B.2 Discussion: How to Choose the Main Modality?
In the proposed method, we use a sequential pipeline to fuse different modalities. In this pipeline, the input of the complementary modality $y$ is first encoded into a representation $\psi_{y}$, which is then used to guide the prompting of modality $x$. This design raises an interesting question: “How to choose the complementary and main modality?”
For tasks that do not require a dense representation, such as classification, our experience is that either modality can be used as the main modality, yielding similar results. In our experiments, we utilize vision as the main modality due to empirically better results, but this does not necessarily mean that the text encoder cannot be the main modality. We have also tested using text as the main modality, and the results are summarized in Tab. 6.
Table 6: Result of Language Model (LM) as main modality.
Method SNLI-VE Food-101 MM-IMDB
Ours(LM) 71.10$\pm$.12 88.01$\pm$.12 58.38$\pm$.11/65.81$\pm$.23
Ours(VM) 73.59$\pm$.15 92.05$\pm$.11 62.01$\pm$.21/68.24$\pm$.12
As the table shows, using BERT as the main encoder results in a decline in performance. Our postulation is that for these three tasks, the text input contains a significant amount of noisy and false positive data. For instance, in the UPMC-Food 101 dataset [41], the text data is derived using a spider, which includes many irrelevant hypertext and website titles. Therefore, treating text as the main modality can lead to overfitting and biased predictions.
For tasks that require a modality-specific dense representation for decoding, such as text generation and segmentation, the main modality must be the one that provides such output. For example, we use the BERT encoder to guide Swin for segmentation, and we use Swin to guide BERT for visual question answering (detail in Appendix 0.C).
0.B.3 Our MoPE-based method yields a better multimodal representation.
We visualize the multimodal representation generated by different methods using t-SNE. For LateConcat, this would be the concatenated feature from both modalities. For the other methods, we visualize the [CLS] token of the main modality. For ease of coloring, we only show the first 20 classes in the test set of the UPMC Food-101 [41] dataset. The results are presented in Fig. 8. As the results show, the representation generated by our method is the most separable.
Figure 8: t-SNE visualization of multimodal representation generated by different methods. The representation generated by our method is the most separable. Better viewed with color.
Appendix 0.C Boarder Applications
In this section, we explore the broader applications of our MoPE-based multimodal fusion method beyond classification tasks by providing additional experiments on dense prediction tasks.
0.C.1 Referring Image Segmentation
The proposed fusion method focuses on utilizing prompts to adapt the Transformer encoders to derive a multimodal representation. Based on this representation, we can tackle tasks beyond classification, such as segmentation. In this section, we briefly introduce how our method can be applied to this scenario and provide experimental results.
Datasets and Task. RefCOCO [13] and RefCOCO+[13] are the commonly used datasets for referring image segmentation (RIS), both built upon the MS-COCO[24] dataset. RefCOCO contains 142,209 referring expressions for 50,000 objects across 19,994 images, while RefCOCO+ includes 141,564 expressions for 49,856 objects in 19,992 images. The task involves predicting a segmentation mask from the input image according to a given referring expression in natural language.
Architecture and Training Details. We utilize ImageNet-pretrained Swin-B-384 as our vision encoder, and we use BERT as the text encoder. For the decoder head, we adopt the standard SETR-PUP [48] head. Following [11], the head is fine-tuned, while we use the MoPE-based tuning method to adapt the encoders. The pyramid feature maps from the 4 Swin layers are fed into the segmentation head for end-to-end training. The optimization uses the same hyperparameters as in our main paper.
Compared Models. To the best of our knowledge, there is no existing prompt-based multimodal fusion method that is compatible with dense visual tasks. Notably, the previous state-of-the-art method, PMF [21], assumes that each encoder uses the same Transformer architecture, and they did not open-source their training code. Furthermore, their prompt tuning is only applied to the deep layers, while SETR-PUP [48] requires pyramid features from internal layers. Therefore, we find it challenging to adapt PMF for dense tasks such as segmentation. Given these challenges, we primarily compare our method with our baselines, SequentialFuse and p-SequentialFuse.
Quantitative Comparison with Baselines. The quantitative results, measured as mean Intersection over Union (mIoU) on the validation and test sets of RefCOCO and RefCOCO+, are reported in Tab. 7. As shown in the table, our MoPE-based method achieves the best performance. We outperform the prompt-based method, p-SequentialFuse, by a significant margin. Additionally, our method performs better than the fine-tuning method. We observe that the fine-tuning method tends to overfit in this experimental setting.
Table 7: Segmentation results on RefCOCO and RefCOCO+. We report the total number of parameters in million (including the trainable segmentation head), and metrics (mIOU) on RefCOCO and RefCOCO+.
Method Param RefCOCO($\uparrow$) RefCOCO+($\uparrow$)
val testA testB val testA testB
SequentialFuse 231.0M 53.48 55.76 52.03 40.22 42.2 37.91
P-SequentialFuse 35.1M 47.69 46.23 45.81 30.66 31.48 28.79
Ours 35.5M 58.40 60.03 53.23 43.80 46.12 38.88
Visualization of segmentation result. We visualize the segmentation results in Figures 9 and 10. As the results show, our MoPE-based segmentation method can better locate the object according to the referring image segmentation task. In contrast, the method based on global-shared prompts, p-SequentialFuse, may struggle to understand the text instructions and provide incorrect results.
Figure 9: Quantitative result of refering image segmentation. Better viewed with color.
Figure 10: Quantitative result of refering image segmentation (cont). Better viewed with color.
0.C.2 Additional Examples of Expert Routing on VQAv2
Our MoPE is designed to scale up the expressiveness of vanilla prompts, and expert specialization is a critical condition to achieve superior results as well as interpretability. To further demonstrate expert specialization, we train our MoPE-based method on an even larger dataset, VQAv2 [8]. This dataset contains 265,016 images and paired questions, covering a wide array of visual and textual concepts. We train our model on this dataset and visualize the routing results in Figures 11, 12, 13, and 14. We observe clear expert specialization in these examples, where different experts capture different concepts.
Figure 11: Additional example of expert routing - 1. We show routing result of expert-4 on VQAv2 dataset, which specialize in animals. Note: Sorted by routing score, images and texts may not in same order.
Figure 12: Additional example of expert routing - 2. We show routing result of expert-6 on VQAv2 dataset, which specialize in lighting conditions. Note: Sorted by routing score, images and texts may not in same order.
Figure 13: Additional example of expert routing - 3. We show routing result of expert-8 on VQAv2 dataset, which specialize in toilet-related concepts and hairstyles. Note: Sorted by routing score, images and texts may not in same order.
Figure 14: Additional example of expert routing - 4. We show routing result of expert-10 on VQAv2 dataset, which specialize in weather conditions. Note: Sorted by routing score, images and texts may not in same order.

### v2

@@
-
+MoPE: Parameter-Efficient and Scalable Multimodal Fusion via Mixture of Prompt Experts
 Abstract
-Prompt-tuning has demonstrated parameter-efficiency in fusing unimodal foundation models for multimodal tasks.
-However, its limited adaptivity and expressiveness lead to suboptimal performance when compared with other tuning methods.
-In this paper, we address this issue by disentangling the vanilla prompts to adaptively capture dataset-level and instance-level features.
-Building upon this disentanglement, we introduce the mixture of prompt experts (MoPE) technique to enhance expressiveness.
+Despite the demonstrated parameter efficiency of prompt-based multimodal fusion methods, their limited adaptivity and expressiveness often result in suboptimal performance compared to other tuning approaches.
+In this paper, we address these limitations by decomposing the vanilla prompts to adaptively capture instance-level features.
+Building upon this decomposition, we introduce the mixture of prompt experts (MoPE) technique to enhance the expressiveness of prompt tuning.
 MoPE leverages multimodal pairing priors to route the most effective prompt on a per-instance basis.
-Compared to vanilla prompting, our MoPE-based conditional prompting exhibits greater expressiveness for multimodal fusion, scaling better with the training data and the overall number of trainable parameters.
-We also study a regularization term for expert routing, leading to emergent expert specialization, where different experts focus on different concepts, enabling interpretable soft prompting.
-Extensive experiments across three multimodal datasets demonstrate that our method achieves state-of-the-art results, matching or even surpassing the performance of fine-tuning, while requiring only 0.8% of the trainable parameters.
-Code will be released: https://github.com/songrise/MoPE.
-Keywords:
-Multimodal Fusion Prompt Tuning Mixture of Experts
-1 Introduction
-Figure 1: High-level motivation of MoPE-based conditional prompting.
-a) vanilla prompt tuning learn a global shared prompt for all instances, which may not be optimal for each instance; b) MoPE-based conditional prompt tuning optimize multiple specialized prompt experts and is instance-wise adaptative.
-Empowered with billion-scale data and highly scalable model architectures, recent unimodal pre-trained models [35, 28, 3, 4, 17], also known as foundation models, have demonstrated their transferable power to various downstream tasks [26, 30, 51].
-However, transferring multimodal foundation models for downstream applications is less flexible.
-Recent explorations, such as CLIP [34], employ two-tower designs that entangle the two encoders, meaning that replacing either one would necessitate expensive paired pre-training from scratch.
-This limitation hinders the broader application that would benefit from a specific unimodal architecture, such as longer context windows.
-To democratize foundation models, a compelling question arises: How can we efficiently combine separately pretrained unimodal foundation models for downstream multimodal tasks?
-Recently, prompt tuning [20, 11] has emerged as a novel paradigm for parameter-efficient transfer learning.
-While originally proposed for low-shot adaptation to tasks with the same modality as they were pretrained on [20, 11, 50], it has also proven effective for multimodal fusion.
-Typically, this is achieved by abstracting the representation from one modality into one or a few prompts, which are fed into the frozen Transformer of another modality for fusion [39, 22, 21, 45].
-Despite the effectiveness of prompt tuning in inter-modality transfer learning, directly adopting it for multimodal fusion could yield suboptimal results, particularly due to its limited adaptivity and scalability.
-Notably, while prompt tuning generally performs well in low-data regimes, it can be less effective when applied to full-shot training on the entire dataset with a challenging objective [7, 22, 39, 44, 19].
+Compared to vanilla prompting, our MoPE-based fusion method exhibits greater expressiveness, scaling more effectively with the training data and the overall number of trainable parameters.
+We also investigate regularization terms for expert routing, which lead to emergent expert specialization during training, paving the way for interpretable soft prompting.
+Extensive experiments across six multimodal datasets spanning four modalities demonstrate that our method achieves state-of-the-art results for prompt fusion, matching or even surpassing the performance of fine-tuning while requiring only 0.8% of the trainable parameters.
+Introduction
+Unimodal pre-trained models like BERT (Radford et al.
+2019) and Swin Transformer (Liu et al.
+2021b) have excelled in transferring capabilities to a wide range of tasks.
+By comparison, extending the pretraining-finetuning paradigm to multimodal applications involves additional complexities, particularly regarding flexibility and cost.
+First, the modality-paired pretraining can be inflexible due to its dependence on specific configurations.
+This can limit flexibility when there is a scarcity of certain modality pairs, or when new architectures emerge, necessitating resource-intensive retraining.
+In contrast, unimodal pretrained models are more readily available and offer greater flexibility, as they can be independently updated and combined for multimodal tasks.
+Second, as these foundation models scale in size, fine-tuning them becomes increasingly expensive, which further restricts their application to downstream tasks.
+To democratize foundation models, a compelling question arises: How can we efficiently combine separately pretrained unimodal models for multimodal tasks?
+Figure 1: High-level motivation of MoPE-based multimodal fusion.
+(a) Vanilla prompt tuning learns a globally shared long prompt for all instances, which may not be optimal for each instance.
+(b) MoPE decomposes the global prompt into multiple short and specialized prompt experts to improve its adaptivity and expressiveness.
+While originally proposed for transfer learning (Li and Liang 2021; Jia et al.
+2022; Zhou et al.
+2022b), recent research has revealed that prompt tuning could also be a parameter-efficient paradigm for multimodal fusion (i.e., prompt fusion).
+Typically, this is achieved by representing one modality with several prompts, which are fed into the frozen Transformer of another modality to achieve fusion (Tsimpoukelli et al.
+2021; Liang, Zhao, and Schütze 2022; Li et al.
+2023; Yang et al.
+2022b).
+However, directly using prompts for fusion could yield suboptimal results, particularly due to their limited adaptivity and scalability.
+Notably, while prompt tuning generally performs well in low-data regimes, it can be less effective when applied to full-shot training on the entire dataset with a challenging objective (Gao et al.
+2022; Liang, Zhao, and Schütze 2022; Tsimpoukelli et al.
+2021; Yang et al.
+2022a; Lester, Al-Rfou, and Constant 2021).
 This reduced efficacy could be attributed to at least two factors.
-First, the vanilla prompt tuning used in previous prompt-based fusion methods [39, 21, 22] is not designed for multimodal tasks.
-These methods simply optimize a globally shared prompt for all instances [11, 20, 39], neglecting the multimodal interplay and potentially failing to capture the locally optimal representation for each input instance.
-Second, the relatively small number of trainable parameters (compared with fine-tuning) can lead to underfitting in multimodal datasets with a long-tail distribution and complex cross-modal mapping [27, 32, 42].
-To address these challenges, increasing the number of learnable prompts, known as “length-scaling,” appears to be a plausible solution.
-Despite its apparent simplicity, a lot of research indicates that the performance gains from length-scaling quickly reach saturation in both unimodal [7, 19] and multimodal cases [22, 39, 44].
-Furthermore, over-length prompts may even lead to worse results [11, 44, 14, 20, 16].
-Recent theoretical analyses have substantiated these empirical observations, highlighting the limited expressiveness of prompt tuning [42, 32].
-A direct implication of their analysis is that simply increasing prompt length is insufficient to address the issues of adaptivity and expressiveness.
-Based on the above observations, we aim to enhance the expressiveness of the vanilla prompt used in multimodal fusion to improve adaptivity and scalability.
-Our motivation is to utilize the multimodal input pairing information as a prior, and to scale up the prompt in another dimension—the number of prompt experts—rather than length, as illustrated in Fig.
+First, the vanilla prompt was originally developed for unimodal learning, which optimizes globally-shared prompts for all instances (Jia et al.
+2022; Li and Liang 2021; Tsimpoukelli et al.
+2021).
+When directly applied for prompt fusion (Tsimpoukelli et al.
+2021; Li et al.
+2023; Liang, Zhao, and Schütze 2022), it fails to adaptively exploit the multimodal interplay and can lead to sub-optimal representations for each input instance.
+Second, the relatively limited representation power (compared to fine-tuning) can lead to underfitting in multimodal datasets with a long-tail distribution and complex cross-modal mapping (Liu et al.
+2021a; Petrov, Torr, and Bibi 2023; Wang et al.
+2023).
+To address these challenges, increasing the number of learnable prompts, known as “length-scaling”, appears to be a straightforward solution.
+However, numerous studies have shown that the performance gains from length-scaling quickly reach saturation in both transfer-learning (Gao et al.
+2022; Lester, Al-Rfou, and Constant 2021) and fusion settings (Liang, Zhao, and Schütze 2022; Tsimpoukelli et al.
+2021; Yang et al.
+2022a).
+Furthermore, over-length prompts may even lead to worse results (Jia et al.
+2022; Yang et al.
+2022a; Khattak et al.
+2023; Li and Liang 2021; Kim et al.
+2023).
+Recent theoretical analyses have substantiated these empirical observations, highlighting the limited expressiveness of prompt tuning (Wang et al.
+2023; Petrov, Torr, and Bibi 2023).
+Overall, evidence from both empirical and theoretical analyses reveals that simply increasing prompt length is insufficient to address the limitations of prompt fusion.
+To improve the adaptivity and expressiveness of prompt fusion, we propose decomposing the unified, long prompt into multiple short prompts with specialized responsibilities, as depicted in Fig.
 1.
-To accomplish this, we first propose a conditional prompt tuning method, which disentangles the vanilla global-shared prompt vector into static and dynamic prompts to adaptively capture global and local features.
-The cornerstone of our dynamic prompt generation is our innovative MoPE (Mixture of Prompt Experts) design.
-In this design, we fix the prompt length used in attention and scale up expressiveness by learning multiple prompt experts and a router at each layer.
-We then use the representation from another modality as a prior to route the most effective dynamic prompt for each instance, enabling adaptive fusion.
-We show that our MoPE-based fusion method alleviates the two issues in previous prompt-based fusion methods by allowing instance-wise adaptive prompting and avoiding over-length deterioration for better scalability.
-We systematically compare our MoPE-based fusion method with previous fusion methods on three multimodal datasets, UPMC Food-101 [41], SNLI-VE [43] and MM-IMDB [1].
-Not only does our method deliver state-of-the-art results on all datasets, it also has higher parameter-efficiency compared with methods based on vanilla prompting, due to our improved adaptivity and expressiveness.
-Specifically, we achieve results comparable to fine-tuning while requiring only $\sim 0.8\%$ of the trainable parameters, and we outperform previous state-of-the-art (SOTA) prompt-based methods by a significant margin.
-Additionally, we demonstrate the superior scalability of our method.
-Our analysis indicates that increasing the number of experts (i.e.
-“expert-scaling”) tends to be a better alternative than length-scaling, which avoids performance deterioration with an overly-long prompt.
-In addition to parameter scaling, our method also demonstrates better scalability with increased training data.
-Finally, as the first endeavor to use prompts as experts, we analyze how different methods of expert routing can affect expert specialization.
-Thanks to the proposed regularization term, we observe the emergence of specialized prompt experts after end-to-end training, achieving high interpretability compared to black-box conditional prompting.
-We summarize our contributions as follows:
-  • We design a conditional prompt tuning method for multimodal fusion, which disentangles the vanilla prompt as static, dynamic and mapped prompts for better adaptiveness.
+We introduce an instance-wise adaptive prompt decomposition that separates the vanilla prompt into static, dynamic, and mapped prompts to adaptively capture global and local features.
+The foundation of our dynamic prompt generation lies in the proposed MoPE (Mixture of Prompt Experts) design, where we maintain a fixed prompt length in the self-attention and enhance expressiveness by learning multiple prompt experts and a multimodal router to select the most suitable prompt for each instance.
+We systematically compare our MoPE-based fusion with previous prompt fusion methods on a total of six datasets, and demonstrate consistently superior performance with higher parameter-efficiency.
+Furthermore, we showcase the superior scalability of our method both in terms of parameter and data.
+Our analysis indicates that increasing the number of experts (i.e., “expert-scaling”) is more scalable than length-scaling, with monotonic performance gains and the avoidance of performance deterioration with an overly-long prompt.
+Our method also scales better with more training data.
+By introducing regularization terms for expert routing, we observe the emergence of specialized prompt experts after end-to-end training, resulting in high interpretability.
+Our key contributions are summarized as follows:
+  • We propose an instance-wise adaptive prompt decomposition, which decomposes the vanilla prompt into static, dynamic and mapped prompt.
   • We introduce the MoPE technique for instance-wise dynamic prompt generation, which scales up the expressiveness of prompt tuning for multimodal fusion.
-  • A combination of regularization terms is studied to aid expert specialization.
-  • Extensive experiments on UPMC Food-101, SNLI-VE and MM-IMDB datasets demonstrate state-of-the-art performance and parameter-efficiency for multimodal fusion.
-2 Related Works
+  • A combination of regularization terms is studied to aid specialization of prompt experts.
+  • Extensive experiments on six datasets spanning four modalities demonstrate state-of-the-art performance and parameter efficiency for prompt-based multimodal fusion.
+Related Works
 Prompt Tuning for Transfer Learning.
-Prompt tuning involves learning continuous token embedding as additional input to a frozen pretrained model for transfer learning.
-This technique was first popularized in NLP [20, 19], then quickly introduced to CV society [11, 2, 25, 23] and multimodal learning [12, 14, 49, 47].
-For either modality, researchers empirically report good transfer learning performance in low-shot adaptation, yet its performance is less comparable to fine-tuning when abundant training instances are available [20, 7].
-Moreover, increasing prompt length quickly reaches performance saturation, and over-length prompts might have an negative impact on performance [44, 11, 16].
-Recent theoretical analysis [32, 42] reveal that the expressiveness of vanilla prompt tuning is lower than fine-tuning and LoRA [10].
-In this work, we scale up the expressiveness of vanilla prompting by employing a Mixture of Experts (MoE)-like design.
+Prompt tuning (Li and Liang 2021; Lester, Al-Rfou, and Constant 2021) learns continuous token embeddings as additional input to a frozen pretrained model for transfer learning of Transformer-based models.
+It is widely used in various modalities (Jia et al.
+2022; Bahng et al.
+2022; Liu et al.
+2022; Lin et al.
+2023; Jiang, Liu, and Chen 2023; Khattak et al.
+2023; Zhou et al.
+2022a).
+A common observation is good transfer learning performance in low-shot scenarios, yet its performance is less comparable to fine-tuning when abundant training instances are available (Li and Liang 2021; Gao et al.
+2022).
+Moreover, increasing prompt length quickly reaches performance saturation, and over-length prompts might lead to worse results (Yang et al.
+2022a; Jia et al.
+2022; Kim et al.
+2023).
+Recent theoretical analyses (Petrov, Torr, and Bibi 2023; Wang et al.
+2023) reveal that the expressiveness of prompt tuning is lower than that of fine-tuning.
+In this work, we tackle this challenge by scaling up the expressiveness of prompts with a Mixture of Experts (MoE)-like design.
 Multimodal Fusion via Prompt.
-Apart from adapting models, prompts can also be used to fuse separately pretrained unimodal models for multimodal tasks.
-This is a different task from the commonly known “multimodal prompt tuning”, which usually refers to using prompts for transfer learning of multimodal pretrained foundation models.
-Frozen [39] first introduces a method where the visual representation is mapped as a few input tokens to query frozen language models (LM).
-PromptFuse and BlindPrompt [22] improved upon this by introducing tunable prompts to the LM for cross-model alignment.
-PICa [45] translate images into discrete text captions to prompt frozen LMs.
+Apart from model adaptation, prompts can also be used to fuse separately pretrained unimodal models for multimodal tasks.
+Frozen (Tsimpoukelli et al.
+2021) first introduced a method where the visual representation is mapped as a few input tokens to query frozen language models (LMs).
+PromptFuse and BlindPrompt (Liang, Zhao, and Schütze 2022) improved upon this by introducing tunable prompts to the LM for cross-modal alignment.
+PICa (Yang et al.
+2022b) translates images into discrete text captions to prompt frozen LMs.
 These methods treat tokens from different modalities independently and lack cross-modal interaction.
-Recently, PMF [21] introduced an interactive prompt-tuning method, but their interactivity is built on the strong assumption that both encoders are white-box and in the same Transformer architecture.
-In this work, we further investigate how to efficiently utilize cross-modal features for prompting while maintaining high modularity by treating one modality as a black box.
+Recently, PMF (Li et al.
+2023) introduced an interactive prompt fusion method for the vision and text modalities, building on the strong assumption that both encoders are white-box and have the same Transformer architecture.
+In this paper, we investigate how to adaptively interact with two or more modalities with minimal assumptions about the model architecture.
 MoE in Transformers.
-MoE is a powerful technique to scale up models, including Transformers [18, 36, 31, 6].
-The fundamental approach involves inserting MoE layers, composed of multiple feed-forward networks (FFNs) acting as experts, into the standard Transformer architecture.
-A router is learned to route each token embedding to the most suitable expert(s).
-Due to the significant computational cost associated with FFN forwarding, a sparse gating function is typically employed to limit the number of experts used per token [33, 36, 38].
-Thanks to those designs, the MoE layer can efficiently scale up model capacity while maintaining computational efficiency.
-Inspired by previous FFN-based MoE, we explore scaling up the vanilla prompting with an MoE design.
-3 Method
-Our objective is to adaptively and efficiently fuse unimodal pretrained models for solving multimodal tasks via prompt tuning.
-In this section, we first review the working mechanism and the limitation of vanilla prompt tuning in Sec.
-3.1.
-In Sec.
-3.2, we explain how we achieve conditional prompt-tuning via a sequential pipeline and disentangled prompt tuning.
-In Sec.
-3.3 we introduce the proposed MoPE, consisting of multiple prompt experts and a mutimodal router for expressiveness scaling.
-Finally, we introduce the regularization methods in Sec.
-3.4.
+MoE is a computationally efficient technique to scale up models, including Transformers (Lepikhin et al.
+2020; Riquelme et al.
+2021; Mustafa et al.
+2022; Fedus, Zoph, and Shazeer 2022).
+The fundamental approach involves inserting MoE layers, usually composed of multiple feed-forward networks (FFNs) acting as experts, into the standard Transformer architecture.
+A router is learned to route each token embedding to the most suitable expert(s) for reducing computational overhead (Qin and Eisner 2021; Riquelme et al.
+2021; Shazeer et al.
+2017).
+In this paper, we do not focus on reducing computational cost; instead, we introduce the MoE design into prompt fusion to scale up its adaptiveness and expressiveness.
+Method
 Figure 2: Architecture overview.
-We illustrate the proposed MoPE-based multimodal fusion method when applied to one frozen Transformer layer.
-a) We sequentially fuse two modalities, using modality y to guide the prompting of modality x through a three-type prompt disentanglement process.
-(b) MoPE is introduced, utilizing the multimodal pairing as a prior to instance-wisely generate the dynamic prompt; (c) Inside the multimodal router, we project the representation from each modalities.
-The concatenated feature (along channel dimension) is used to match the routing embedding paired with each experts for routing score calculation.
+a) A sequential fusion pipeline is used, using modality $\mathbb{Y}$ to guide the prompting of modality $\mathbb{X}$ through our prompt decomposition.
+(b) We introduce MoPE, which utilize the multimodal pairing as a prior to route the most suitable dynamic prompt for each instance; (c) Inside the multimodal router, we project the representation from each modalities.
+The concatenated embedding is used to match the routing embedding paired with each experts for routing score calculation.
 Better viewed with color.
-3.1 Preliminary: Vanilla Prompt Tuning
-Consider a Transformer [40] or its variants used to extract features from an embedded input sequence $\mathbf{x}^{0}\in\mathbb{R}_{s\times d_{x}}$, where $s$ is the sequence length and $d_{x}$ is the embedding dimension of the Transformer.
-Prompt-tuning [^1] In this paper we assume that soft prompts are applied to all layers, which corresponds to “VPT-deep” in CV and “prefix-tuning” in NLP.
-freezes all pre-trained model weights and optimize a small amount of continuous embeddings (i.e., prompts) $\mathbf{P}\in\mathbb{R}_{l\times d}$ inserted into the input, where $l$ is the number of prompts or “prompt length”.
+Preliminary: Vanilla Prompt Tuning
+Consider a Transformer (Vaswani et al.
+2017) or its variants used to extract features from an embedded input sequence $\mathbf{x}^{0}\in\mathbb{R}_{s\times d_{x}}$, where $s$ is the sequence length and $d_{x}$ is the embedding dimension.
+Prompt-tuning freezes all pre-trained parameters and optimize a small amount of continuous embeddings (i.e., prompts) $\mathbf{P}\in\mathbb{R}_{l\times d}$ concatenated to the input of each layer, where $l$ is length of the prompt.
 The input of the $i$-th layer layer $L^{i}$ could be denoted as:
@@
 where $x^{i-1}_{0}\in\mathbb{R}_{d_{x}}$ denotes the [CLS] token, $\mathbf{T}^{i-1}\in\mathbb{R}_{s\times d_{x}}$ is the token embedding from the previous layer, and [,][,][ , ] denotes the concatenation operation.
-Succinctly, prompt tuning works by biasing the pretrained attention pattern in each Transformer layers [32].
-As opposed to fine-tuning, this biasing has strictly limited expressiveness [32, 42], meaning that there are tasks that are un-learnable even with $l\to\infty$.
-The empirical expressiveness of vanilla prompting is even lower than its theoretical upper bound, due to competing optimization when $l>1$, as postulated by Petrov et al.
-[32].
-In other words, although longer prompts provide the model with more capability, finding such a long prompt becomes challenging, and a suboptimal long prompt can lead to performance degradation.
+Succinctly, prompt tuning works by biasing the pretrained attention pattern in each Transformer layers (Petrov, Torr, and Bibi 2023).
+As opposed to fine-tuning, this biasing has strictly limited expressiveness in theory (Petrov, Torr, and Bibi 2023; Wang et al.
+2023), meaning that there are tasks that are un-learnable even with $l\to\infty$.
+The empirical expressiveness of vanilla prompting is even lower than its theoretical upper bound, due to competing optimization when $l>1$ (Petrov, Torr, and Bibi 2023).
+In other words, although longer prompt gives more trainable parameters, finding such a long prompt becomes challenging, and a suboptimal long prompt can lead to performance degradation instead.
 In this paper, we do not aim to increase the theoretical upper bound of prompt-tuning; instead, we approach the problem by narrowing the gap between its empirical and theoretical expressiveness.
-3.2 Conditional Prompt Tuning for Multimodal Fusion
-Previous prompt-based multimodal fusion methods [22, 21] optimize a global prompt that is shared across all instances.
-However, this approach neglects the interplay between different modalities.
-With paired multimodal input as a prior, we aim to condition the prompting of one modality on the other for more adaptive and efficient transfer learning.
-Sequential Pipeline.
-A sequential pipeline for multimodal fusion is adopted to achieve prompt conditioning.
-Specifically, let $(x,y)$ represent a pair of multimodal inputs, and $\mathcal{E}_{y}$ be the encoder of modality $y$.
-Depending on the intrinsics of the specific task, a fusion direction could be assigned from the complementary modality to the main modality, and the proposed method will be used for tuning the main modality only.
-Without loss of generality, we consider $x$ to be the main modality in the following discussion.
-Prompt Disentanglement.
-Based on the sequential fusing pipeline, we disentangle the vanilla prompt vector $\mathbf{P}$ used in Eqn.
-1 into three types of specialized prompts $[\mathbf{P}_{s},\mathbf{P}_{d},P_{m}]$.
-The static prompt $\mathbf{P}_{s}\in\mathbb{R}_{l\times d_{x}}$ is a globally-shared prompt vector that is non-conditional to input.
-To synthesize the dynamic prompt $\mathbf{P}_{d}\in\mathbb{R}_{l\times d_{x}}$, we first extract the global-level feature from the complementary modality $\mathcal{E}_{y}(y)=\psi_{y}\in\mathbb{R}_{d_{y}}$, and utilize a MoPE module $R(\cdot)$ to synthesize the instance-wise dynamic prompt (entailed in Sec 3.3).
-Additionally, we apply a lightweight mapper $f_{m}(\cdot)$ to map the complementary feature $\psi_{y}$ into a single prompt $P_{m}\in\mathbb{R}_{d_{x}}$, which injects fine-grained cross-modal information.
-In summary, the input of layer $L^{i}$ of $\mathcal{E}_{x}$ becomes:
-$\mathbf{\hat{x}}^{i}=[x^{i-1}_{0},\underbrace{\mathbf{P}_{s},R(\psi_{y}),f_{m}(\psi_{y})}_{\text{Disentangled prompts}},\mathbf{T}^{i-1}]$ (2)
-It is important to note that our three types of prompting serve as a plug-in module to replace the vanilla prompt vector of the main modality, while treating the encoder of the complementary modality $\mathcal{E}_{y}$ as a black-box for modularity.
-The whole process is illustrated in Fig.
-2-(a).
+Instance-wise Adaptive Prompt Decomposition
+Previous prompt fusion methods (Liang, Zhao, and Schütze 2022; Yang et al.
+2022b) optimize a global prompt that is shared across all instances, neglecting the interplay between different modalities and instance-specific features.
+With paired multimodal input as a prior, we aim to instance-wisely condition the prompting of one modality on the other(s) for better adaptivity.
+To achieve this goal, we first adopt a sequential fusion pipeline.
+Let $x\in\mathbb{X},y\in\mathbb{Y}$ be a pair of multimodal inputs, and $\mathcal{E}_{\mathbb{X}}$, $\mathcal{E}_{\mathbb{Y}}$ be the encoder of each modality.
+Depending on the task intrinsic, we assign a fusion direction where the encoding of main modality $\mathbb{X}$ is conditioned on representation of complementary modality(ies) $\mathbb{Y}$.
+Building upon the sequential pipeline, we decompose the vanilla prompt vector $\mathbf{P}$ used in $\mathcal{E}_{\mathbb{X}}$ into three types of specialized prompts $[\mathbf{P}_{s},\mathbf{P}_{d},P_{m}]$.
+The static prompt $\mathbf{P}_{s}\in\mathbb{R}_{l\times d_{x}}$ is a globally-shared prompt vector that is shared for all instances.
+The dynamic prompt $\mathbf{P}_{d}\in\mathbb{R}_{l\times d_{x}}$ are adaptive to different instances.
+To synthesize it, we first encode global-level feature from the complementary modality $\psi_{y}=\mathcal{E}_{\mathbb{Y}}(y)\in\mathbb{R}_{d_{y}}$, and utilize a MoPE module $R(\cdot,\cdot)$ to generate the dynamic prompt.
+Additionally, we apply a lightweight mapper $f_{m}(\cdot)$ to map the complementary feature into a single prompt $P_{m}\in\mathbb{R}_{d_{x}}$, which injects fine-grained cross-modal information.
+In summary, the input of layer $L^{i}$ of $\mathcal{E}_{\mathbb{X}}$ becomes:
+$\mathbf{\hat{x}}^{i}=[x^{i-1}_{0},\underbrace{\mathbf{P}_{s},R(x_{0}^{i-1},\psi_{y}),f_{m}(\psi_{y})}_{\text{Decomposed prompts}},\mathbf{T}^{i-1}]$ (2)
+The whole process is illustrated in Fig. 2-(a).
 Figure 3: Interpretable prompting through MoPE routing.
@@
 In this example, expert-4 is specialized for children while expert-12 focuses on crowds.
-3.3 Mixture of Prompt Experts
-As length-scaling has been proven ineffective due to quick saturation and over-length deterioration, we propose to fix the prompt length used in attention to suppress the negative impact of longer prompts.
-To inject more learnable parameters into the framework and increase the expressiveness of the prompts, we condition the prompts on multiple prompt experts for end-to-end training.
-Prompt as Expert.
-We randomly initialize $k$ prompt experts and their associated routing embedding $\{(\mathbf{E}_{j},\mathbf{v}_{j})\}_{j=1}^{k}$ to be optimized end-to-end in each prompt-tuned Transformer layer $L^{i}$.
-For each pair, $\mathbf{E}_{i}\in\mathbb{R}_{l\times d}$ is the expert and $\mathbf{v}_{i}\in\mathbb{R}_{d_{r}}$ is the routing embedding, where $d_{r}$ is dimension of routing embedding.
-Multimodal Expert Routing.
-In multimodal tasks, missing information from either modality can lead to ambiguous and incorrect results.
-Therefore, we route each instance based on representations from both modalities.
-Specifically, for layer $L^{i}$, we apply a layer-wise learned linear transformation $\mathbf{W}_{y}^{i}\in\mathbb{R}_{d_{y}\times d_{c}}$ to get the cross-modal routing embedding based on cross-modal feature $\psi_{y}$, where $d_{c}$ is the embedding dimension for cross-modal routing.
-We also utilize the global-level feature from previous layer $L^{i-1}$ to compute a inter-modal routing embedding through $\mathbf{W}_{x}^{i}\in\mathbb{R}_{d_{x}\times d_{i}}$, where $d_{i}$ is the inter-modal routing embedding, and $d_{i}+d_{c}=d_{r}$.
-Finally, we concatenate both embeddings along the embedding dimension to get an adjoint embedding $\mathbf{k}\in\mathbb{R}_{d_{r}}$, and dot product them with the routing embedding of all available experts.
+Mixture of Prompt Experts
+To mitigate the negative impact on optimizing a long prompt and to improve the adaptiveness of dynamic prompt, we propose to learn multiple short prompts with a MoE-like design.
+To be more specific, we use MoPE module in each prompt-tuned Transformer layers $L^{i}$ to generate the dynamic prompt in a scalable way.
+A MoPE module consists of a router, $k$ prompt experts and their associated routing embeddings $\{(\mathbf{E}_{j},\mathbf{k}_{j})\}_{j=1}^{k}$ where $\mathbf{E}_{i}\in\mathbb{R}_{l\times d}$ is an expert, $\mathbf{k}_{i}\in\mathbb{R}_{d_{r}}$ is its routing embedding, and $d_{r}$ is the dimension of routing embedding.
+We route each instance based on representations of all modalities, as illustrated in Fig.
+2-(b,c).
+The router is parameterized as two layer-specific linear transformations $\mathbf{W}_{y}^{i}\in\mathbb{R}_{d_{y}\times d_{c}}$ and $\mathbf{W}_{x}^{i}\in\mathbb{R}_{d_{x}\times d_{i}}$, where $d_{c}$, $d_{i}$ are the dimension of cross-modal and inter-modal routing embedding, respectively, and $d_{i}+d_{c}=d_{r}$.
+The cross-model embedding is projected from $\psi_{y}$, while inter-model embedding is calculated over the global-level feature (typically represented as the [CLS] token) from previous layer $L^{i-1}$.
+Finally, we concatenate both embeddings to get an multimodal query embedding $\mathbf{q}\in\mathbb{R}_{d_{r}}$, and compute the dot product with the routing embeddings of all available experts.
 The routing score $\mathbf{r}$ is calculated by:
-$\mathbf{k}=[\psi_{y}\mathbf{W}_{y}^{i},x_{0}^{i}\mathbf{W}_{x}^{i}],\quad\mathbf{r}_{j}=\frac{\operatorname{exp}(\mathbf{k}^{\top}\mathbf{v}_{j}/\tau+\epsilon)}{\sum_{n=1}^{k}\operatorname{exp}(\mathbf{k}^{\top}\mathbf{v}_{n}/\tau+\epsilon)}$ (3)
-where $\tau=0.1$ is the temperature hyper-parameter, and $\epsilon$ is sampled noise [36].
-The dynamic prompt is calculated by a convex combination of all experts at this layer according to the routing score:
+$\mathbf{q}=[\psi_{y}\mathbf{W}_{y}^{i},x_{0}^{i-1}\mathbf{W}_{x}^{i}],\quad\mathbf{r}_{j}=\frac{\operatorname{exp}(\mathbf{q}^{\top}\mathbf{k}_{j}/\tau+\epsilon)}{\sum_{n=1}^{k}\operatorname{exp}(\mathbf{q}^{\top}\mathbf{k}_{n}/\tau+\epsilon)}$ (3)
+where $\tau=0.1$ is the temperature hyper-parameter, and $\epsilon$ is sampled noise (Riquelme et al.
+2021).
+When there are more than one complementary modality, the additional embedding could be easily extended by learning additional projections.
+The dynamic prompt is obtained by a convex combination of all experts at this layer according to the routing score:
 $\mathbf{P}_{d}=\sum_{j=1}^{k}\mathbf{r}_{j}\mathbf{E}_{j}$ (4)
-3.4 Regularizing Expert Routing
-The proposed MoPE scales up vanilla prompt tuning by dividing the problem space into subspaces governed by specialized experts.
-However, we empirically found that a few experts tend to dominate across all instances during training, resulting in non-specialization.
-This phenomenon is self-reinforcing and is due to degenerated routing, where experts receiving high scores during initialization are exposed to larger gradient scales, causing them to be selected more frequently during optimization [5, 38].
-In this section, we introduce methods to circumvent this issue and aid expert specialization.
+Regularizing Expert Routing
+We empirically find that a few experts tend to dominate across all instances during training, a phenomenon that has also been observed in previous MoE methods (Eigen, Ranzato, and Sutskever 2013; Shazeer et al.
+2017).
+In this section, we introduce methods to circumvent degenerated routing and aid expert specialization.
 Orthogonal Routing Embedding.
-Inspired by the analysis of degenerated routing, we freeze the routing embeddings to avoid self-reinforcement.
-We find that an orthogonal initialized [37] and non-learnable routing embedding $\{\mathbf{v}\}_{j=1}^{k}$ performs as well as (and even slightly better than) learned embeddings, while requiring fewer trainable parameters.
+To avoid the expert being selected in initialization being exposed to a larger gradient signal, we freeze the routing embeddings $\{\mathbf{k}\}_{j=1}^{k}$ to suppress self-reinforcing expert domination.
+Our finding indicate that an orthogonal initialized (Saxe, McClelland, and Ganguli 2013) and non-learnable routing embedding outperform learned one, while requiring fewer trainable parameters.
 Importance Loss.
-To aid expert specialization, we add an additional importance loss [38, 36] to penalize dominant experts.
-For a batch of input $\mathbf{B}=\{(x_{0},y_{0}),(x_{1},y_{1}),...,(x_{b},y_{b})\}$, the importance of expert-$j$ is defined as the summed routing score in this batch.
-The importance loss is calculated as the mean coefficient of variation of all experts’ importance averaged across all layers::
-$\operatorname{Imp}(\mathbf{E}_{j})=\sum_{(x,y)\in\mathbf{B}}\mathbf{r}_{j},\quad\mathcal{L}_{imp}=\operatorname{stopgrad}\left(\left(\frac{\operatorname{std}(\{\operatorname{Imp}(\mathbf{E}_{j})\}_{j}^{k})}{\operatorname{mean}(\{\operatorname{Imp}(\mathbf{E}_{j})\}_{j}^{k})}\right)^{2};\gamma\right)$ (5)
-where $\operatorname{stopgrad}(\cdot)$ is the stop-gradient operator to prevent error propagation of this loss term when the coefficient of variation is less than a pre-defined threshold $\gamma=0.1$.
-While this loss was initially proposed for balancing computational budgets [38, 36], we adapt its use for promoting expert specialization.
+To further aid expert specialization, we add an additional importance loss (Shazeer et al.
+2017; Riquelme et al.
+2021) to penalize dominant experts.
+For a batch of input $\mathbf{B}=\{(x_{0},y_{0}),(x_{1},y_{1}),...,(x_{b},y_{b})\}$, the importance of expert-$j$ is defined as the summed routing score in this batch, $\operatorname{Imp}(\mathbf{E}_{j})=\sum_{(x,y)\in\mathbf{B}}\mathbf{r}_{j}$.
+The importance loss is calculated as the mean coefficient of variation of all experts’ importance averaged across all layers:
+$\mathcal{L}_{imp}$ $=\sum_{All\ layers}\operatorname{\sigma}\left(\left(\frac{\operatorname{std}(\{\operatorname{Imp}(\mathbf{E}_{j})\}_{j}^{k})}{\operatorname{mean}(\{\operatorname{Imp}(\mathbf{E}_{j})\}_{j}^{k})}\right)^{2};\gamma\right)$ (5)
+where $\operatorname{\sigma}(\cdot)$ is the stop-gradient operator to prevent error propagation of this loss term when the coefficient of variation is less than a pre-defined threshold $\gamma=0.1$.
+While this loss was initially proposed for balancing computational budgets (Shazeer et al.
+2017; Riquelme et al.
+2021), we adapt its use for promoting expert specialization.
 We add an additional threshold constraint due to our instance-wise routing, which is more likely to have a larger coefficient of variation than per-token routing.
-4 Experiments
-4.1 Implementation Details
+I+L A+L A+V+L
+Method Param Speed (ms) SNLI-VE UPMC Food MM-IMDB MUsTARD MUsTARD
+Acc ($\uparrow$) Acc ($\uparrow$) F1-macro/F1-micro ($\uparrow$) Pre ($\uparrow$) Pre ($\uparrow$)
+fine-tuning ImgOnly 86.9M 16.45$\pm$0.2 33.34 75.64 39.21/53.85 - -
+TextOnly 109.0M 7.28$\pm$0.9 69.58 86.92 58.80/65.37 65.41 65.41
+LateConcat 196.0M 22.79$\pm$0.9 72.01 93.19 60.43/67.77 68.82 69.40
+SeqFuse 197.0M 23.32$\pm$1.2 74.28 93.73 59.22/66.34 68.13 71.35
+MMBT 196.5M 15.91$\pm$1.2 67.58 94.10 60.80/66.10 - -
+prompt-tuning P-ImgOnly 0.1M 21.01$\pm$1.1 33.34 76.65 33.70/50.04 - -
+P-TextOnly 0.1M 7.57$\pm$1.2 64.86 81.01 52.19/61.16 59.27 59.27
+P-LateConcat 1.3M 28.59$\pm$2.1 64.29 90.27 56.95/64.23 60.71 65.17
+P-SeqFuse 1.1M 28.52$\pm$1.9 65.01 81.27 55.57/63.98 63.72 65.73
+P-MMBT 0.9M 16.71$\pm$1.3 67.58 81.07 52.95/59.30 - -
+PromptFuse <0.1M 28.75$\pm$0.9 64.53 82.21 48.59/54.49 63.76 64.20
+BlindPrompt <0.1M 29.52$\pm$1.2 65.54 84.56 50.18/56.46 62.01 63.80
+PromptFuse(†) 0.1M 29.26$\pm$1.7 64.94 82.14 50.78/60.96 64.73 66.29
+PMF 2.5M 32.21$\pm$0.8 71.92 91.51 58.77/64.51 - -
+Ours ($k\!=\!4$) 1.6M 30.47$\pm$1.4 73.14 91.54 61.93/68.19 67.12 67.35
+Ours ($k\!=\!16$) 2.6M 30.44$\pm$1.3 73.59 92.05 62.01/68.24 68.73 69.94
+Table 1: Quantitative results on multimodal classification.
+Our method achieve the best performance and parameter-efficiency against all prompt-fusion methods.
+I: Image, L: Language, A: Audio, V: Video.
+The main modality is underlined, and the best prompt tuning method is in bold, and (†): Our re-implementation with prompt applied to all layers.
+Method Param RefCOCO, mIOU($\uparrow$) RefCOCO+, mIOU($\uparrow$)
+val testA testB val testA testB
+SeqFuse 231.0M 53.48 55.76 52.03 40.22 42.2 37.91
+P-SeqFuse 35.1M 47.69 46.23 45.81 30.66 31.48 28.79
+PromptFuse(†) 35.1M 43.23 39.71 47.74 27.72 33.37 23.67
+Ours 35.5M 58.40 60.03 53.23 43.80 46.12 38.88
+Table 2: Quantitative result on referring image segmentation We report the total number of parameters in million (including the trainable segmentation head), and metrics (mIOU) on RefCOCO and RefCOCO+.
+Experiments
+Datasets.
+Following previous prompt fusion methods (Liang, Zhao, and Schütze 2022; Li et al.
+2023), we conduct experiments on MM-IMDB (Arevalo et al.
+2017) and UPMC Food-101 (Wang et al.
+2015) for classification, SNLI-VE (Xie et al.
+2019) for visual entailment (with image premise only), MUsTARD (Castro et al.
+2019) for sarcasm detection in speaker-dependent setting.
+We additionally test our method on the RefCOCO and RefCOCO+ (Kazemzadeh et al.
+2014) dataset for referring expression segmentation.
 Architecture Details.
-For all experiments, unless otherwise specified, we use Swin-B [28] as the vision encoder and Bert-base [3] as the text encoder.
-Vision is used as main modality.
-Following the experiment setup in [11, 21], we also finetune a linear head for each dataset.
-We implement the mapper $f_{m}(\cdot)$ as a two-layer MLP with GeLU [9] nonlinearily.
-Regarding prompt tuning, we use $l=6$ prompts and $k=16$ experts by default, and the tunable prompt is applied to all layers of the main modality encoder.
-Vanilla prompt tuning [20] with $l^{\prime}=4$ is used to tune the complementary modality.
+In main experiment, we use Swin-B (Liu et al.
+2021b) as the image encoder, Bert-base (Devlin et al.
+2018) as the text encoder, Wav2Vec2 (Baevski et al.
+2020) as the audio encoder.
+For the video modality, we use the same frame averaging technique by adapting a ViT (Dosovitskiy et al.
+2020) encoder, which follows PromptFuse (Liang, Zhao, and Schütze 2022).
+Image is used as main modality unless otherwise specified.
+Following the experiment setup in (Jia et al.
+2022; Li et al.
+2023; Liang, Zhao, and Schütze 2022), we finetune dataset-specific heads.
+We implement the mapper as a two-layer MLP with GeLU nonlinearily.
+Regarding prompt, we set $l=6$ and use $k=16$ experts by default, and the prompts are applied to all layers of the main modality encoder.
+Vanilla prompt tuning (Li and Liang 2021) with $l^{\prime}=4$ is used to tune the $\mathcal{E}_{\mathbb{Y}}$.
 Training Details.
-All images are resized and cropped to the size of $224\times 224$.
-All models are trained for 15 epoches, using the AdamW [29] optimizer with a learning rate of $4\times 10^{-4}$ for vision and $5\times 10^{-4}$ for text.
-We use a constant step decay scheduler that scales the learning rate at epochs 3 and 6 by a factor of $\gamma=0.4$.
-All models are trained on a server with a RTX4090 GPU.
-The mean and standard deviation of metrics, measured over 4 different experiments are reported when possible.
-4.2 Datasets
-MM-IMDB [1] is a multimodal movie classification dataset.
-It consists of 25,956 image-text pairs, where each pair includes a movie poster and a plot summary The dataset supports multi-label genre classification across a spectrum of 23 genres with a long-tail distribution.
-UPMC Food-101 [41] is a comprehensive multimodal dataset designed for fine-grained recipe classification with 90,840 image-text pairs for 101 food classes.
-SNLI-VE [43] is a large-scale multimodal dataset with 565,286 image-text pairs.
-The task for this dataset is visual entailment, which requires the model to infer whether a hypothesis matches the given premise.
-In our experiments, following [21], we only use the image premise.
-However, it is worth noting that some other works may also include the text premise.
-4.3 Compared Methods
-First, we compare our method with several baselines for multimodal fusion.
-ImgOnly / TextOnly. Only fine-tune one encoder, and the input of the other modality is discarded.
-P-ImgOnly / P-TextOnly. Only prompt-tune one encoder, and the input of the other modality is discarded.
+All models are trained for 20 epoches, using the AdamW (Loshchilov and Hutter 2017) optimizer with a learning rate of $4\times 10^{-4}$ for main modality and $5\times 10^{-4}$ for complementary modality.
+All models are trained with a RTX-4090 GPU.
+Further experiment details could be found in the Supplementary Material.
+Compared Methods. We consider following baselines:
+ImgOnly / TextOnly. Fine-tune one encoder only, and the input of the other modality is discarded.
+P-ImgOnly / P-TextOnly. Only prompt-tune one encoder.
 LateConcat. This baseline involves fine-tuning both encoders, concatenating their features, and learn a linear head for classification.
-P-LateConcat. Similar as LateConcat but prompt-tune each encoder instead of fine-tuning. However, the linear head is still fine-tuned.
-SequentialFuse.
-This method first extracts features from the complementary modality and maps them to the embedding space of the main modality encoder in the same way as our mapped prompt for end-to-end training.
+P-LateConcat. Similar as LateConcat but prompt-tune each encoder instead of fine-tuning.
+SeqFuse.
+This baseline first extracts features from the complementary modality and maps them to the embedding space of the main modality encoder by a MLP.
 Both encoders are fine-tuned.
 This is a strong baseline that can be considered as our method without MoPE, but with all parameters fine-tuned.
-P-SequentialFuse. Prompt-tune each encoders in SequentialFuse.
-For a fair comparison, all the prompt-tuning methods listed above use the same prompt length ($l=6$) as our method.
-In addition to these baselines, we also compare our methods with existing prompt-based fusion methods, including MMBT [15], Frozen [39], PromptFuse and BlindPrompt [22], and PMF [21].
-Among these methods, the setting in PromptFuse [22] is the most similar to ours, as it also assumes one modality to be a black-box.
-Table 1: Quantitative results on three multimodal dataset.
-We compare all baseline methods, existing prompt-based fusion methods, and our method with different expert numbers.
-Our method achieve state-of-the-art performance and parameter-efficiency.
-(†): Our implementation with prefix-tuning.
-Method Param SNLI-VE($\uparrow$) Food-101($\uparrow$) MM-IMDB($\uparrow$)
-finetuning ImgOnly 86.9M 33.34$\pm$.03 75.64$\pm$.21 39.21$\pm$.32/53.85$\pm$.21
-TextOnly 109.0M 69.58$\pm$.24 86.92$\pm$.21 58.80$\pm$.31/65.37$\pm$.13
-LateConcat 196.0M 72.01$\pm$.23 93.19$\pm$.21 60.43$\pm$.68/67.77$\pm$.36
-SequentialFuse 197.0M 74.28$\pm$.46 93.73$\pm$.16 59.22$\pm$.38/66.34$\pm$.19
-MMBT [15] 196.5M 67.58 94.10 60.80/66.10
-prompt-tuning P-ImgOnly 0.1M 33.34$\pm$.01 76.65$\pm$.07 33.70$\pm$.55/50.04$\pm$.27
-P-TextOnly 0.1M 64.86$\pm$.32 81.01$\pm$.50 52.19$\pm$.32/61.16$\pm$.43
-P-LateConcat 1.3M 64.29$\pm$.28 90.27$\pm$.40 56.95$\pm$.37/64.23$\pm$.26
-P-SequentialFuse 1.1M 65.01$\pm$.18 81.27$\pm$.22 55.57$\pm$.63/63.98$\pm$.35
-P-MMBT [21] 0.9M 67.58 81.07 52.95/59.30
-PromptFuse [22] <0.1M 64.53 82.21 48.59/54.49
-BlindPrompt [22] <0.1M 65.54 84.56 50.18/56.46
-PromptFuse(†) [22] 0.1M 64.94 82.14 50.78/60.96
-PMF [21] 2.5M 71.92 91.51 58.77/64.51
-Ours (ViT-b, $k=4$) 1.6M 73.47$\pm$.11 91.55$\pm$.12 62.37$\pm$.35/68.73$\pm$.24
-Ours (Swin-b, $k=4$) 1.6M 73.14±.$\pm.21$ 91.54$\pm$.21 61.93$\pm$.37/68.19$\pm$.14
-Ours (Swin-b, $k=16$) 2.6M 73.59$\pm$.15 92.05$\pm$.11 62.01$\pm$.21/68.24$\pm$.12
-Table 2: Ablation on three types of prompt disentanglement.
-The results from all combinations of mapped prompt, dynamic prompt, and static prompt are presented.
-Our full method with all prompts achieves the best result.
-Prompt SNLI-VE($\uparrow$) Food-101($\uparrow$) MM-IMDB($\uparrow$)
+P-SeqFuse. Prompt-tune each encoders in SeqFuse.
+In addition to these baselines, we also compare our methods with existing prompt-based fusion methods, including MMBT (Kiela et al.
+2019), Frozen (Tsimpoukelli et al.
+2021), PromptFuse and BlindPrompt (Liang, Zhao, and Schütze 2022), and PMF (Li et al.
+2023).
+Among these methods, the setting in PromptFuse (Liang, Zhao, and Schütze 2022) is the most similar to ours, as it also assumes one modality to be a black-box and is by design compatible with more than two modalities.
+To ensure a fair comparison, all prompt-tuning baselines utilize the same prompt length ($l=6$) as our method.
+Prompt SNLI-VE($\uparrow$) UPMC Food($\uparrow$) MM-IMDB($\uparrow$)
 […]
@@
 […]
-$[\mathbf{P}_{s},\mathbf{P}_{d},P_{m}]$ $\mathbf{73.59\!\pm\!.15}$ $\mathbf{92.05\!\pm\!.11}$ $\mathbf{62.01\!\pm\!.21}$/$\mathbf{68.24\!\pm\!.12}$
-4.4 Main Results
-The quantitative results of all baselines, compared methods, and our method with different expert numbers $k$ are summarized in Tab 1.
-The metric on SNLI-VE and Food-101 is accuracy (%), and MM-IMDB is F1-Macro and F1-Micro (%).
-We also list the number of trainable parameters (million) for each method.
+$[\mathbf{P}_{s},\mathbf{P}_{d},P_{m}]$ $\mathbf{73.59}{\color[rgb]{.5,.5,.5}{\pm.15}}$15 $\mathbf{92.05}{\color[rgb]{.5,.5,.5}{\pm.11}}$11 $12\mathbf{62.01}{\color[rgb]{.5,.5,.5}{\pm.21}}/\mathbf{68.24}{\color[rgb]{.5,.5,.5}{\pm.12}}$12
+Table 3: Ablation on prompt decomposition.
+The results from all combinations of mapped prompt, dynamic prompt, and static prompt are presented.
+Our full method with all prompts achieves the best result.
+Quantitative Comparison with Previous Methods
+Multimodal Classification.
+The quantitative results on multimodal classification on all methods are summarized in Tab 1.
+The metric on SNLI-VE, UPMC Food-101 is accuracy (%), MM-IMDB is F1-Macro and F1-Micro (%), and on MuSTARD is precision (%).
+We also list the inference time (in milliseconds) and number of trainable parameters (in million) for each method.
 Our method outperforms all prompt-based methods, and is competitive with fine-tuning.
-Specifically, when compared with the fine-tuning baselines, SequentialFuse and LateConcat, our method delivers competitive accuracy on the Food-101 dataset and superior results on the SNLI-VE and MM-IMDB datasets, while requiring as few as 0.8% of the trainable parameters.
-Apart from surpassing baselines, the proposed method also outperforms existing prompt-based fusion methods, including PromptFuse [22], BlindPrompt [22], and PMF [21].
-Notably, our method outperforms the current SOTA, PMF [21], by a significant margin on all datasets, while requiring $37\%$ fewer parameters.
-Moreover, PMF assumes both encoders share the same Transformer architecture.
-By contrast, our method has weaker assumptions, which could be easily extended to any Transfomer-based architecture, such as ViT [4] and Swin [28].
-4.5 Qualitative Result of MoPE Routing
-During end-to-end tuning, prompt experts in MoPE spontaneously specialize in different concepts.
-Fig.
-3 presents two examples of this specialization.
-Specifically, we visualize the expert with the highest score as the routed expert for each instance pair.
+Specifically, when compared with the fine-tuning baselines, SeqFuse and LateConcat, our method delivers competitive accuracy on the UPMC Food-101 dataset and superior results on the SNLI-VE, MM-IMDB and MusTARD datasets, while requiring as few as 0.8% of the trainable parameters.
+The proposed method also outperforms all existing prompt fusion methods, including PromptFuse (Liang, Zhao, and Schütze 2022), BlindPrompt (Liang, Zhao, and Schütze 2022), and PMF (Li et al.
+2023).
+Notably, our method ($k=4$) outperforms the current SOTA, PMF, by a significant margin on all datasets, while requiring $37\%$ fewer parameters.
+Moreover, our method has a weak assumption than PMF, which assumes both encoders share the same Transformer architecture, making it hard to be extended to other modalities and other model architectures.
+Referring Expression Segmentation.
+Previous prompt fusion methods mainly test their performance on classification (Li et al.
+2023; Liang, Zhao, and Schütze 2022; Tsimpoukelli et al.
+2021).
+The improved expressiveness of MoPE allows learning of harder tasks, for example, referring image segmentation.
+Following experiment setting in VPT, We additionally finetune a SeTR-PUP (Zheng et al.
+2021) head for segmentation.
+The result is summarized in Tab.
+2.
+Our proposed method achieves the best result among all compared methods.
+In particular we outperform the fine-tuning baseline SeqFuse, which suffers from catastrophic forgetting problem on this task.
+Qualitative Result of MoPE Routing
+Through training, prompt experts in MoPE spontaneously specialize, and we present two examples in Fig.
+3.
+We visualize the expert with the highest score as the routed expert for each instance pair.
 In the provided examples, we observe that expert-4 specialize in children-related concepts, while expert-12 focuses on crowds.
-These examples also demonstrate the interpretability of our MoPE-based prompting, as opposed to black-box conditional prompt such as CoOp [50] and CoCoOp [49].
-We encourage readers for a more comprehensive visualizations in Section 0.C.2.
-5 Analysis and Discussion
+These examples also demonstrate the interpretability of MoPE, as opposed to black-box conditional prompt such as CoOp (Zhou et al.
+2022b).
+More visualizations could be found in supplementary material.
+Analysis and Discussion
 In this section, we systematically analyze the effects of our proposed design through ablation studies.
 We show that our MoPE-based fusion method is more adaptive and scalable for multimodal fusion compared to existing approaches.
-Our three-types of disentangled prompts are collaborative.
-We first ablate each type of prompt in our prompt disentanglement, and the results are summarized in Tab.
-2.
-Our full method achieves the best performance, indicating that the prompts collaborate effectively with each other.
-Specifically, the dynamic prompt is essential for multimodal fusion as it learns the interplay between different modalities.
-This is in contrast with previous methods [22, 39, 45], where each modality do not explicitly interact.
-Adding the dynamic prompt induces a significant gain of 13.5%, 13.26, 8.4%/6.5% on all datasets.
-Without the mapped prompt, the fine-grained information from complimentary modality is lost, leading to a significant drop in all datasets.
-The static prompt is also necessary, serving as a special expert that is always routed.
-Adding this special expert allows the other experts to focus more on capturing instance-level features, resulting in a slight performance gain.
+Our decomposed prompts are collaborative.
+We ablate each type of prompt in our instance-wise prompt decomposition, with results across 3 random seeds reported in Tab.
+3 .
+Our full method achieves the best performance, indicating effective prompt collaboration.
+Adding the dynamic prompt yields gains of 13.5%, 13.26%, 8.4%/6.5% on all datasets, this is different from previous methods (Liang, Zhao, and Schütze 2022; Tsimpoukelli et al.
+2021) where modalities don’t explicitly interact.
+Without the mapped prompt, fine-grained information from complementary modalities is lost, causing significant drops.
+The static prompt is also necessary, acting as a special expert always routed for capturing global-level features.
 Figure 4: More experts v.s.
@@
 To study the expressiveness of MoPE, we compare expert-scaling with length-scaling.
-Our starting point for MoPE is $l=6$ prompts and $k=2$ experts (when $k\leq 1$, MoPE degenerate into vanilla prompt), which account for $(2+1)\times 6+1=19$ tunable prompts in total per-layer.
-We increase the number of $k$, and at each step we also report the result of simply increasing $l$ in vanilla prompt, up to the same total number of prompts.
-The results are summarised in Fig.
+Our starting point for MoPE is $l=6$ prompts and $k=2$ experts, which account for $(2+1)\times 6+1=19$ tunable prompts in total per-layer.
+We increase the number of $k$, and at each step we also report the result of increasing $l$ in vanilla prompt to the same total number of prompts.
+The results are summarized in Fig.
 4.
-As illustrated in the figure, adding the MoPE design with as few as $k=2$ experts boosts performance, and increasing the number of experts leads to a monotonic performance improvement.
-Although our method may also saturate when the number of experts becomes excessive, it does not face the over-length deterioration issue observed in vanilla prompt tuning.
-In contrast, merely augmenting the number of prompts $l$ in vanilla prompt tuning does not lead to a linear performance improvement and is prone to performance deterioration.
-This observation is consistent with previous findings [11, 44, 14, 20].
-This makes the prompt length $l$ a hyperparameter that is difficult to tune manually.
-In addition to accuracy considerations, longer prompts in vanilla prompt tuning can exacerbate computational overhead due to the quadratic time complexity of self-attention.
+As illustrated in the figure, adding the MoPE design with as few as $k=2$ experts boosts performance, and increasing the number of experts leads to a monotonic performance improvement which avoids over-length deterioration.
+In contrast, length-scaling does not lead to a linear performance improvement and is prone to performance deterioration, which is consistent with previous findings (Jia et al.
+2022; Yang et al.
+2022a; Khattak et al.
+2023; Li and Liang 2021).
+Additionally, longer prompts in vanilla prompt tuning can exacerbate computational overhead due to the quadratic time complexity of self-attention.
 This is why we cannot experiment with $l>199$ for the other two methods.
-On the other hand, MoPE scales by conditioning the dynamic prompt on more experts while maintaining a fixed length of learnable prompts during self-attention.
+By contrast, MoPE scales by conditioning the dynamic prompt on more experts while maintaining a fixed prompt length during actual self-attention.
 As a result, our method maintains a nearly constant time complexity.
-Table 3: Result of dense routing v.s.
-sparse routing.
-Dense routing achieves slightly better results on all three datasets.
-Routing SNLI-VE $\uparrow$ Food-101 $\uparrow$ MM-IMDB $\uparrow$ Dense 73.55 91.74 62.01/68.25 Sparse 71.15 90.89 60.28/66.01 Table 4: Different way of generating routing embedding.
-Frozen routing embedding with orthogonal initialization is as good as learned one.
-Routing Embed SNLI-VE $\uparrow$ Food-101 $\uparrow$ MM-IMDB $\uparrow$ Frozen 73.55 91.74 62.01/68.25 Learned 73.13 91.20 61.64/67.97
-[Uncaptioned image] Figure 5: Scaling performance with increased training data.
+Figure 5: Scaling performance with increased training data.
 We show the performance of our method and representative methods as we progressively increase the amount of training data, or “#shots”.
-Our method outperform other prompt-tuning methods at all data scales.
-MoPE scales better when more training data is available.
-Previous prompting methods have been shown to not scale well with respect to increased training data [7, 22, 39, 44].
-We study how the proposed MoPE method performs on different data scales to assess its scalability.
-Specifically, we sub-sample the training-set with different shots to represent a range of low-shot to high-shot learning scenarios.
-We then train our method and other prompt-tuning and fine-tuning methods with the same subsampled data.
-The mean accuracy and F1-Micro score on SNLI-VE and MM-IMDB, respectively, are reported in Fig.
-5.
-As the figure shows, our MoPE-based method is the most scalable, achieving superior results across all data scales.
-Specifically, we consistently match the results of the fine-tuning method, SequentialFuse, on the SNLI-VE dataset, and we achieve superior results on the MM-IMDB dataset.
-By contrast, the method based on vanilla-prompt does not scale well with respect to increased training data, and there is a consistent performance gap between ours and the compared method.
-This performance gap becomes more significant on larger datasets, such as when training with $10^{5}$ shots or using the full training set.
-Refer to caption Figure 6: Routing determinism. We show the entropy of routing score during training. Dense routing are more deterministic as optimization progress. Better viewed with color.
+Our method outperform other prompt fusion methods at all data scales.
+MoPE scales better with more training data.
+Previous prompting methods have been shown to not scale well with respect to increased training data (Gao et al.
+2022; Liang, Zhao, and Schütze 2022; Tsimpoukelli et al.
+2021; Yang et al.
+2022a).
+To assess the scalability of MoPE, we sub-sample the training set with different shots to simulate a range of low-shot to high-shot learning scenarios.
+We then train our method and other prompt-tuning and fine-tuning methods on the same subset of SNLI-VE and MM-IMDB.
+As Fig.
+5 shows, our MoPE-based method demonstrates superior scalability compared to other prompt tuning methods.
+Specifically, we consistently match the results of the fine-tuning method, SeqFuse, on the SNLI-VE dataset, while surpassing it on the MM-IMDB dataset.
+By contrast, the methods based on vanilla prompts do not scale well with respect to increased training data, resulting in a consistent performance gap between our method and the compared methods.
+This performance disparity becomes more pronounced on larger datasets, such as when training with $10^{5}$ shots or using the full training set.
 Figure 6: Effect of the importance loss.
@@
 (a) Without importance loss, only a few experts are used throughout training (b) The importance loss ensures balanced utilization of all experts.
-Dense routing with frozen routing embedding achieves the best result.
-Previous MoE with FFN acting as experts usually employ a sparse gating function that only select one or a few experts per token, while we use a dense router (Eqn.
-4) that weigh all experts.
-To compare the effect, we follow  [36] to implement a sparse router by adding a Top-1 gate: $\mathbf{r}^{\prime}=\operatorname{TOP}_{1}(\mathbf{r})$.
-The results of using sparse routing versus dense routing are reported in Tab. 5. Our experiments indicate that using the dense router yields marginally better results.
-To further understand how dense routing benefits the model, we evaluated the degree of uncertainty in routing by calculating the empirical entropy of the routing score $\mathcal{H}(\mathbf{r})=-\sum_{i=1}^{n}r_{i}log_{2}(r_{i})$, averaged across training batches and all layers.
-In essence, high entropy indicates that routing behaves more randomly.
-The relationship between entropy and optimization steps is illustrated in Fig.
-6.
-As demonstrated in the figure, dense routing is more deterministic.
-In addition to routing sparsity, we also investigate the effect of using frozen routing embeddings. As Tab. 5 shows, fixed routing embedding could achieve the same result as the trained one.
+Frozen routing embedding outperform learned embedding.
+We ablate the effect of using frozen routing embeddings.
+As Tab.
+4 shows, fixed routing embedding could achieve the same result as the trained one.
+Routing Embed SNLI-VE $\uparrow$ UPMC Food $\uparrow$ MM-IMDB $\uparrow$
+Frozen 73.55 91.74 62.01/68.25
+Learned 73.13 91.20 61.64/67.97
+Table 4: Result of frozen routing embedding. Frozen routing embedding are slightly better than learned one.
 Importance loss aids expert specialization.
@@
 In Fig.
-7, we visualize how the importance of each expert (i.e., average routing score) changes when training on the SNLI-VE dataset.
-Without the importance loss, routing adheres to its initial state, resulting in a skewed distribution where a few experts dominate.
-The importance loss allows other experts to be used by penalizing highly unbalanced expert importance.
-Architecture Tuning MM-IMDB ($\uparrow$)↑ )
-BoW* Fine-tune 48.20/57.50
-Bert [3] Frozen 58.86/66.13
-GPT-2 [35] Frozen 34.03/50.84
-Longformer [3] Bitfit [46] 60.34/67.27
-Table 5: Our fusing method is highly modular. We treat complementary modality as black-box, allowing different models to be plug-in for fusion. (*): Bag-of-word initialized with Bert word embeddings.
-Our fusion method are highly modular.
-Our method abstracts the complimentary modality as a black-box, allowing high modality of both modalities.
-In particular, our method allow arbitrary models to be seamless plug-in for multimodal fusion.
-We give some results in Tab.
+6, we visualize how the importance (i.e., average routing score) of each expert changes when training on the SNLI-VE dataset.
+Without the importance loss, routing adheres to its initial state, resulting in a skewed distribution where a few experts are always being routed.
+Adding it aids expert specialization by penalizing highly unbalanced expert importance.
+How MoPE could be more expressive.
+We have demonstrated that the proposed MoPE scales better with trainable parameters and data.
+This increased expressiveness comes from two factors.
+First, previous methods use a shared prompt for all instances and neglect multimodal interplay.
+By contrast, our dynamic prompt with MoPE is adaptive, and finding an instance-wise optimal prompt can lead to better results.
+Secondly, MoPE circumvents the challenge of optimizing a long prompt (Petrov, Torr, and Bibi 2023).
+By doing so, we shift the difficulty from learning a universal long prompt to learning a parameterized router with multiple short prompts.
+As a result, our empirical expressiveness is closer to the theoretical upper bound of prompt tuning.
+Conclusion
+In this paper, we propose an adaptive and expressive prompt-based method for fusing unimodal models.
+Our method involves decomposing the prompt for instance-adaptive prompt learning.
+To effectively scale up the prompt while avoiding performance deterioration, we introduce MoPE, which improves the expressiveness of prompt tuning.
+Extensive experiments demonstrate that our method is highly parameter-efficient and scales better with dataset size and prompt length.
+We believe MoPE could be an efficient fusion method for a wide array of multimodal tasks.
+Appendix A Additional Details on Experiments
+In this section, we provide additional details on our model architecture, experimental setup, and data preprocessing.
+Architecture Details
+Prompt Vector.
+Our prompt implementation closely follows VPT (Jia et al.
+2022).
+Specifically, for static prompts and prompt experts, we use uniform initialization $\mathcal{U}\sim(-\eta,\eta)$, where $\eta$ is calculated according to the embedding dimension and patch size of the Transformer (Jia et al.
+2022).
+Dropout with $p=0.1$ is applied to all prompts.
+However, we do not use the reparameterization trick for prompts introduced in the original prompt tuning methods (Li and Liang 2021), as the gradients of our dynamic prompt and mapped prompt are already rectified by MLPs (i.e., $\mathbf{W}_{x}$, $\mathbf{W}_{y}$, $f_{m}(\cdot)$).
+For Transformer architectures that employ a window attention mechanism (e.g., Swin (Liu et al.
+2021b)), we duplicate the same prompt and prepend it to all windows for self-attention calculation, following the approach in VPT (Jia et al.
+2022).
+Mapper.
+We learn a mapper to map representations from complementary modality $\mathbb{Y}$ to the embedding dimension of main modality $\mathbb{X}$.
+Generally speaking, the mapper is implemented as a two-layer MLP with a bottleneck design, which shares similarities with previous work (Li et al.
+2023).
+In our experiments, we set the bottleneck dimension as half of the dimension of the complementary representation, i.e., $d_{bot}=\lceil d_{y}/2\rceil$.
+Then, we apply a batch normalization layer and a GeLU activation to obtain a bottleneck feature $\psi_{bot}\in\mathbb{R}_{d_{bot}}$.
+After obtaining this bottleneck feature, we apply another linear layer to project it to the dimension of $d_{x}$.
+However, some Transformer architectures use inconsistent $d_{x}$ in different layers.
+For example, in Swin-b (Liu et al.
+2021b), the embedding dimension changes from $[128,256,512,1024]$, making it challenging to fit a single linear transformation.
+To circumvent this, we instead learn four separate linear projections, each used to project the shared bottleneck feature to different embedding dimensions.
+As a result, there is one single down-sampling layer and multiple up-sampling layers.
+The design is illustrated in Fig 7.
+Figure 7: Architecture of the Mapper for Swin. The cross model feature is first projected to a low-dimensional bottleneck, then further mapped to different embeddding dimension of each Swin blocks.
+MoPE.
+As discussed in the main body of this paper, we learn a per-layer linear projection $\mathbf{W}_{x}^{i}$ to obtain the cross-modal embedding.
+Here, we would like to further clarify that the weight is not shared with the one used in the mapper $f_{m}$.
+In our experiments, we use $d_{c}=8$ and $d_{i}=2$, resulting in $d_{r}=10$.
+The ablation on the two dimensions can be found in the second part of this supplementary material.
+For the importance loss, we scale it by a factor of $0.01$ and linearly combine it with the task-specific loss(es) for optimization.
+Dataset
+UPMC Food-101 (Wang et al.
+2015) serves as a comprehensive multimodal dataset designed for fine-grained recipe classification.
+The dataset contains 90,840 image-text pairs for 101 food classes.
+We follow previous methods (Li et al.
+2023; Kiela et al.
+2019) to create a validation set of 5000 samples.
+MM-IMDB (Arevalo et al.
+2017) is a multimodal movie classification dataset.
+It comprises 25,956 pairs of images and texts, each pair including a movie poster and a plot summary.
+The dataset supports multi-label genre classification across a spectrum of 23 genres with highly imbalanced classes.
+SNLI-VE (Xie et al.
+2019) is a large-scale multimodal dataset with 565,286 image-text pairs.
+The task for this dataset is visual entailment, which means that the model should decide whether a hypothesis matches the given premise or not.
+This dataset provide image and text premise, while the hypothesis is always in text modality.
+MUsTARD (Castro et al.
+2019) is a dataset for multimodal sarcasm detection.
+The original dataset contains 690 video clips in MP4 format with an even number of sarcastic and non-sarcastic labels.
+Each video is provided with annotations in the language modality, describing the speaker and sentence in that video.
+Collectively, there are three modalities available: video, audio extracted from the video, and the language modality.
+RefCOCO and RefCOCO+ (Kazemzadeh et al.
+2014) are two datasets for referring image segmentation.
+Both datasets are built based on MSCOCO and contain around 19,900 images with 50,000 objects and 140,000 referring expressions.
+RefCOCO allows referring expressions of any type, while RefCOCO+ features expressions that do not contain object positions.
+In Fig. 8, we provide examples of each dataset.
+Figure 8: Examples of each dataset used in the main paper.
+Data Processing
+For all input images, we perform RandAug with $N=2$ and $M=5$, resize them to $(256,256)$, and perform center cropping to obtain images of size $(224,224)$.
+For the SNLI-VE dataset, we follow Frozen and PMF and use only the image premise, which means that the input to the model is the image premise + text hypothesis.
+Note that this setting may differ from other works that also use the text premise.
+For the MuSTARD dataset, the overall processing pipeline follows PromptFuse (Liu et al.
+2022).
+For the audio modality, we extract WAV audio from the video using a sampling rate of 16,000 Hz.
+We also remove background noise using the Librosa package.
+For a batch of input, we zero-pad the WAV input to the same length.
+As for the video, while PromptFuse (Liu et al.
+2022) uses a face detection model to only sample frames where the speaker’s face is visible, we evenly sample $8$ frames for all videos in time.
+The sampled frames are resized to $(224,224)$ for encoding.
+For the text modality, we use the speaker-dependent setting, which means that the speaker’s name is visible to the model during both training and inference.
+To achieve this, we simply concatenate the speaker’s name to the input utterance string.
+Appendix B Extended Analysis and Discussion
+This section aims to provide further ablation study on the design choices and hyper-parameters. We also provided extended discussion on the effectiveness of proposed MoPE.
+Additional Ablations
+Ablation on dimension of $d_{i}$ and $d_{c}$.
+Our finding indicate that $d_{c}>>d_{i}$ results in better performance in general, and we set $d_{c}=8,d_{i}=2$ in our main experiment as ablated in Tab.
 5.
-How could MoPE be more expressive & Limitations.
-We have demonstrated that the MoPE-based method is more expressive for multimodal fusion.
-Compared with existing prompt-based methods, our approach scales better with data and trainable parameters.
-This increased expressiveness comes from two factors.
-First, the prompts used in previous methods are not adaptive, meaning that a single prompt is used for all instances.
-Finding a instance-wise optimal prompt could lead to better results.
-Secondly, we scale the prompt by using more experts, since a long prompt can be difficult to optimize [32] for optimal performance.
-By doing so, we shift the difficulty from learning a universal long prompt to learning a parameterized router with multiple short prompts.
-As a result, our empirical expressiveness is closer to the theoretical upper bound of prompt-tuning.
-Despite improving empirical expressiveness, our prompt still functions in the same way as vanilla prompts during attention, resulting in same theoretical upper-bound of expressiveness.
-This is also reflected by the small gap between ours and fine-tuning in large datasets such as SNLI-VE and UPMC Food-101.
-Future studies could focus on designing a more expressive prompting technique.
-6 Conclusion
-In this paper, we enhance the scalability and adaptivity of prompts used in multimodal fusion.
-Our method involves disentangling the prompt into static and dynamic components to enable instance-adaptive prompt learning.
-To circumvent performance deterioration due to overly long prompts, we introduce the mixture-of-prompt-expert technique, which improves the expressiveness of prompt-tuning.
-Extensive experiments demonstrate that our method is highly parameter-efficient and scales better with dataset size and prompt length.
-We have also shown that our method is interpretable and highly modular.
-Appendix 0.A Additional Experiment Details
-In this section, we provide additional details of our implementation and training.
-Prompt Vector.
-Our prompt implementation closely follows [11].
-Specifically, for static prompts and prompt experts, we use uniform initialization $\mathcal{U}\sim(-\eta,\eta)$, where $\eta$ is calculated according to the embedding dimension and patch size of the Transformer [11].
-Dropout with $p=0.1$ is applied to all prompts.
-However, we do not use the reparameterization trick for prompts [20], as the gradients of our dynamic prompt and mapped prompt are already rectified by MLPs.
-For Transformer architectures that employ a window attention mechanism, we duplicate the same prompt and prepend it to all windows for attention calculation, following the approach in  [11].
-Mapper.
-The main objective of the mapper is to map representations from dimension $y$ to the embedding dimension of modality $x$.
-Generally speaking, the mapper is implemented as a two-block MLP, which shares similarities with previous work [21].
-In our experiments, we set the bottleneck dimension as half of the dimension of the complementary representation, i.e., $d_{bot}=\lceil d_{y}/2\rceil$.
-Then, we apply a batch normalization layer and a GeLU activation to obtain a bottleneck feature $\psi_{bot}\in\mathbb{R}{d_{bot}}$.
-After obtaining this bottleneck feature, we apply another MLP to project it to the embedding dimension of $d_{x}$.
-However, some Transformer architectures use inconsistent $d_{x}$ in their internal layers.
-For example, in Swin-b[28], the embedding dimension changes from $[128,256,512,1024]$, making it challenging to fit a single MLP.
-To circumvent this, we instead learn four separate linear projections, each used to project the shared bottleneck feature to different embedding dimensions.
-As a result, there is one single down-sampling MLP and multiple up-sampling MLPs.
-To ensure fairness, we also use a similar design in plain Transformers such as ViT[4] and BERT [3].
-MoPE.
-As discussed in the main body of this paper, we learn a per-layer linear projection to obtain the cross-model embedding.
-Here, we would like to further clarify that this projection weight is not shared with the one used in the mapper.
-In our experiments, we use $d_{c}=8$ and $d_{i}=2$, resulting in $d_{r}=10$.
-For the importance loss, we scale it by a factor of $0.01$ and apply it together with the task-specific losses.
-Appendix 0.B Extended Analysis and Discussion
-0.B.1 Analysis on Adaptivity of vanilla prompt and MoPE
+$k=4$ $d_{i}=10$, $d_{c}=0$ $d_{i}=8$, $d_{c}=2$ $d_{i}=5$, $d_{c}=5$ $d_{i}=2$, $d_{c}=8$ $d_{i}=0$, $d_{c}=10$
+MM-IMDB($\uparrow$) 54.50/64.63 60.12/67.02 61.22/67.51 61.93/68.19 60.91/67.93
+Table 5: Ablation on $d_{c}$ and $d_{i}$. $d_{c}>>d_{i}$ is better.
+Ablation on vision encoder.
+In main experiment we use swin-base as the vision encoder.
+To have a fair comparision on previous prompt fusion methods (Li et al.
+2023; Liang, Zhao, and Schütze 2022) that are based on ViT (Dosovitskiy et al.
+2020), we provide the result of our method with ViT-base as vision encoder in Tab.
+6, with results measured using 3 random seeds, As the table shows, using the ViT as vision encoder gives similar results.
+The result demonstrate that our superior performance is not due to more advanced vision encoder.
+Method Param SNLI-VE($\uparrow$) UPMC Food($\uparrow$) MM-IMDB($\uparrow$)
+Ours(ViT-B) 1.6M 73.47$\pm$.11 91.55$\pm$.12 62.37$\pm$.35/68.73$\pm$.24
+Ours(Swin-B) 1.6M 73.14$\pm$.21 91.54$\pm$.21 61.93$\pm$.37/68.19$\pm$.14
+Table 6: Ablation on ViT as the main vision encoder.
+Analysis on Adaptivity of Vanilla Prompt and MoPE
 In the main body of the paper, we have empirically demonstrated that the dynamic prompt generated by MoPE exhibits greater adaptivity compared to global-shared prompt tuning.
@@
 Through this analysis, we will also provide intuitive insights into the significance of expert specialization in MoPE-based fusion.
-Following the approach adopted in previous papers [42, 32], we will focus our analysis on the case of a single prompt within a single Transformer layer.
+Following the approach adopted in previous papers (Wang et al.
+2023; Petrov, Torr, and Bibi 2023), we will focus our analysis on the case of a single prompt within a single Transformer layer.
 Since our method is designed to generate an effective prompt while leaving the attention calculation the same as previous prompt tuning, our MoPE will only affect the forward behavior up to the calculation of attention map.
@@
 The objective is to find a prompt $\mathbf{P}$ that enables the attention map $\mathbf{A}(\mathbf{X},\mathbf{P})$ to closely match the desired target attention pattern for each input instance.
-Theorem 0.B.1 (Limited Adaptivity of Global-shared Prompt)
+Theorem 1 (Limited Adaptivity of Global-shared Prompt)
 Let $\mathcal{X}$ be the input space, and $\mathcal{A}$ be the space of attention matrices.
@@
 Then, for vanilla prompting with a single shared prompt $\mathbf{P}$, there exists no $\mathbf{P}\in\mathcal{P}$ such that $\mathcal{A}^{*}(\mathbf{x})\subseteq\mathcal{A}(\mathbf{x},\mathbf{P})$ for all $\mathbf{x}\in\mathcal{X}$, where $\mathcal{P}$ is the prompt space.
-Proof
+Proof B.1
 Let $\mathbf{x}_{1},\mathbf{x}_{2}\in\mathcal{X}$ be two distinct input instances with disjoint optimal attention sets, i.e., $\mathcal{A}^{*}(\mathbf{x}_{1})\cap\mathcal{A}^{*}(\mathbf{x}_{2})=\emptyset$.
@@
 $\Delta(\mathbf{x}_{1},\mathbf{P}_{\text{shared}}^{*})+\Delta(\mathbf{x}_{2},\mathbf{P}_{\text{shared}}^{*})\geq\Delta(\mathbf{x}_{1},\mathbf{P}_{1}^{*})+\Delta(\mathbf{x}_{2},\mathbf{P}_{2}^{*})$ (10)
-Furthermore, by the limited expressiveness of prompt-tuning [32], we have:
+Furthermore, by the limited expressiveness of prompt-tuning (Petrov, Torr, and Bibi 2023), we have:
 $\Delta(\mathbf{x}_{2},\mathbf{P}_{2}^{*})\geq 0;\\ \Delta(\mathbf{x}_{1},\mathbf{P}_{1}^{*})\geq 0$ (11)
@@
 $\Delta(\mathbf{x}_{1},\mathbf{P}_{\text{shared}}^{*})+\Delta(\mathbf{x}_{2},\mathbf{P}_{\text{shared}}^{*})\geq\Delta(\mathbf{x}_{1},\mathbf{P}_{1}^{*})+\Delta(\mathbf{x}_{2},\mathbf{P}_{2}^{*})\geq 0$ (12)
-Theorem 0.B.2 (Improved Adaptivity of MoPE)
+Theorem 2 (Improved Adaptivity of MoPE)
 Let =$\mathcal{E}=\{\mathcal{E}_{1},\mathcal{E}_{2},\ldots,\mathcal{E}_{k}\}$ be a set of $k$ expert attention mappings, where each $\mathcal{E}_{i}:\mathcal{X}\rightarrow\mathcal{A}$ maps an input $\mathbf{x}\in\mathcal{X}$ to a set of attention patterns $\mathcal{E}_{i}(\mathbf{x})\subseteq\mathcal{A}$.
 Define the induced attention mapping of MoPE as:
-[…]
+$\mathcal{A}_{\text{MoPE}}(\mathbf{x},\mathcal{E},\mathbf{r})=\left\{\sum_{i=1}^{k}r_{i}(\mathbf{x})\mathbf{A}_{i}~{}\bigg{|}~{}\mathbf{A}_{i}\in\mathcal{E}_{i}(\mathbf{x}),\forall i\right\}$ (13)
 where $\mathbf{r}(\mathbf{x})=[r_{1}(\mathbf{x}),\ldots,r_{k}(\mathbf{x})]$ are the routing weights for input $\mathbf{x}$.
 Let $\mathcal{X}^{\prime}\subseteq\mathcal{X}$ be a set of instances. If the convex hull of the expert attention mappings, denoted as:
-[…]
-[…]
+$\text{conv}(\mathcal{E})=$ (14)
+$\left\{\sum_{i=1}^{k}\alpha_{i}\mathbf{A}_{i}~{}\bigg{|}~{}\mathbf{A}_{i}\in\mathcal{E}_{i}(\mathbf{x}),\forall i,\forall\mathbf{x}\in\mathcal{X}^{\prime},\sum_{i=1}^{k}\alpha_{i}=1,\alpha_{i}\geq 0\right\}$
+contains an optimal attention pattern for each instance in $\mathcal{X}^{\prime}$, i.e., $\mathcal{A}^{*}(\mathbf{x})\subseteq\text{conv}(\mathcal{E}),\forall\mathbf{x}\in\mathcal{X}^{\prime}$, then there exists a routing score $\mathbf{r}^{*}$ such that the accumulated attention discrepancy for MoPE under $\mathbf{r}^{*}$ across instances in $\mathcal{X}^{\prime}$ is equal to the sum of the optimal instance-wise attention discrepancies, i.e.,
 $\sum_{\mathbf{x}\in\mathcal{X}^{\prime}}\Delta(\mathbf{x},\mathcal{E},\mathbf{r}^{*})=\sum_{\mathbf{x}\in\mathcal{X}^{\prime}}\inf_{\mathbf{A}\in\mathcal{A}^{*}(\mathbf{x})}\|\mathbf{A}-\mathbf{A}^{*}\|_{\mathcal{A}}$ (15)
-Proof
-When there are more experts than instances, i.e.
-$|\mathcal{X}^{\prime}|\leq k$ , the proof is trivial.
+Proof B.2
+When there are more experts than instances, i.e., $|\mathcal{X}^{\prime}|\leq k$ , the proof is trivial.
 This is because the the optimal prompt for each instance could simply be “stored” in one or a few experts.
@@
 This implies that the attention discrepancy for each instance $\mathbf{x}\in\mathcal{X}^{\prime}$ under the routing score $\mathbf{r}^{*}$ is equal to the optimal instance-wise attention discrepancy:
-$\Delta(\mathbf{x},\mathcal{E},\mathbf{r}^{*})=\inf_{\mathbf{A}\in\mathcal{A}_{\text{MoPE}}(\mathbf{x},\mathcal{E},\mathbf{r}^{*})}\|\mathbf{A}-\mathbf{A}^{*}\|_{\mathcal{A}}=\inf_{\mathbf{A}\in\mathcal{A}^{*}(\mathbf{x})}\|\mathbf{A}-\mathbf{A}^{*}\|_{\mathcal{A}},\quad\forall\mathbf{x}\in\mathcal{X}^{\prime}$ (18)
+$\Delta(\mathbf{x},\mathcal{E},\mathbf{r}^{*})$ $=\inf_{\mathbf{A}\in\mathcal{A}_{\text{MoPE}}(\mathbf{x},\mathcal{E},\mathbf{r}^{*})}\|\mathbf{A}-\mathbf{A}^{*}\|_{\mathcal{A}}$ (18)
+$=\inf_{\mathbf{A}\in\mathcal{A}^{*}(\mathbf{x})}\|\mathbf{A}-\mathbf{A}^{*}\|_{\mathcal{A}},\quad\forall\mathbf{x}\in\mathcal{X}^{\prime}$
 Therefore, the accumulated attention discrepancy for MoPE under the routing score $\mathbf{r}^{*}$ across instances in $\mathcal{X}^{\prime}$ is equal to the sum of the optimal instance-wise attention discrepancies:
@@
 Thus, when the number of instances exceeds the number of experts, if the convex hull of the expert attention mappings contains an optimal attention pattern for each instance, then there exists a routing score $\mathbf{r}^{*}$ that allows MoPE to achieve an accumulated attention discrepancy equal to the sum of the optimal instance-wise attention discrepancies across those instances.
-Putting it together, the Theorem 0.B.1 state that the global-shared prompt (i.e.
-vanilla prompt) could not achieve the best result, when the input instances require different attention patterns to perform well.
-Equation 15 state that it is possible for MoPE to achieve the best result on all of the input instances, conditioned on the appropriate specialization of experts.
-0.B.2 Discussion: How to Choose the Main Modality?
+Putting it together, the Theorem.
+1 state that the global-shared prompt (i.e., vanilla prompt) could not achieve the best result, when the input instances require different attention patterns to perform well.
+Theorem.
+2 state that it is possible for MoPE to achieve the best result on all of the input instances, conditioned on the appropriate specialization of experts.
+Discussion: How to Choose the Main Modality?
 In the proposed method, we use a sequential pipeline to fuse different modalities.
-In this pipeline, the input of the complementary modality $y$ is first encoded into a representation $\psi_{y}$, which is then used to guide the prompting of modality $x$.
+In this pipeline, the input of the complementary modality $\mathbb{Y}$ is first encoded into a representation $\psi_{y}$, which is then used to guide the prompting of modality $\mathbb{X}$.
 This design raises an interesting question: “How to choose the complementary and main modality?”
@@
 We have also tested using text as the main modality, and the results are summarized in Tab.
-6.
-Table 6: Result of Language Model (LM) as main modality.
+7.
 Method SNLI-VE Food-101 MM-IMDB
@@
 Ours(VM) 73.59$\pm$.15 92.05$\pm$.11 62.01$\pm$.21/68.24$\pm$.12
+Table 7: Result of Language Model (LM) as main modality.
+Figure 9: t-SNE visualization of multimodal representation generated by different methods. The representation generated by our method is the most separable. Better viewed with color.
 As the table shows, using BERT as the main encoder results in a decline in performance.
 Our postulation is that for these three tasks, the text input contains a significant amount of noisy and false positive data.
-For instance, in the UPMC-Food 101 dataset [41], the text data is derived using a spider, which includes many irrelevant hypertext and website titles.
-Therefore, treating text as the main modality can lead to overfitting and biased predictions.
-For tasks that require a modality-specific dense representation for decoding, such as text generation and segmentation, the main modality must be the one that provides such output.
-For example, we use the BERT encoder to guide Swin for segmentation, and we use Swin to guide BERT for visual question answering (detail in Appendix 0.C).
-0.B.3 Our MoPE-based method yields a better multimodal representation.
+For instance, in the UPMC-Food 101 dataset (Wang et al.
+2015), the text data is derived using a spider, which includes many irrelevant hypertext and website titles.
+Therefore, treating the text as the main modality can lead to overfitting and biased predictions.
+For tasks that require a modality-specific dense representation for decoding, the main modality must be the one that provides such output for compatibility.
+For example, in referring expression segmentation, we have to use the image as main modality, because the segmentation head expects feature maps instead of a global representation.
+Our MoPE-based method yields a better multimodal representation.
 We visualize the multimodal representation generated by different methods using t-SNE.
@@
 For the other methods, we visualize the [CLS] token of the main modality.
-For ease of coloring, we only show the first 20 classes in the test set of the UPMC Food-101 [41] dataset.
+For ease of coloring, we only show the first 20 classes in the test set of the UPMC Food-101 (Wang et al.
+2015) dataset.
 The results are presented in Fig.
-8.
+9.
 As the results show, the representation generated by our method is the most separable.
-Figure 8: t-SNE visualization of multimodal representation generated by different methods. The representation generated by our method is the most separable. Better viewed with color.
-Appendix 0.C Boarder Applications
-In this section, we explore the broader applications of our MoPE-based multimodal fusion method beyond classification tasks by providing additional experiments on dense prediction tasks.
-0.C.1 Referring Image Segmentation
-The proposed fusion method focuses on utilizing prompts to adapt the Transformer encoders to derive a multimodal representation.
-Based on this representation, we can tackle tasks beyond classification, such as segmentation.
-In this section, we briefly introduce how our method can be applied to this scenario and provide experimental results.
-Datasets and Task.
-RefCOCO [13] and RefCOCO+[13] are the commonly used datasets for referring image segmentation (RIS), both built upon the MS-COCO[24] dataset.
-RefCOCO contains 142,209 referring expressions for 50,000 objects across 19,994 images, while RefCOCO+ includes 141,564 expressions for 49,856 objects in 19,992 images.
-The task involves predicting a segmentation mask from the input image according to a given referring expression in natural language.
-Architecture and Training Details.
-We utilize ImageNet-pretrained Swin-B-384 as our vision encoder, and we use BERT as the text encoder.
-For the decoder head, we adopt the standard SETR-PUP [48] head.
-Following [11], the head is fine-tuned, while we use the MoPE-based tuning method to adapt the encoders.
-The pyramid feature maps from the 4 Swin layers are fed into the segmentation head for end-to-end training.
-The optimization uses the same hyperparameters as in our main paper.
-Compared Models.
-To the best of our knowledge, there is no existing prompt-based multimodal fusion method that is compatible with dense visual tasks.
-Notably, the previous state-of-the-art method, PMF [21], assumes that each encoder uses the same Transformer architecture, and they did not open-source their training code.
-Furthermore, their prompt tuning is only applied to the deep layers, while SETR-PUP [48] requires pyramid features from internal layers.
-Therefore, we find it challenging to adapt PMF for dense tasks such as segmentation.
-Given these challenges, we primarily compare our method with our baselines, SequentialFuse and p-SequentialFuse.
-Quantitative Comparison with Baselines.
-The quantitative results, measured as mean Intersection over Union (mIoU) on the validation and test sets of RefCOCO and RefCOCO+, are reported in Tab.
-7.
-As shown in the table, our MoPE-based method achieves the best performance.
-We outperform the prompt-based method, p-SequentialFuse, by a significant margin.
-Additionally, our method performs better than the fine-tuning method.
-We observe that the fine-tuning method tends to overfit in this experimental setting.
-Table 7: Segmentation results on RefCOCO and RefCOCO+. We report the total number of parameters in million (including the trainable segmentation head), and metrics (mIOU) on RefCOCO and RefCOCO+.
-Method Param RefCOCO($\uparrow$) RefCOCO+($\uparrow$)
-val testA testB val testA testB
-SequentialFuse 231.0M 53.48 55.76 52.03 40.22 42.2 37.91
-P-SequentialFuse 35.1M 47.69 46.23 45.81 30.66 31.48 28.79
-Ours 35.5M 58.40 60.03 53.23 43.80 46.12 38.88
-Visualization of segmentation result.
-We visualize the segmentation results in Figures 9 and 10.
-As the results show, our MoPE-based segmentation method can better locate the object according to the referring image segmentation task.
-In contrast, the method based on global-shared prompts, p-SequentialFuse, may struggle to understand the text instructions and provide incorrect results.
-Figure 9: Quantitative result of refering image segmentation. Better viewed with color.
-Figure 10: Quantitative result of refering image segmentation (cont). Better viewed with color.
-0.C.2 Additional Examples of Expert Routing on VQAv2
+Appendix C Additional Visualizations
+Visualization of Referring Expression Segmentation Results
+In Fig.
+10, Fig.
+11 we provide a qualitative comparision on our method and the P-SeqFuse baseline.
+As illustrated in the figures, our methods could correctly understand the referring expression and localize object mask accordingly.
+By contrast, the compared method may fail to follow the text guidance, leading to less accurate segmentation results.
+Figure 10: Visualization on Referring Expression Segmentation - 1
+Figure 11: Visualization on Referring Expression Segmentation - 2
+Additional Examples of Expert Routing on VQAv2
 Our MoPE is designed to scale up the expressiveness of vanilla prompts, and expert specialization is a critical condition to achieve superior results as well as interpretability.
-To further demonstrate expert specialization, we train our MoPE-based method on an even larger dataset, VQAv2 [8].
+In the main body of the paper, we have provided visualization results of expert routing on the SNLI-VE dataset.
+We found the expert specialization are more observable on large dataset with a high heteriogenity.
+Hence, to further demonstrate expert specialization, we train our MoPE-based method on an even larger dataset, VQAv2 (Goyal et al.
+2017).
 This dataset contains 265,016 images and paired questions, covering a wide array of visual and textual concepts.
-We train our model on this dataset and visualize the routing results in Figures 11, 12, 13, and 14.
+We train our model on this dataset and visualize the routing results in Figures 12, 13, 14, and 15.
 We observe clear expert specialization in these examples, where different experts capture different concepts.
 Figure 12: Additional example of expert routing - 1.
-We show routing result of expert-4 on VQAv2 dataset, which specialize in animals.
-Note: Sorted by routing score, images and texts may not in same order.
+We show routing result of expert-4 on the VQAv2 dataset, which specialize in animals.
+Note: Sorted by routing score, images and texts may not be in the same order.
 Figure 13: Additional example of expert routing - 2.
-We show routing result of expert-6 on VQAv2 dataset, which specialize in lighting conditions.
-Note: Sorted by routing score, images and texts may not in same order.
+We show routing result of expert-6 on the VQAv2 dataset, which specialize in lighting conditions.
+Note: Sorted by routing score, images and texts may not be in the same order.
 Figure 14: Additional example of expert routing - 3.
-We show routing result of expert-8 on VQAv2 dataset, which specialize in toilet-related concepts and hairstyles.
-Note: Sorted by routing score, images and texts may not in same order.
+We show routing result of expert-8 on the VQAv2 dataset, which specialize in toilet-related concepts and hairstyles.
+Note: Sorted by routing score, images and texts may not be in the same order.
 Figure 15: Additional example of expert routing - 4.
-We show routing result of expert-10 on VQAv2 dataset, which specialize in weather conditions.
-Note: Sorted by routing score, images and texts may not in same order.
+We show routing result of expert-10 on the VQAv2 dataset, which specialize in weather conditions.
+Note: Sorted by routing score, images and texts may not be in the same order.

### v3

@@
-MoPE: Parameter-Efficient and Scalable Multimodal Fusion via Mixture of Prompt Experts
+MoPE: Mixture of Prompt Experts for Parameter-Efficient and Scalable Multimodal Fusion
 Abstract
 Despite the demonstrated parameter efficiency of prompt-based multimodal fusion methods, their limited adaptivity and expressiveness often result in suboptimal performance compared to other tuning approaches.
-In this paper, we address these limitations by decomposing the vanilla prompts to adaptively capture instance-level features.
-Building upon this decomposition, we introduce the mixture of prompt experts (MoPE) technique to enhance the expressiveness of prompt tuning.
-MoPE leverages multimodal pairing priors to route the most effective prompt on a per-instance basis.
+In this paper, we introduce the Mixture of Prompt Experts (MoPE), the first technique designed to overcome these limitations by decomposing standard prompts to capture instance-level features adaptively.
+Building on this decomposition, MoPE enhances prompt fusion’s expressiveness by leveraging multimodal pairing priors to route the most effective prompt for each instance dynamically.
 Compared to vanilla prompting, our MoPE-based fusion method exhibits greater expressiveness, scaling more effectively with the training data and the overall number of trainable parameters.
-We also investigate regularization terms for expert routing, which lead to emergent expert specialization during training, paving the way for interpretable soft prompting.
-Extensive experiments across six multimodal datasets spanning four modalities demonstrate that our method achieves state-of-the-art results for prompt fusion, matching or even surpassing the performance of fine-tuning while requiring only 0.8% of the trainable parameters.
-Introduction
-Unimodal pre-trained models like BERT (Radford et al.
-2019) and Swin Transformer (Liu et al.
-2021b) have excelled in transferring capabilities to a wide range of tasks.
+We also investigate regularization terms for expert routing, which lead to emergent expert specialization with enhanced adaptiveness and interpretablity.
+Extensive experiments across six multimodal datasets spanning four modalities demonstrate state-of-the-art performance for prompt fusion, matching or even surpassing the performance of fine-tuning while requiring only 0.8% of the trainable parameters.
+Project homepage: https://github.com/songrise/MoPE.
+1 Introduction
+Unimodal pre-trained models like Bert [30] and Swin Transformer [25] have excelled in transferring capabilities to a wide range of tasks.
 By comparison, extending the pretraining-finetuning paradigm to multimodal applications involves additional complexities, particularly regarding flexibility and cost.
 First, the modality-paired pretraining can be inflexible due to its dependence on specific configurations.
-This can limit flexibility when there is a scarcity of certain modality pairs, or when new architectures emerge, necessitating resource-intensive retraining.
+When there is a scarcity of certain modality pairs, or when new architectures emerge, it necessitates resource-intensive retraining.
 In contrast, unimodal pretrained models are more readily available and offer greater flexibility, as they can be independently updated and combined for multimodal tasks.
 Second, as these foundation models scale in size, fine-tuning them becomes increasingly expensive, which further restricts their application to downstream tasks.
-To democratize foundation models, a compelling question arises: How can we efficiently combine separately pretrained unimodal models for multimodal tasks?
+To democratize foundation models, a compelling question arises:
+How can we efficiently combine separately pretrained unimodal models for multimodal tasks
 Figure 1: High-level motivation of MoPE-based multimodal fusion.
 (a) Vanilla prompt tuning learns a globally shared long prompt for all instances, which may not be optimal for each instance.
-(b) MoPE decomposes the global prompt into multiple short and specialized prompt experts to improve its adaptivity and expressiveness.
-While originally proposed for transfer learning (Li and Liang 2021; Jia et al.
-2022; Zhou et al.
-2022b), recent research has revealed that prompt tuning could also be a parameter-efficient paradigm for multimodal fusion (i.e., prompt fusion).
-Typically, this is achieved by representing one modality with several prompts, which are fed into the frozen Transformer of another modality to achieve fusion (Tsimpoukelli et al.
-2021; Liang, Zhao, and Schütze 2022; Li et al.
-2023; Yang et al.
-2022b).
-However, directly using prompts for fusion could yield suboptimal results, particularly due to their limited adaptivity and scalability.
-Notably, while prompt tuning generally performs well in low-data regimes, it can be less effective when applied to full-shot training on the entire dataset with a challenging objective (Gao et al.
-2022; Liang, Zhao, and Schütze 2022; Tsimpoukelli et al.
-2021; Yang et al.
-2022a; Lester, Al-Rfou, and Constant 2021).
-This reduced efficacy could be attributed to at least two factors.
-First, the vanilla prompt was originally developed for unimodal learning, which optimizes globally-shared prompts for all instances (Jia et al.
-2022; Li and Liang 2021; Tsimpoukelli et al.
-2021).
-When directly applied for prompt fusion (Tsimpoukelli et al.
-2021; Li et al.
-2023; Liang, Zhao, and Schütze 2022), it fails to adaptively exploit the multimodal interplay and can lead to sub-optimal representations for each input instance.
-Second, the relatively limited representation power (compared to fine-tuning) can lead to underfitting in multimodal datasets with a long-tail distribution and complex cross-modal mapping (Liu et al.
-2021a; Petrov, Torr, and Bibi 2023; Wang et al.
-2023).
+(b) MoPE decomposes the unified long prompt into multiple specialized and short prompt experts to improve its adaptivity and expressiveness.
+While originally proposed for transfer learning [19, 11, 43], recent research has revealed that prompt tuning could adapted for parameter-efficient multimodal fusion (i.e., prompt fusion).
+Typically, this is achieved by representing one modality with several prompts and feed into the frozen Transformer of another modality [34, 21, 20, 41].
+However, directly using prompts for fusion could yield suboptimal results.
+It has been observed in various papers that while prompt tuning generally performs well in low-data regimes, it can be less effective when applied to full-shot training on larger datasets with a challenging objective [9, 21, 34, 40, 18].
+The reduced efficacy of prompt fusion could be attributed to its limited adaptivity and expressiveness.
+Specifically, these methods typically employ globally shared prompts [34, 20, 21] adopted from vanilla prompt tuning for all instances, which fails to capture the nuanced, instance-specific multimodal interactions.
+Additionally, the constrained expressiveness of prompt (compared to fine-tuning) can lead to underfitting in multimodal datasets with a long-tail distribution and complex cross-modal mapping [24, 28, 37].
 To address these challenges, increasing the number of learnable prompts, known as “length-scaling”, appears to be a straightforward solution.
-However, numerous studies have shown that the performance gains from length-scaling quickly reach saturation in both transfer-learning (Gao et al.
-2022; Lester, Al-Rfou, and Constant 2021) and fusion settings (Liang, Zhao, and Schütze 2022; Tsimpoukelli et al.
-2021; Yang et al.
-2022a).
-Furthermore, over-length prompts may even lead to worse results (Jia et al.
-2022; Yang et al.
-2022a; Khattak et al.
-2023; Li and Liang 2021; Kim et al.
-2023).
-Recent theoretical analyses have substantiated these empirical observations, highlighting the limited expressiveness of prompt tuning (Wang et al.
-2023; Petrov, Torr, and Bibi 2023).
-Overall, evidence from both empirical and theoretical analyses reveals that simply increasing prompt length is insufficient to address the limitations of prompt fusion.
-To improve the adaptivity and expressiveness of prompt fusion, we propose decomposing the unified, long prompt into multiple short prompts with specialized responsibilities, as depicted in Fig.
+Nevertheless, the performance gains from length-scaling quickly reach saturation in both transfer-learning [9, 18] and fusion settings [21, 34, 40].
+Furthermore, over-length prompts may even lead to worse results [11, 40, 14, 19, 16].
+Recent theoretical analyses have substantiated these empirical observations [37, 28], particularly regarding the difficulty in optimizing a unified long prompt vector.
+In this work, we present the first approach that specifically targets the enhancement of both adaptivity and expressiveness in prompt fusion.
+Our high-level idea is to divide-and-conquer the problem space utilizing multimodal pairing information, as depicted in Fig.
 1.
-We introduce an instance-wise adaptive prompt decomposition that separates the vanilla prompt into static, dynamic, and mapped prompts to adaptively capture global and local features.
-The foundation of our dynamic prompt generation lies in the proposed MoPE (Mixture of Prompt Experts) design, where we maintain a fixed prompt length in the self-attention and enhance expressiveness by learning multiple prompt experts and a multimodal router to select the most suitable prompt for each instance.
-We systematically compare our MoPE-based fusion with previous prompt fusion methods on a total of six datasets, and demonstrate consistently superior performance with higher parameter-efficiency.
-Furthermore, we showcase the superior scalability of our method both in terms of parameter and data.
-Our analysis indicates that increasing the number of experts (i.e., “expert-scaling”) is more scalable than length-scaling, with monotonic performance gains and the avoidance of performance deterioration with an overly-long prompt.
-Our method also scales better with more training data.
-By introducing regularization terms for expert routing, we observe the emergence of specialized prompt experts after end-to-end training, resulting in high interpretability.
+Specifically, we propose decomposing the unified, long prompt into three types of specialized, short prompts that are instance-wisely adaptive.
+To synthesize the most effective prompt per-instance, the MoPE module (Mixture of Prompt Experts) is introduced.
+Instead of increasing the prompt length, MoPE scales the parameter capacity of prompt fusion by adding more experts.
+This approach enhances expressiveness while avoiding the negative side effects associated with overly long prompt.
+We conduct systematic experiments on a total of six datasets spanning four modalities.
+Compared to existing prompt fusion methods, MoPE not only has better performance but also achieves a higher parameter efficiency.
+In our ablation analysis, we reveal that increasing the number of experts (i.e., “expert-scaling”) is more scalable than length-scaling, with monotonic performance gains and the avoidance of performance deterioration with an overly-long prompt.
+Furthermore, by introducing regularization terms for expert routing, we observe the emergence of specialized prompt experts after end-to-end training, resulting in high interpretability.
 Our key contributions are summarized as follows:
-  • We propose an instance-wise adaptive prompt decomposition, which decomposes the vanilla prompt into static, dynamic and mapped prompt.
-  • We introduce the MoPE technique for instance-wise dynamic prompt generation, which scales up the expressiveness of prompt tuning for multimodal fusion.
+  • We propose a novel instance-wise adaptive prompt decomposition to augment the adaptiveness of prompt fusion.
+  • We introduce the MoPE technique for instance-wise dynamic prompt generation, which scales up the expressiveness of vanilla prompt.
   • A combination of regularization terms is studied to aid specialization of prompt experts.
-  • Extensive experiments on six datasets spanning four modalities demonstrate state-of-the-art performance and parameter efficiency for prompt-based multimodal fusion.
-Related Works
+  • Extensive experiments on six datasets spanning four modalities demonstrate SOTA performance and parameter efficiency for prompt-based multimodal fusion.
+2 Related Works
 Prompt Tuning for Transfer Learning.
-Prompt tuning (Li and Liang 2021; Lester, Al-Rfou, and Constant 2021) learns continuous token embeddings as additional input to a frozen pretrained model for transfer learning of Transformer-based models.
-It is widely used in various modalities (Jia et al.
-2022; Bahng et al.
-2022; Liu et al.
-2022; Lin et al.
-2023; Jiang, Liu, and Chen 2023; Khattak et al.
-2023; Zhou et al.
-2022a).
-A common observation is good transfer learning performance in low-shot scenarios, yet its performance is less comparable to fine-tuning when abundant training instances are available (Li and Liang 2021; Gao et al.
-2022).
-Moreover, increasing prompt length quickly reaches performance saturation, and over-length prompts might lead to worse results (Yang et al.
-2022a; Jia et al.
-2022; Kim et al.
-2023).
-Recent theoretical analyses (Petrov, Torr, and Bibi 2023; Wang et al.
-2023) reveal that the expressiveness of prompt tuning is lower than that of fine-tuning.
+Prompt tuning [19, 18] learns continuous token embeddings as additional input to a frozen pretrained model for transfer learning of Transformer-based models.
+It is widely used in various modalities [11, 3, 23, 22, 12, 14, 42].
+A common observation is good transfer learning performance in low-shot scenarios, yet its performance is less comparable to fine-tuning when abundant training instances are available [19, 9].
+Moreover, increasing prompt length quickly reaches performance saturation, and over-length prompts might lead to worse results [40, 11, 16].
+Recent theoretical analyses [28, 37] reveal that the expressiveness of prompt tuning is lower than that of fine-tuning.
 In this work, we tackle this challenge by scaling up the expressiveness of prompts with a Mixture of Experts (MoE)-like design.
-Multimodal Fusion via Prompt.
+Prompt Fusion.
 Apart from model adaptation, prompts can also be used to fuse separately pretrained unimodal models for multimodal tasks.
-Frozen (Tsimpoukelli et al.
-2021) first introduced a method where the visual representation is mapped as a few input tokens to query frozen language models (LMs).
-PromptFuse and BlindPrompt (Liang, Zhao, and Schütze 2022) improved upon this by introducing tunable prompts to the LM for cross-modal alignment.
-PICa (Yang et al.
-2022b) translates images into discrete text captions to prompt frozen LMs.
-These methods treat tokens from different modalities independently and lack cross-modal interaction.
-Recently, PMF (Li et al.
-2023) introduced an interactive prompt fusion method for the vision and text modalities, building on the strong assumption that both encoders are white-box and have the same Transformer architecture.
+Frozen [34] first introduced a method where the visual representation is mapped as a few input tokens to query frozen language models (LMs).
+PromptFuse and BlindPrompt [21] improved upon this by introducing tunable prompts to the LM for cross-modal alignment.
+PICa [41] translates images into discrete text captions to prompt frozen LMs.
+These methods treat tokens from different modalities independently and lack explicit cross-modal interaction.
+Recently, PMF [20] introduced an interactive prompt fusion method for the vision and text modalities, building on the strong assumption that both encoders are white-box and have the same Transformer architecture.
 In this paper, we investigate how to adaptively interact with two or more modalities with minimal assumptions about the model architecture.
 MoE in Transformers.
-MoE is a computationally efficient technique to scale up models, including Transformers (Lepikhin et al.
-2020; Riquelme et al.
-2021; Mustafa et al.
-2022; Fedus, Zoph, and Shazeer 2022).
+MoE is a computationally efficient technique to scale up models, including Transformers [17, 31, 27, 8].
 The fundamental approach involves inserting MoE layers, usually composed of multiple feed-forward networks (FFNs) acting as experts, into the standard Transformer architecture.
-A router is learned to route each token embedding to the most suitable expert(s) for reducing computational overhead (Qin and Eisner 2021; Riquelme et al.
-2021; Shazeer et al.
-2017).
+A router is learned to route each token embedding to the most suitable expert(s) for reducing computational overhead [29, 31, 33].
 In this paper, we do not focus on reducing computational cost; instead, we introduce the MoE design into prompt fusion to scale up its adaptiveness and expressiveness.
-Method
+3 Method
 Figure 2: Architecture overview.
@@
 Better viewed with color.
-Preliminary: Vanilla Prompt Tuning
-Consider a Transformer (Vaswani et al.
-2017) or its variants used to extract features from an embedded input sequence $\mathbf{x}^{0}\in\mathbb{R}_{s\times d_{x}}$, where $s$ is the sequence length and $d_{x}$ is the embedding dimension.
-Prompt-tuning freezes all pre-trained parameters and optimize a small amount of continuous embeddings (i.e., prompts) $\mathbf{P}\in\mathbb{R}_{l\times d}$ concatenated to the input of each layer, where $l$ is length of the prompt.
+3.1 Preliminary: Vanilla Prompt Tuning
+Consider a Transformer [35] or its variants used to extract features from an embedded input sequence $\mathbf{x}^{0}\in\mathbb{R}_{s\times d_{x}}$, where $s$ is the sequence length and $d_{x}$ is the embedding dimension.
+Prompt-tuning freezes all pre-trained parameters and optimize a small number of continuous embeddings (i.e., prompts) $\mathbf{P}\in\mathbb{R}_{l\times d}$ concatenated to the input of each layer, where $l$ is the length of the prompt.
 The input of the $i$-th layer layer $L^{i}$ could be denoted as:
@@
 where $x^{i-1}_{0}\in\mathbb{R}_{d_{x}}$ denotes the [CLS] token, $\mathbf{T}^{i-1}\in\mathbb{R}_{s\times d_{x}}$ is the token embedding from the previous layer, and [,][,][ , ] denotes the concatenation operation.
-Succinctly, prompt tuning works by biasing the pretrained attention pattern in each Transformer layers (Petrov, Torr, and Bibi 2023).
-As opposed to fine-tuning, this biasing has strictly limited expressiveness in theory (Petrov, Torr, and Bibi 2023; Wang et al.
-2023), meaning that there are tasks that are un-learnable even with $l\to\infty$.
-The empirical expressiveness of vanilla prompting is even lower than its theoretical upper bound, due to competing optimization when $l>1$ (Petrov, Torr, and Bibi 2023).
+Succinctly, prompt tuning works by biasing the pretrained attention pattern in each Transformer layers [28].
+As opposed to fine-tuning, this biasing has strictly limited expressiveness in theory [28, 37], meaning that there are tasks that are un-learnable even with $l\to\infty$.
+This limitation characterizes the upper-bound of prompt tuning.
+On the other hand, the empirical expressiveness of vanilla prompting is even lower, which means that in practice, the prompt tuning tends to perform worse than its theoretical upper bound.
+This is due to competing optimization scheme when a long prompt $l>1$ [28] is employed.
 In other words, although longer prompt gives more trainable parameters, finding such a long prompt becomes challenging, and a suboptimal long prompt can lead to performance degradation instead.
 In this paper, we do not aim to increase the theoretical upper bound of prompt-tuning; instead, we approach the problem by narrowing the gap between its empirical and theoretical expressiveness.
-Instance-wise Adaptive Prompt Decomposition
-Previous prompt fusion methods (Liang, Zhao, and Schütze 2022; Yang et al.
-2022b) optimize a global prompt that is shared across all instances, neglecting the interplay between different modalities and instance-specific features.
+3.2 Instance-wise Adaptive Prompt Decomposition
+Previous prompt fusion methods [21, 41] optimize a global prompt that is shared across all instances, neglecting the interplay between different modalities and instance-specific features.
 With paired multimodal input as a prior, we aim to instance-wisely condition the prompting of one modality on the other(s) for better adaptivity.
@@
 In this example, expert-4 is specialized for children while expert-12 focuses on crowds.
-Mixture of Prompt Experts
+3.3 Mixture of Prompt Experts
 To mitigate the negative impact on optimizing a long prompt and to improve the adaptiveness of dynamic prompt, we propose to learn multiple short prompts with a MoE-like design.
@@
 $\mathbf{q}=[\psi_{y}\mathbf{W}_{y}^{i},x_{0}^{i-1}\mathbf{W}_{x}^{i}],\quad\mathbf{r}_{j}=\frac{\operatorname{exp}(\mathbf{q}^{\top}\mathbf{k}_{j}/\tau+\epsilon)}{\sum_{n=1}^{k}\operatorname{exp}(\mathbf{q}^{\top}\mathbf{k}_{n}/\tau+\epsilon)}$ (3)
-where $\tau=0.1$ is the temperature hyper-parameter, and $\epsilon$ is sampled noise (Riquelme et al.
-2021).
+where $\tau=0.1$ is the temperature hyper-parameter, and $\epsilon$ is sampled noise [31].
 When there are more than one complementary modality, the additional embedding could be easily extended by learning additional projections.
@@
 $\mathbf{P}_{d}=\sum_{j=1}^{k}\mathbf{r}_{j}\mathbf{E}_{j}$ (4)
-Regularizing Expert Routing
-We empirically find that a few experts tend to dominate across all instances during training, a phenomenon that has also been observed in previous MoE methods (Eigen, Ranzato, and Sutskever 2013; Shazeer et al.
-2017).
+We visualize the MoPE routing mechanism in Fig. 2-(b, c).
+3.4 Regularizing Expert Routing
+The proposed MoPE scales up vanilla prompt tuning by dividing the problem space into subspaces governed by specialized experts.
+However, we empirically find that a few experts tend to dominate across all instances during training, a phenomenon that has also been observed in previous MoE methods [7, 33].
 In this section, we introduce methods to circumvent degenerated routing and aid expert specialization.
@@
 To avoid the expert being selected in initialization being exposed to a larger gradient signal, we freeze the routing embeddings $\{\mathbf{k}\}_{j=1}^{k}$ to suppress self-reinforcing expert domination.
-Our finding indicate that an orthogonal initialized (Saxe, McClelland, and Ganguli 2013) and non-learnable routing embedding outperform learned one, while requiring fewer trainable parameters.
+Our finding indicate that an orthogonal initialized [32] and non-learnable routing embedding outperform learned one, while requiring fewer trainable parameters.
 Importance Loss.
-To further aid expert specialization, we add an additional importance loss (Shazeer et al.
-2017; Riquelme et al.
-2021) to penalize dominant experts.
+To further aid expert specialization, we add an additional importance loss [33, 31] to penalize dominant experts.
 For a batch of input $\mathbf{B}=\{(x_{0},y_{0}),(x_{1},y_{1}),...,(x_{b},y_{b})\}$, the importance of expert-$j$ is defined as the summed routing score in this batch, $\operatorname{Imp}(\mathbf{E}_{j})=\sum_{(x,y)\in\mathbf{B}}\mathbf{r}_{j}$.
@@
 where $\operatorname{\sigma}(\cdot)$ is the stop-gradient operator to prevent error propagation of this loss term when the coefficient of variation is less than a pre-defined threshold $\gamma=0.1$.
-While this loss was initially proposed for balancing computational budgets (Shazeer et al.
-2017; Riquelme et al.
-2021), we adapt its use for promoting expert specialization.
+While this loss was initially proposed for balancing computational budgets [33, 31], we adapt its use for promoting expert specialization.
 We add an additional threshold constraint due to our instance-wise routing, which is more likely to have a larger coefficient of variation than per-token routing.
@@
 Method Param Speed (ms) SNLI-VE UPMC Food MM-IMDB MUsTARD MUsTARD
-Acc ($\uparrow$) Acc ($\uparrow$) F1-macro/F1-micro ($\uparrow$) Pre ($\uparrow$) Pre ($\uparrow$)
+Acc % ($\uparrow$) Acc %($\uparrow$) F1-macro/F1-micro ($\uparrow$) Pre % ($\uparrow$) Pre % ($\uparrow$)
 fine-tuning ImgOnly 86.9M 16.45$\pm$0.2 33.34 75.64 39.21/53.85 - -
@@
 Table 2: Quantitative result on referring image segmentation We report the total number of parameters in million (including the trainable segmentation head), and metrics (mIOU) on RefCOCO and RefCOCO+.
-Experiments
-Datasets.
-Following previous prompt fusion methods (Liang, Zhao, and Schütze 2022; Li et al.
-2023), we conduct experiments on MM-IMDB (Arevalo et al.
-2017) and UPMC Food-101 (Wang et al.
-2015) for classification, SNLI-VE (Xie et al.
-2019) for visual entailment (with image premise only), MUsTARD (Castro et al.
-2019) for sarcasm detection in speaker-dependent setting.
-We additionally test our method on the RefCOCO and RefCOCO+ (Kazemzadeh et al.
-2014) dataset for referring expression segmentation.
+4 Experiments
+4.1 Dataset and Tasks
+UPMC Food-101 [36] serves as a comprehensive multimodal dataset designed for fine-grained recipe classification.
+The dataset contains 90,840 image-text pairs for 101 food classes.
+We follow previous methods [20, 15] to create a validation set of 5000 samples.
+MM-IMDB [1] is a multimodal movie classification dataset.
+It comprises 25,956 pairs of images and texts, each pair including a movie poster and a plot summary.
+The dataset supports multi-label genre classification across a spectrum of 23 genres with highly imbalanced classes.
+SNLI-VE [39] is a large-scale multimodal dataset with 565,286 image-text pairs.
+The task for this dataset is visual entailment, which means that the model should decide whether a hypothesis matches the given premise or not.
+This dataset provide image and text premise, while the hypothesis is always in text modality.
+MUStARD [4] is a dataset for multimodal sarcasm detection.
+The original dataset contains 690 video clips in MP4 format with an even number of sarcastic and non-sarcastic labels.
+Each video is provided with annotations in the language modality, describing the speaker and sentence in that video.
+Collectively, there are three modalities available: video, audio extracted from the video, and the language modality.
+RefCOCO and RefCOCO+ [13] are two datasets for referring image segmentation.
+Both datasets are built based on MSCOCO and contain around 19,900 images with 50,000 objects and 140,000 referring expressions.
+RefCOCO allows referring expressions of any type, while RefCOCO+ features expressions that do not contain object positions.
+4.2 Experiment Setups
+Metrics.
+The metric on SNLI-VE, UPMC Food-101 is accuracy (%), MM-IMDB is F1-Macro and F1-Micro, and on MuSTARD is precision (%).
+Mean Intersection over Union (mIoU) is the metric for RefCOCO and RefCOCO+.
 Architecture Details.
-In main experiment, we use Swin-B (Liu et al.
-2021b) as the image encoder, Bert-base (Devlin et al.
-2018) as the text encoder, Wav2Vec2 (Baevski et al.
-2020) as the audio encoder.
-For the video modality, we use the same frame averaging technique by adapting a ViT (Dosovitskiy et al.
-2020) encoder, which follows PromptFuse (Liang, Zhao, and Schütze 2022).
+In main experiment, we use pretrained Swin-B [25] as the image encoder, pretrained Bert-base [5] as the text encoder, and pretrained Wav2Vec2 [2] as the audio encoder.
+For video encoding, we follow PromptFuse [21] to use a ViT [6] for extracting image features from $n=8$ sampled frames, the averaged feature is used to represent the whole video.
 Image is used as main modality unless otherwise specified.
-Following the experiment setup in (Jia et al.
-2022; Li et al.
-2023; Liang, Zhao, and Schütze 2022), we finetune dataset-specific heads.
-We implement the mapper as a two-layer MLP with GeLU nonlinearily.
-Regarding prompt, we set $l=6$ and use $k=16$ experts by default, and the prompts are applied to all layers of the main modality encoder.
-Vanilla prompt tuning (Li and Liang 2021) with $l^{\prime}=4$ is used to tune the $\mathcal{E}_{\mathbb{Y}}$.
+Following the experiment setup in [11, 20, 21], we finetune dataset-specific head.
+Linear head is used for all classification tasks, and we use a standard UperNet [38] head for segmentation task.
+We implement the mapper as a two-layer MLP with GeLU nonlinearity.
+Regarding the prompt, we set $l=6$ and use $k=16$ experts by default, which strike a balance between performance and parameter size.
+The prompts are applied to all layers of the main modality encoder.
+Vanilla prompt tuning [19] with $l^{\prime}=4$ is used to tune the $\mathcal{E}_{\mathbb{Y}}$.
+Further implementation details could be found in the Supplementary Material.
 Training Details.
-All models are trained for 20 epoches, using the AdamW (Loshchilov and Hutter 2017) optimizer with a learning rate of $4\times 10^{-4}$ for main modality and $5\times 10^{-4}$ for complementary modality.
+All models are trained for 20 epoch, using the AdamW [26] optimizer with a learning rate of $4\times 10^{-4}$ for main modality and $5\times 10^{-4}$ for complementary modality.
 All models are trained with a RTX-4090 GPU.
-Further experiment details could be found in the Supplementary Material.
 Compared Methods. We consider following baselines:
@@
 This is a strong baseline that can be considered as our method without MoPE, but with all parameters fine-tuned.
-P-SeqFuse. Prompt-tune each encoders in SeqFuse.
-In addition to these baselines, we also compare our methods with existing prompt-based fusion methods, including MMBT (Kiela et al.
-2019), Frozen (Tsimpoukelli et al.
-2021), PromptFuse and BlindPrompt (Liang, Zhao, and Schütze 2022), and PMF (Li et al.
-2023).
-Among these methods, the setting in PromptFuse (Liang, Zhao, and Schütze 2022) is the most similar to ours, as it also assumes one modality to be a black-box and is by design compatible with more than two modalities.
+P-SeqFuse. Similar as SeqFuse but prompt-tune each encoders. This baseline is comparable to CoCoOp [42] but with additional static prompts.
+In addition to these baselines, we also compare our methods with existing prompt-based fusion methods, including MMBT [15], Frozen [34], PromptFuse and BlindPrompt [21], and PMF [20].
+Among these methods, the setting in PromptFuse [21] is the most similar to ours, as it also assumes one modality to be a black-box and is by design compatible with more than two modalities.
 To ensure a fair comparison, all prompt-tuning baselines utilize the same prompt length ($l=6$) as our method.
-Prompt SNLI-VE($\uparrow$) UPMC Food($\uparrow$) MM-IMDB($\uparrow$)
+Prompt SNLI-VE Acc($\uparrow$) UPMC Food Acc($\uparrow$) MM-IMDB F1($\uparrow$)
 […]
@@
 […]
-$[\mathbf{P}_{s},\mathbf{P}_{d},P_{m}]$ $\mathbf{73.59}{\color[rgb]{.5,.5,.5}{\pm.15}}$15 $\mathbf{92.05}{\color[rgb]{.5,.5,.5}{\pm.11}}$11 $12\mathbf{62.01}{\color[rgb]{.5,.5,.5}{\pm.21}}/\mathbf{68.24}{\color[rgb]{.5,.5,.5}{\pm.12}}$12
+[…]
 Table 3: Ablation on prompt decomposition.
@@
 Our full method with all prompts achieves the best result.
-Quantitative Comparison with Previous Methods
+4.3 Quantitative Comparisons
 Multimodal Classification.
-The quantitative results on multimodal classification on all methods are summarized in Tab 1.
-The metric on SNLI-VE, UPMC Food-101 is accuracy (%), MM-IMDB is F1-Macro and F1-Micro (%), and on MuSTARD is precision (%).
+The quantitative results of multimodal classification with all methods are summarized in Tab 1.
 We also list the inference time (in milliseconds) and number of trainable parameters (in million) for each method.
-Our method outperforms all prompt-based methods, and is competitive with fine-tuning.
+Our method outperforms all prompt-based fusion methods, and is competitive with fine-tuning.
 Specifically, when compared with the fine-tuning baselines, SeqFuse and LateConcat, our method delivers competitive accuracy on the UPMC Food-101 dataset and superior results on the SNLI-VE, MM-IMDB and MUStARD datasets, while requiring as few as 0.8% of the trainable parameters.
-The proposed method also outperforms all existing prompt fusion methods, including PromptFuse (Liang, Zhao, and Schütze 2022), BlindPrompt (Liang, Zhao, and Schütze 2022), and PMF (Li et al.
-2023).
+The proposed method also outperforms all existing prompt fusion methods, including PromptFuse [21], BlindPrompt [21], and PMF [20].
 Notably, our method ($k=4$) outperforms the current SOTA, PMF, by a significant margin on all datasets, while requiring $37\%$ fewer parameters.
-Moreover, our method has a weak assumption than PMF, which assumes both encoders share the same Transformer architecture, making it hard to be extended to other modalities and other model architectures.
+Moreover, our method is built upon weak assumption than PMF, which assumes both encoders share the same Transformer architecture with an ad-hoc interaction layer.
+Their method is hard be extended for the task with three modalities, such as MUStARD.
+Extending PMF to heterogeneous model architectures is also are challenge.
+By contrast, our method is compatible with more than two modality and different architecture.
 Referring Expression Segmentation.
-Previous prompt fusion methods mainly test their performance on classification (Li et al.
-2023; Liang, Zhao, and Schütze 2022; Tsimpoukelli et al.
-2021).
-The improved expressiveness of MoPE allows learning of harder tasks, for example, referring image segmentation.
-Following experiment setting in VPT, We additionally finetune a SeTR-PUP (Zheng et al.
-2021) head for segmentation.
+Previous prompt fusion methods mainly test their performance on classification [20, 21, 34].
+The improved expressiveness of MoPE allows learning of dense prediction tasks, for example, referring image segmentation.
 The result is summarized in Tab.
 2.
-Our proposed method achieves the best result among all compared methods.
-In particular we outperform the fine-tuning baseline SeqFuse, which suffers from catastrophic forgetting problem on this task.
-Qualitative Result of MoPE Routing
+Our proposed method achieves the best result among all compared methods by a significant margin.
+In particular the proposed method even outperform the fine-tuning baseline SeqFuse using only 15.3% parameters.
+This experiment further demonstrate the improved expressiveness of proposed method.
+4.4 Qualitative visualization of MoPE Routing
 Through training, prompt experts in MoPE spontaneously specialize, and we present two examples in Fig.
@@
 In the provided examples, we observe that expert-4 specialize in children-related concepts, while expert-12 focuses on crowds.
-These examples also demonstrate the interpretability of MoPE, as opposed to black-box conditional prompt such as CoOp (Zhou et al.
-2022b).
+These examples also demonstrate the interpretability of MoPE, as opposed to black-box conditional prompt such as CoOp [43].
 More visualizations could be found in supplementary material.
-Analysis and Discussion
+5 Analysis and Discussion
 In this section, we systematically analyze the effects of our proposed design through ablation studies.
@@
 Our decomposed prompts are collaborative.
-We ablate each type of prompt in our instance-wise prompt decomposition, with results across 3 random seeds reported in Tab.
-3 .
+We ablate each type of prompt in our instance-wise adaptive prompt decomposition, with results across 3 random seeds reported in Tab.
+3.
 Our full method achieves the best performance, indicating effective prompt collaboration.
-Adding the dynamic prompt yields gains of 13.5%, 13.26%, 8.4%/6.5% on all datasets, this is different from previous methods (Liang, Zhao, and Schütze 2022; Tsimpoukelli et al.
-2021) where modalities don’t explicitly interact.
+Adding the dynamic prompt yields gains of 13.5%, 13.26%, 8.4%/6.5% on all datasets, this is fundamentally different from previous methods [21, 34] where modalities don’t explicitly interact.
 Without the mapped prompt, fine-grained information from complementary modalities is lost, causing significant drops.
-The static prompt is also necessary, acting as a special expert always routed for capturing global-level features.
+The static prompt is also necessary, it could be interpreted as a special expert always routed to capture global-level features.
+Adding this prompt allows the other experts to focus more on capturing instance-level features, leading to a slight performance gain.
 Figure 4: More experts v.s.
@@
 We compare increasing the number of experts, $k$, versus lengthen prompt, $l$.
-Expert-scaling consistently outperforms length-scaling, exhibiting a linear growth trend and do not deteriorate even when $k\to\infty$.
-Conversely, excessive prompt length has a negative impact on accuracy and speed.
+Expert-scaling consistently outperforms length-scaling, exhibiting a linear growth trend.
+Conversely, length-scaling suffer from deterioration.
 Expert-scaling is more expressive than length-scaling.
-To study the expressiveness of MoPE, we compare expert-scaling with length-scaling.
-Our starting point for MoPE is $l=6$ prompts and $k=2$ experts, which account for $(2+1)\times 6+1=19$ tunable prompts in total per-layer.
+We compare expert-scaling with length-scaling.
+Our starting point for MoPE is $l=6$ prompts and $k=2$ experts (when $k\leq 1$, the MoPE degenerates into vanilla prompt), which account for $(2+1)\times 6+1=19$ tunable prompts in total per-layer.
 We increase the number of $k$, and at each step we also report the result of increasing $l$ in vanilla prompt to the same total number of prompts.
@@
 4.
-As illustrated in the figure, adding the MoPE design with as few as $k=2$ experts boosts performance, and increasing the number of experts leads to a monotonic performance improvement which avoids over-length deterioration.
-In contrast, length-scaling does not lead to a linear performance improvement and is prone to performance deterioration, which is consistent with previous findings (Jia et al.
-2022; Yang et al.
-2022a; Khattak et al.
-2023; Li and Liang 2021).
+Adding the MoPE design with as few as $k=2$ experts boosts performance, and increasing the number of experts leads to a monotonic performance improvement that avoids over-length deterioration.
+In contrast, length-scaling does not lead to a linear performance improvement and is prone to performance deterioration, which is consistent with previous findings [11, 40, 14, 19].
 Additionally, longer prompts in vanilla prompt tuning can exacerbate computational overhead due to the quadratic time complexity of self-attention.
-This is why we cannot experiment with $l>199$ for the other two methods.
-By contrast, MoPE scales by conditioning the dynamic prompt on more experts while maintaining a fixed prompt length during actual self-attention.
+By contrast, MoPE scales by conditioning the dynamic prompt on more experts.
+As a result, we scale up the empirical expressiveness while maintaining a fixed prompt length during actual self-attention.
 As a result, our method maintains a nearly constant time complexity.
@@
 Our method outperform other prompt fusion methods at all data scales.
+The performance gap between the best compared prompt fusion method is shaded.
 MoPE scales better with more training data.
-Previous prompting methods have been shown to not scale well with respect to increased training data (Gao et al.
-2022; Liang, Zhao, and Schütze 2022; Tsimpoukelli et al.
-2021; Yang et al.
-2022a).
+Previous prompting methods have been shown to not scale well with respect to increased training data [9, 21, 34, 40].
 To assess the scalability of MoPE, we sub-sample the training set with different shots to simulate a range of low-shot to high-shot learning scenarios.
 We then train our method and other prompt-tuning and fine-tuning methods on the same subset of SNLI-VE and MM-IMDB.
-As Fig.
-5 shows, our MoPE-based method demonstrates superior scalability compared to other prompt tuning methods.
+Our MoPE-based method demonstrates superior scalability compared to other prompt tuning methods, as demonstrated in Fig.
+5.
 Specifically, we consistently match the results of the fine-tuning method, SeqFuse, on the SNLI-VE dataset, while surpassing it on the MM-IMDB dataset.
-By contrast, the methods based on vanilla prompts do not scale well with respect to increased training data, resulting in a consistent performance gap between our method and the compared methods.
+By contrast, the methods based on vanilla prompts, PromptFuse and P-SeqFuse are less scalable with respect to increased training data, resulting in a consistent performance gap between our method and the compared prompt fusion methods.
 This performance disparity becomes more pronounced on larger datasets, such as when training with $10^{5}$ shots or using the full training set.
+This result demonstrates the effect of proposed MoPE in scaling up the expressiveness of prompt fusion.
 Figure 6: Effect of the importance loss.
@@
 (a) Without importance loss, only a few experts are used throughout training (b) The importance loss ensures balanced utilization of all experts.
+Our prompt fusion method is highly modular.
+The proposed fusion method abstracts the complimentary modality as a representation, allowing high modality of both modalities.
+This is different from PMF [20], which is not modular.
+In particular, our method allow arbitrary models to be seamless plug-ins for multimodal fusion.
+Such modularity is at least threefold: model architecture, the pre-training scheme, and the specific transfer learning method of the model.
+We exemplify each in Tab.
+4.
+Architecture Pretraining Transfer MM-IMDB F1($\uparrow$) BoW Bert* Fine-tuning 48.20/57.50 Transformer Bert [5] Frozen 58.86/66.13 Transformer GPT-2 [30] Frozen 34.03/50.84 Transformer Bert [5] Fine-tuning 60.34/67.27
+Table 4: Our prompt tuning method are highly modular.
+We offer flexibility for complementary modality in at least three dimensions: model architecture, pretraining scheme, as well as transfer learning technique.
+(*): Bag-of-word initialized with Bert word embeddings.
 Frozen routing embedding outperform learned embedding.
@@
 As Tab.
-4 shows, fixed routing embedding could achieve the same result as the trained one.
-Routing Embed SNLI-VE $\uparrow$ UPMC Food $\uparrow$ MM-IMDB $\uparrow$
+5 shows, fixed routing embeddings could slightly better performance, while not requiring training.
+Routing Embd SNLI-VE Acc($\uparrow$) UPMC Acc($\uparrow$) Food $\uparrow$ MM-IMDB F1($\uparrow$)
 Frozen 73.55 91.74 62.01/68.25
@@
 By contrast, our dynamic prompt with MoPE is adaptive, and finding an instance-wise optimal prompt can lead to better results.
-Secondly, MoPE circumvents the challenge of optimizing a long prompt (Petrov, Torr, and Bibi 2023).
+Secondly, MoPE circumvents the challenge of optimizing a long prompt [28].
 By doing so, we shift the difficulty from learning a universal long prompt to learning a parameterized router with multiple short prompts.
 As a result, our empirical expressiveness is closer to the theoretical upper bound of prompt tuning.
-Conclusion
-In this paper, we propose an adaptive and expressive prompt-based method for fusing unimodal models.
-Our method involves decomposing the prompt for instance-adaptive prompt learning.
+Limitations and future works.
+Despite improving empirical expressiveness, our prompt still functions in the same way as the vanilla prompts during the self-attention.
+This means that our method is closer to the theoretical upper-bound of expressiveness for prompt tuning but not surpass it.
+This is also reflected by the small gap between ours and fine-tuning in large datasets such as SNLI-VE and UPMC Food-101.
+Future work could focus on augmenting the way that prompt attends to the token embedding to scale up its theoretical expressiveness.
+6 Conclusion
+In this paper, we propose an method to mitigate the issue of lacking adaptiveness and expressiveness in existing prompt-based multimodal fusion method.
+Our method involves decomposing the unified long prompt for instance-adaptive prompt learning.
 To effectively scale up the prompt while avoiding performance deterioration, we introduce MoPE, which improves the expressiveness of prompt tuning.
 Extensive experiments demonstrate that our method is highly parameter-efficient and scales better with dataset size and prompt length.
-We believe MoPE could be an efficient fusion method for a wide array of multimodal tasks.
+We believe MoPE could be an efficient fusion method for a wide array of downstream multimodal tasks.
 Appendix A Additional Details on Experiments
 In this section, we provide additional details on our model architecture, experimental setup, and data preprocessing.
-Architecture Details
+A.1 Architecture Details
 Prompt Vector.
-Our prompt implementation closely follows VPT (Jia et al.
-2022).
-Specifically, for static prompts and prompt experts, we use uniform initialization $\mathcal{U}\sim(-\eta,\eta)$, where $\eta$ is calculated according to the embedding dimension and patch size of the Transformer (Jia et al.
-2022).
+Our prompt implementation closely follows VPT [11].
+Specifically, for static prompts and prompt experts, we use uniform initialization $\mathcal{U}\sim(-\eta,\eta)$, where $\eta$ is calculated according to the embedding dimension and patch size of the Transformer [11].
 Dropout with $p=0.1$ is applied to all prompts.
-However, we do not use the reparameterization trick for prompts introduced in the original prompt tuning methods (Li and Liang 2021), as the gradients of our dynamic prompt and mapped prompt are already rectified by MLPs (i.e., $\mathbf{W}_{x}$, $\mathbf{W}_{y}$, $f_{m}(\cdot)$).
-For Transformer architectures that employ a window attention mechanism (e.g., Swin (Liu et al.
-2021b)), we duplicate the same prompt and prepend it to all windows for self-attention calculation, following the approach in VPT (Jia et al.
-2022).
+However, we do not use the reparameterization trick for prompts introduced in the original prompt tuning methods [19], as the gradients of our dynamic prompt and mapped prompt are already rectified by MLPs (i.e., $\mathbf{W}_{x}$, $\mathbf{W}_{y}$, $f_{m}(\cdot)$).
+For Transformer architectures that employ a window attention mechanism (e.g., Swin [25]), we duplicate the same prompt and prepend it to all windows for self-attention calculation, following the approach in VPT [11].
 Mapper.
 We learn a mapper to map representations from complementary modality $\mathbb{Y}$ to the embedding dimension of main modality $\mathbb{X}$.
-Generally speaking, the mapper is implemented as a two-layer MLP with a bottleneck design, which shares similarities with previous work (Li et al.
-2023).
+Generally speaking, the mapper is implemented as a two-layer MLP with a bottleneck design, which shares similarities with previous work [20].
 In our experiments, we set the bottleneck dimension as half of the dimension of the complementary representation, i.e., $d_{bot}=\lceil d_{y}/2\rceil$.
@@
 However, some Transformer architectures use inconsistent $d_{x}$ in different layers.
-For example, in Swin-b (Liu et al.
-2021b), the embedding dimension changes from $[128,256,512,1024]$, making it challenging to fit a single linear transformation.
+For example, in Swin-b [25], the embedding dimension changes from $[128,256,512,1024]$, making it challenging to fit a single linear transformation.
 To circumvent this, we instead learn four separate linear projections, each used to project the shared bottleneck feature to different embedding dimensions.
@@
 For the importance loss, we scale it by a factor of $0.01$ and linearly combine it with the task-specific loss(es) for optimization.
-Dataset
-UPMC Food-101 (Wang et al.
-2015) serves as a comprehensive multimodal dataset designed for fine-grained recipe classification.
+A.2 Dataset
+UPMC Food-101 [36] serves as a comprehensive multimodal dataset designed for fine-grained recipe classification.
 The dataset contains 90,840 image-text pairs for 101 food classes.
-We follow previous methods (Li et al.
-2023; Kiela et al.
-2019) to create a validation set of 5000 samples.
-MM-IMDB (Arevalo et al.
-2017) is a multimodal movie classification dataset.
+We follow previous methods [20, 15] to create a validation set of 5000 samples.
+MM-IMDB [1] is a multimodal movie classification dataset.
 It comprises 25,956 pairs of images and texts, each pair including a movie poster and a plot summary.
 The dataset supports multi-label genre classification across a spectrum of 23 genres with highly imbalanced classes.
-SNLI-VE (Xie et al.
-2019) is a large-scale multimodal dataset with 565,286 image-text pairs.
+SNLI-VE [39] is a large-scale multimodal dataset with 565,286 image-text pairs.
 The task for this dataset is visual entailment, which means that the model should decide whether a hypothesis matches the given premise or not.
 This dataset provide image and text premise, while the hypothesis is always in text modality.
-MUsTARD (Castro et al.
-2019) is a dataset for multimodal sarcasm detection.
+MUsTARD [4] is a dataset for multimodal sarcasm detection.
 The original dataset contains 690 video clips in MP4 format with an even number of sarcastic and non-sarcastic labels.
@@
 Collectively, there are three modalities available: video, audio extracted from the video, and the language modality.
-RefCOCO and RefCOCO+ (Kazemzadeh et al.
-2014) are two datasets for referring image segmentation.
+RefCOCO and RefCOCO+ [13] are two datasets for referring image segmentation.
 Both datasets are built based on MSCOCO and contain around 19,900 images with 50,000 objects and 140,000 referring expressions.
@@
 Figure 8: Examples of each dataset used in the main paper.
-Data Processing
+A.3 Data Processing
 For all input images, we perform RandAug with $N=2$ and $M=5$, resize them to $(256,256)$, and perform center cropping to obtain images of size $(224,224)$.
@@
 Note that this setting may differ from other works that also use the text premise.
-For the MuSTARD dataset, the overall processing pipeline follows PromptFuse (Liu et al.
-2022).
+For the MuSTARD dataset, the overall processing pipeline follows PromptFuse [23].
 For the audio modality, we extract WAV audio from the video using a sampling rate of 16,000 Hz.
@@
 For a batch of input, we zero-pad the WAV input to the same length.
-As for the video, while PromptFuse (Liu et al.
-2022) uses a face detection model to only sample frames where the speaker’s face is visible, we evenly sample $8$ frames for all videos in time.
+As for the video, while PromptFuse [23] uses a face detection model to only sample frames where the speaker’s face is visible, we evenly sample $8$ frames for all videos in time.
 The sampled frames are resized to $(224,224)$ for encoding.
@@
 This section aims to provide further ablation study on the design choices and hyper-parameters. We also provided extended discussion on the effectiveness of proposed MoPE.
-Additional Ablations
+B.1 Additional Ablations
 Ablation on dimension of $d_{i}$ and $d_{c}$.
 Our finding indicate that $d_{c}>>d_{i}$ results in better performance in general, and we set $d_{c}=8,d_{i}=2$ in our main experiment as ablated in Tab.
-5.
+6.
 $k=4$ $d_{i}=10$, $d_{c}=0$ $d_{i}=8$, $d_{c}=2$ $d_{i}=5$, $d_{c}=5$ $d_{i}=2$, $d_{c}=8$ $d_{i}=0$, $d_{c}=10$
@@
 In main experiment we use swin-base as the vision encoder.
-To have a fair comparision on previous prompt fusion methods (Li et al.
-2023; Liang, Zhao, and Schütze 2022) that are based on ViT (Dosovitskiy et al.
-2020), we provide the result of our method with ViT-base as vision encoder in Tab.
-6, with results measured using 3 random seeds, As the table shows, using the ViT as vision encoder gives similar results.
+To have a fair comparision on previous prompt fusion methods [20, 21] that are based on ViT [6], we provide the result of our method with ViT-base as vision encoder in Tab.
+7, with results measured using 3 random seeds, As the table shows, using the ViT as vision encoder gives similar results.
 The result demonstrate that our superior performance is not due to more advanced vision encoder.
@@
 Table 7: Ablation on ViT as the main vision encoder.
-Analysis on Adaptivity of Vanilla Prompt and MoPE
+B.2 Analysis on Adaptivity of Vanilla Prompt and MoPE
 In the main body of the paper, we have empirically demonstrated that the dynamic prompt generated by MoPE exhibits greater adaptivity compared to global-shared prompt tuning.
@@
 Through this analysis, we will also provide intuitive insights into the significance of expert specialization in MoPE-based fusion.
-Following the approach adopted in previous papers (Wang et al.
-2023; Petrov, Torr, and Bibi 2023), we will focus our analysis on the case of a single prompt within a single Transformer layer.
+Following the approach adopted in previous papers [37, 28], we will focus our analysis on the case of a single prompt within a single Transformer layer.
 Since our method is designed to generate an effective prompt while leaving the attention calculation the same as previous prompt tuning, our MoPE will only affect the forward behavior up to the calculation of attention map.
@@
 The objective is to find a prompt $\mathbf{P}$ that enables the attention map $\mathbf{A}(\mathbf{X},\mathbf{P})$ to closely match the desired target attention pattern for each input instance.
-Theorem 1 (Limited Adaptivity of Global-shared Prompt)
+Theorem 1 (Limited Adaptivity of Global-shared Prompt).
 Let $\mathcal{X}$ be the input space, and $\mathcal{A}$ be the space of attention matrices.
@@
 Then, for vanilla prompting with a single shared prompt $\mathbf{P}$, there exists no $\mathbf{P}\in\mathcal{P}$ such that $\mathcal{A}^{*}(\mathbf{x})\subseteq\mathcal{A}(\mathbf{x},\mathbf{P})$ for all $\mathbf{x}\in\mathcal{X}$, where $\mathcal{P}$ is the prompt space.
-Proof B.1
+Proof.
 Let $\mathbf{x}_{1},\mathbf{x}_{2}\in\mathcal{X}$ be two distinct input instances with disjoint optimal attention sets, i.e., $\mathcal{A}^{*}(\mathbf{x}_{1})\cap\mathcal{A}^{*}(\mathbf{x}_{2})=\emptyset$.
@@
 $\Delta(\mathbf{x}_{1},\mathbf{P}_{\text{shared}}^{*})+\Delta(\mathbf{x}_{2},\mathbf{P}_{\text{shared}}^{*})\geq\Delta(\mathbf{x}_{1},\mathbf{P}_{1}^{*})+\Delta(\mathbf{x}_{2},\mathbf{P}_{2}^{*})$ (10)
-Furthermore, by the limited expressiveness of prompt-tuning (Petrov, Torr, and Bibi 2023), we have:
+Furthermore, by the limited expressiveness of prompt-tuning [28], we have:
 $\Delta(\mathbf{x}_{2},\mathbf{P}_{2}^{*})\geq 0;\\ \Delta(\mathbf{x}_{1},\mathbf{P}_{1}^{*})\geq 0$ (11)
@@
 $\Delta(\mathbf{x}_{1},\mathbf{P}_{\text{shared}}^{*})+\Delta(\mathbf{x}_{2},\mathbf{P}_{\text{shared}}^{*})\geq\Delta(\mathbf{x}_{1},\mathbf{P}_{1}^{*})+\Delta(\mathbf{x}_{2},\mathbf{P}_{2}^{*})\geq 0$ (12)
-Theorem 2 (Improved Adaptivity of MoPE)
+Theorem 2 (Improved Adaptivity of MoPE).
 Let =$\mathcal{E}=\{\mathcal{E}_{1},\mathcal{E}_{2},\ldots,\mathcal{E}_{k}\}$ be a set of $k$ expert attention mappings, where each $\mathcal{E}_{i}:\mathcal{X}\rightarrow\mathcal{A}$ maps an input $\mathbf{x}\in\mathcal{X}$ to a set of attention patterns $\mathcal{E}_{i}(\mathbf{x})\subseteq\mathcal{A}$.
@@
 $\sum_{\mathbf{x}\in\mathcal{X}^{\prime}}\Delta(\mathbf{x},\mathcal{E},\mathbf{r}^{*})=\sum_{\mathbf{x}\in\mathcal{X}^{\prime}}\inf_{\mathbf{A}\in\mathcal{A}^{*}(\mathbf{x})}\|\mathbf{A}-\mathbf{A}^{*}\|_{\mathcal{A}}$ (15)
-Proof B.2
+Proof.
 When there are more experts than instances, i.e., $|\mathcal{X}^{\prime}|\leq k$ , the proof is trivial.
@@
 Thus, when the number of instances exceeds the number of experts, if the convex hull of the expert attention mappings contains an optimal attention pattern for each instance, then there exists a routing score $\mathbf{r}^{*}$ that allows MoPE to achieve an accumulated attention discrepancy equal to the sum of the optimal instance-wise attention discrepancies across those instances.
+∎
 Putting it together, the Theorem.
@@
 2 state that it is possible for MoPE to achieve the best result on all of the input instances, conditioned on the appropriate specialization of experts.
-Discussion: How to Choose the Main Modality?
+B.3 Discussion: How to Choose the Main Modality?
 In the proposed method, we use a sequential pipeline to fuse different modalities.
@@
 We have also tested using text as the main modality, and the results are summarized in Tab.
-7.
+8.
 Method SNLI-VE Food-101 MM-IMDB
@@
 Our postulation is that for these three tasks, the text input contains a significant amount of noisy and false positive data.
-For instance, in the UPMC-Food 101 dataset (Wang et al.
-2015), the text data is derived using a spider, which includes many irrelevant hypertext and website titles.
+For instance, in the UPMC-Food 101 dataset [36], the text data is derived using a spider, which includes many irrelevant hypertext and website titles.
 Therefore, treating the text as the main modality can lead to overfitting and biased predictions.
@@
 For example, in referring expression segmentation, we have to use the image as main modality, because the segmentation head expects feature maps instead of a global representation.
-Our MoPE-based method yields a better multimodal representation.
+B.4 Our MoPE-based method yields a better multimodal representation.
 We visualize the multimodal representation generated by different methods using t-SNE.
@@
 For the other methods, we visualize the [CLS] token of the main modality.
-For ease of coloring, we only show the first 20 classes in the test set of the UPMC Food-101 (Wang et al.
-2015) dataset.
+For ease of coloring, we only show the first 20 classes in the test set of the UPMC Food-101 [36] dataset.
 The results are presented in Fig.
@@
 Appendix C Additional Visualizations
-Visualization of Referring Expression Segmentation Results
+C.1 Visualization of Referring Expression Segmentation Results
 In Fig.
 10, Fig.
-11 we provide a qualitative comparision on our method and the P-SeqFuse baseline.
+11 we provide a qualitative comparison on our method and the P-SeqFuse baseline.
 As illustrated in the figures, our methods could correctly understand the referring expression and localize object mask accordingly.
@@
 Figure 11: Visualization on Referring Expression Segmentation - 2
-Additional Examples of Expert Routing on VQAv2
+C.2 Additional Examples of Expert Routing on VQAv2
 Our MoPE is designed to scale up the expressiveness of vanilla prompts, and expert specialization is a critical condition to achieve superior results as well as interpretability.
@@
 We found the expert specialization are more observable on large dataset with a high heteriogenity.
-Hence, to further demonstrate expert specialization, we train our MoPE-based method on an even larger dataset, VQAv2 (Goyal et al.
-2017).
+Hence, to further demonstrate expert specialization, we train our MoPE-based method on an even larger dataset, VQAv2 [10].
 This dataset contains 265,016 images and paired questions, covering a wide array of visual and textual concepts.